### v1

# Computable Thermodynamic Formalism

Abstract.
We investigate the theory of thermodynamic formalism from the perspective of computable analysis, with a special focus on the computability of equilibrium states. Specifically, we develop two complementary general approaches to verify the computability of equilibrium states for non-uniformly expanding computable dynamical systems. The first approach applies to dynamical systems whose topological pressure functions admit effective approximations and whose measure-theoretic entropy functions are upper semi-continuous. As a concrete application, we establish the computability of the equilibrium states for Misiurewicz–Thurston rational maps with Hölder continuous potentials. The second approach exploits prescribed Jacobians of equilibrium states through a local analysis and applies to settings where the measure-theoretic entropy functions may lack upper semi-continuity.
Key words and phrases:
computability, computable analysis, equilibrium state, thermodynamic formalism, expanding Thurston map, Misiurewicz–Thurston rational map.
2020 Mathematics Subject Classification:
Primary: 03D78, 37D35; Secondary: 37D25, 37F15.
1. Introduction
The study of computability in dynamical systems lies at the intersection of mathematics, physics, and computer science, and has become increasingly vital for understanding complex physical phenomena through computational perspectives. A fundamental insight driving this field is that while chaotic systems exhibit sensitive dependence on initial conditions, their typical statistical behaviors are often described by computable objects. This interplay between chaos and computability underscores a fundamental principle: the more expansive and chaotic a system’s dynamics, the more tractable its typical behavior becomes.
This perspective builds on the well-established principle that higher dynamical complexity, particularly through uniform expansion or hyperbolicity, facilitates the analysis of associated dynamical invariants. In such uniformly expanding systems, the dynamical mechanisms driving long-term behavior become sufficiently regular to facilitate effective computation of key quantities such as entropy, pressure, and key invariant measures. Thus, the central challenge, and our primary focus, lies in extending these computational methodologies to systems where uniform expansion fails.
Thermodynamic formalism emerges as the natural framework for studying chaotic dynamics from a statistical viewpoint. This theory, which draws inspiration from statistical mechanics, was pioneered by Ruelle, Sinai, Bowen, and others in the 1970s ([Do68, Si72, Bo75, Wa82]). Since its birth, thermodynamic formalism has been extensively applied in various classical contexts and has remained at the frontier and core of research in dynamical systems. It focuses on equilibrium states, which are invariant measures that maximize pressure functionals encoding both entropy [^1] The concepts of entropy in dynamical systems have their roots in the early works on the notions of entropy by Boltzmann and Gibbs (statistical mechanics 1875), von Neumann (quantum mechanics 1932), and Shannon (information theory 1948). These notions of entropy are all designed to describe the complexity of their respective systems or objects. In recent years, there have been exciting developments and diverse applications of entropy and complexity theory, see e.g. Braverman’s report at the International Congress of Mathematicians in 2022 [Bra23]. and integrals of potentials. Crucially, equilibrium states characterize the typical behavior of the dynamics and thus possess central mathematical and physical significance. In many settings, equilibrium states describe the weighted distribution of random backward orbits (see e.g. [HT03, Li18]), iterated preimages, and periodic orbits.
In complex dynamics, Brolin–Lyubich measures [Bro65, Ly82] are measures of maximal entropy for rational maps. A uniform algorithm to compute such measures was developed in [BBRY11]. This complements the discovery of polynomials with computable coefficients but non-computable Julia sets, as explored in the pioneering works of Braverman and Yampolsky [BY06, BY09], which can be traced back to a question posed by Milnor (see [BY06, Section 1]). For further research on algorithmic aspects of Julia sets, we refer the reader to recent works of Rojas and Yampolsky [RY21b] and Dudko and Yampolsky [DY21] and references therein.
The computability of Brolin–Lyubich measures presents an apparent paradox. Intuitively, one might expect a measure to contain more information than its support. However, computable analysis reveals the existence of a computable invariant probability measure (the Brolin–Lyubich measure) whose support is non-computable. Indeed, this paradox can be reconciled by interpreting these results as reflecting distinct computability properties of the system from geometric and statistical perspectives. From this statistical viewpoint, questions regarding the computability of equilibrium states via thermodynamic formalism gain critical significance.
In this article, we study thermodynamic formalism from the point of view of computable analysis, with a special focus on the computability of equilibrium states in dynamical systems. Classical hyperbolic systems admit well-developed techniques and are generally regarded as well-understood. By contrast, non-uniformly hyperbolic systems resist conventional approaches. The relaxation of uniform expansion requirements therefore represents a fundamental challenge in dynamical systems. Our central contribution lies in demonstrating that statistical computability can coexist with dynamical complexity, even in non-uniformly hyperbolic regimes.
We develop two complementary approaches that establish fundamental links between the computability of thermodynamic quantities and equilibrium states. Each approach offers distinct advantages and ranges of applicability, designed to cover a broad class of dynamical systems.
The first approach utilizes the set of tangent functionals to the topological pressure function at the potential to link the computability of the topological pressure to that of the equilibrium state. This method is suitable for dynamical systems where the measure-theoretic entropy map is upper semi-continuous and the topological pressure function can be effectively approximated. For dynamical systems whose measure-theoretic entropy functions may lack upper semi-continuity, we implement the second approach, which establishes the computability of equilibrium states through local (rather than global) considerations. The novelties of our second approach include the following two aspects: (i) on the thermodynamic formalism side, using general transfer operators to study equilibrium states of full measure away from singular points with prescribed Jacobians, and (ii) on the computable analysis side, the verification of Jacobians away from singular points, and the construction of a recursively compact set of measures to exclude atomic measures supported on singular points.
To demonstrate applications of these approaches, we focus on expanding Thurston maps as primary case studies, whose ergodic theory has been actively studied. A Thurston map is a branched covering map on a topological $2$-sphere that is not a homeomorphism and satisfies the postcritically-finite condition (i.e., every critical point has a finite forward orbit). Such maps play a central role in the study of complex dynamics, and the most important examples are given by postcritically-finite rational maps on the Riemann sphere [^2] There has been active research on algorithmic aspects of Thurston maps. For example, Bonnet, Braverman, and Yampolsky [BBY12] devised an algorithm to determine whether a Thurston map is Thurston equivalent to a rational map; Hubbard and Schleicher [HS94], in the setting of quadratic rational maps, provided an algorithm that, given a convenient description of the Thurston map, outputs the coefficients of the rational map; Selinger, Rafi, and Yampolsky [SY15, RSY20] investigated the decidability of Thurston equivalence. . Inspired by Sullivan’s dictionary and their interest in Cannon’s Conjecture [Ca94], Bonk and Meyer [BM10, BM17], as well as Haïssinsky and Pilgrim [HP09], studied a subclass of Thurston maps, called expanding Thurston maps, by imposing some additional condition of weak expansion (see Definition 6.2).
Based on these works, ergodic theory for expanding Thurston maps has been actively investigated in [BM10, BM17, HP09, Li15, Li17, Li18, LZ25, LSZ25, LS24a, LS24b]. In particular, the third-named author [Li18] developed the thermodynamic formalism and investigated the existence, uniqueness, and other ergodic properties of equilibrium states for expanding Thurston maps333cf. the monograph [Li17].. Notably, expanding Thurston maps exhibit weak expansion; they are neither expansive nor $h$-expansive. Furthermore, those with at least one periodic critical point are not even asymptotically $h$-expansive [Li15]. Recent advances by the third-named and fourth-named authors [LS24b] further demonstrated that the measure-theoretic entropy function is upper semi-continuous if and only if the expanding Thurston map has no periodic critical points. Leveraging these results, we apply our methods to investigate the computability of equilibrium states for expanding Thurston maps, demonstrating the applicability of our two distinct approaches.
1.1. Main results
As mentioned earlier, two complementary approaches link the computability of thermodynamic quantities to that of equilibrium states.
The first approach (Theorem 1.1) applies to systems with upper semi-continuous entropy. It shows that certain computability properties of the topological pressures guarantee the computability of the equilibrium states, with a concrete application to Misiurewicz–Thurston rational maps in Theorem 1.2. The second approach, established in Theorem 1.3, applies to systems without upper semi-continuous entropy, where the computability of equilibrium states arises from the computability of their prescribed Jacobians. As an application, we establish the computability of the measures of maximal entropy for computable expanding Thurston maps with computable critical points in Theorem 1.4.
We adopt the conventions and terminology for computable analysis from [We00] and refer the reader to Section 3 for a more detailed introduction. Below, we introduce the uniformly computable systems that form the setting for our approaches.
We say that the quadruple $(X,\,\rho,\,\mathcal{S},\,\{X_{n}\}_{n\in\mathbb{N}},\,\{T_{n}\}_{n\in\mathbb{N}})$ is a uniformly computable system if the following properties are satisfied:
  1. (i)
    $(X,\,\rho,\,\mathcal{S})$ is a computable metric space in which $X$ is recursively compact.
  2. (ii)
    $X_{n}\subseteq X$ is open and $T_{n}\colon X\rightarrow X$ is a Borel-measurable function for each $n\in\mathbb{N}$.
  3. (iii)
    In $(X,\,\rho,\,\mathcal{S})$, $\{X_{n}\}_{n\in\mathbb{N}}$ is a sequence of uniformly lower semi-computable open sets and $\{T_{n}\}_{n\in\mathbb{N}}$ is a sequence of uniformly computable functions with respect to $\{X_{n}\}_{n\in\mathbb{N}}$.
Here the computable structure of metric spaces is given in Definition 3.4, the lower semi-computable openness of a set is an effective version of openness given in Definition 3.7, and the recursive compactness of a set is an effective version of compactness (with an additional algorithmic covering procedure); see Definition 3.18. Moreover, the computability of functions is given in Definition 3.12.
In the first approach, we exploit the set of tangent functionals to the topological pressure function $P_{T}$ (see Section 4) at the potential $\phi$ and demonstrate that the computability of the equilibrium state follows from the certain computability properties of the topological pressure. We note that this set of tangent functionals is naturally identified with a subset of the set $\mathcal{P}(X)$ of Borel probability measures (see Remark 4.4). We denote by $\mathcal{E}(T,\phi)$ the set of equilibrium state(s) for a map $T$ and a potential $\phi$. We refer the reader to Subsection 3.1 for a detailed discussion on the computability of measures.
Theorem 1.1.
Let $(X,\,\rho,\,\mathcal{S},\,\{X_{n}\}_{n\in\mathbb{N}},\,\{T_{n}\}_{n\in\mathbb{N}})$ be a uniformly computable system with $X_{n}=X$ for all $n\in\mathbb{N}$, and $\{\phi_{n}\}_{n\in\mathbb{N}}$ be a sequence of uniformly computable functions $\phi_{n}\colon X\rightarrow\mathbb{R}$. Suppose $T_{n}$ has finite topological entropy, and the measure-theoretic entropy map $\nu\mapsto h_{\nu}(T_{n})$ is upper semi-continuous on $\mathcal{M}(X,T_{n})$ for each $n\in\mathbb{N}$. Assume that the following statements are true:
  1. (i)
    There exists a sequence $\{\psi_{n,i}\}_{(n,i)\in\mathbb{N}^{2}}$ of uniformly computable functions $\psi_{n,i}\colon X\rightarrow\mathbb{R}$ such that the closure $\overline{D}_{n}$ of $D_{n}\coloneqq\{\psi_{n,i}:i\in\mathbb{N}\}$ in $C(X)$ contains a neighborhood of $\phi_{n}$ for each $n\in\mathbb{N}$ and $\{P(T_{n},\psi_{n,i})\}_{(n,i)\in\mathbb{N}^{2}}$ is a sequence of uniformly upper semi-computable real numbers.
  2. (ii)
    $\{P(T_{n},\phi_{n})\}_{n\in\mathbb{N}}$ is a sequence of uniformly lower semi-computable real numbers.
  3. (iii)
    There exists a unique equilibrium state $\mu_{n}$ for $T_{n}$ and $\phi_{n}$ for each $n\in\mathbb{N}$.
Then $\{\mu_{n}\}_{n\in\mathbb{N}}$ is a sequence of uniformly computable measures.
Here the computability properties of real numbers is given in Definitions 3.3 and 3.6, and the computable structure on the space $\mathcal{P}(X)$ of Borel probability measures is specified in Proposition 3.26.
Applying Theorem 1.1, we establish the computability of the equilibrium states for computable Misiurewicz–Thurston rational maps (i.e., postcritically-finite rational maps without periodic critical points which are computable) and computable Hölder continuous potentials. The existence and the uniqueness of the equilibrium states for expanding Thurston maps follow from Theorem 6.3.
Theorem 1.2.
Let $f\colon\widehat{\mathbb{C}}\rightarrow\widehat{\mathbb{C}}$ be a computable Misiurewicz–Thurston rational map, $\sigma$ be the chordal metric, and $\alpha\in(0,1]$. Assume that $\{\phi_{n}\}_{n\in\mathbb{N}}$ is a sequence of uniformly computable functions on $\widehat{\mathbb{C}}$, and $\{Q_{n}\}_{n\in\mathbb{N}}$ is a sequence of uniformly computable real numbers. Suppose $\phi_{n}\in C^{0,\alpha}\bigl(\widehat{\mathbb{C}},\sigma\bigr)$, $\mathcal{E}(f,\phi_{n})=\{\mu_{n}\}$, and $\mathopen{\lvert}\phi_{n}\mathclose{\rvert}_{\alpha,\sigma}\leqslant Q_{n}$ for each $n\in\mathbb{N}$. Then $\{\mu_{n}\}_{n\in\mathbb{N}}$ is a sequence of uniformly computable measures.
The computable structure in $\widehat{\mathbb{C}}$ is given in Proposition 6.5. Moreover, $\mathopen{\lvert}\phi\mathclose{\rvert}_{\alpha,\sigma}$ denotes the Hölder constant of $\phi\in C^{0,\alpha}\bigl(\widehat{\mathbb{C}},\sigma\bigr)$ (see (2.1)).
It is worth noting that Theorem 1.1 can be applied to establish the computability of the equilibrium states for a wide range of dynamical systems with a unique equilibrium state and an upper semi-continuous measure-theoretic entropy map, such as rational maps with Hölder continuous hyperbolic potentials (see e.g. [DU91]). Due to space limitations, we focus on the current examples and postpone further investigations to future works.
For many dynamical systems, the measure-theoretic entropy map may not be upper semi-continuous, or verifying this property may be difficult. To overcome this issue, we establish a second approach by considering the prescribed Jacobians for equilibrium states.
For a compact metric space $(X,\rho)$, a Borel-measurable transformation $T\colon X\rightarrow X$, we say that $A\subseteq X$ is admissible (for $T$) if $A,\,T(A)\in\mathcal{B}(X)$ and $T|_{A}$ is injective. Given a Borel subset $Y\subseteq X$, and a Borel function $J\colon X\rightarrow[0,+\infty)$, define
$\mathcal{M}(X,T;Y)$ $\coloneqq\bigl\{\mu\in\mathcal{P}(X):\mu\bigl(T^{-1}(A)\cap Y\bigr)\leqslant\mu(A)\text{ for each Borel }A\subseteq X\bigr\},\quad\text{and}$ (1.1)
$\mathcal{M}(X,T;Y,J)$ $\coloneqq\biggl\{\mu\in\mathcal{P}(X):\mu(T(A))\geqslant\int_{A}\!J\,\mathrm{d}\mu\text{ for each admissible set }A\subseteq Y\text{ for }T\biggr\}.$ (1.2)
Theorem 1.3.
Let $(X,\,\rho,\,\mathcal{S},\,\{X_{n}\}_{n\in\mathbb{N}},\,\{T_{n}\}_{n\in\mathbb{N}})$ be a uniformly computable system, and $Y_{n}$ be an open subset of $X_{n}$ for each $n\in\mathbb{N}$. Assume that there exist two recursively enumerable sets K,LK,\,L with $L\subseteq\mathbb{N}\times K$, and a sequence $\{Y_{n,k}\}_{(n,k)\in L}$ of uniformly lower semi-computable open sets in $(X,\,\rho,\,\mathcal{S})$ with the properties that $Y_{n,k}$ is admissible for $T_{n}$, and $Y_{n}=\bigcup_{(n,k)\in L_{n}}Y_{n,k}$, where $L_{n}\coloneqq\{(n,k)\in L:k\in K\}$ for each $n\in\mathbb{N}$.
Assume that $\{J_{n}\}_{n\in\mathbb{N}}$ is a sequence of uniformly lower semi-computable functions $J_{n}\colon X\rightarrow[0,+\infty)$ with respect to $\{Y_{n}\}_{n\in\mathbb{N}}$ satisfying that $J_{n}$ is nonnegative on $Y_{n}$ and Borel for each $n\in\mathbb{N}$. Suppose there exists a sequence $\{\mathcal{K}_{n}\}_{n\in\mathbb{N}}$ of uniformly recursively compact sets in $(\mathcal{P}(X),\,W_{\rho},\,\mathcal{Q}_{\mathcal{S}})$ such that
$\mathcal{M}(X,T_{n};Y_{n},J_{n})\cap\mathcal{K}_{n}=\{\mu_{n}\}\quad\text{for each }n\in\mathbb{N}.$ (1.3)
Then $\{\mu_{n}\}_{n\in\mathbb{N}}$ is a sequence of uniformly computable measures.
By Theorem 5.6, in some settings (cf. Definition 5.2), the set defined in (1.2) could describe the set of equilibrium states. Moreover, as applications of Theorem 1.3, in Theorems 5.9 and 5.10, we construct two corresponding families of uniformly recursively compact sets $\{\mathcal{K}_{n}\}_{n\in\mathbb{N}}$ satisfying the additional assumptions for two classes of dynamical systems.
Theorem 1.3 extends the methodologies developed in [BBRY11, Theorem A] and [BHLZ25, Theorem 1.1], introducing new techniques to handle both non-uniform expansion and the failure of upper semi-continuity of entropy. Compared to the previous results, our approach significantly broadens applicability by relaxing the computability requirement on prescribed Jacobians to only lower semi-computability.
To demonstrate the applicability of Theorem 1.3 while minimizing unnecessary technicalities, we employ it to prove the computability of the measures of maximal entropy for computable expanding Thurston maps.
Theorem 1.4.
[…]
We provide examples of such expanding Thurston maps and apply Theorem 1.4 to them at the end of Subsection 6.3. It is worth noting that the result in Theorem 1.4 is not uniform and considers only computable expanding Thurston maps on $\widehat{\mathbb{C}}$. Moreover, Theorem 1.3 can be leveraged to extend the above result to the uniform computability of the equilibrium states of computable expanding Thurston maps on a general topological $2$-sphere $S^{2}$. However, such applications would require a discussion on the computability of visual metrics (see e.g. [BM17, Chapter 8]), which would take us too far astray from the core topic of this article.
The conditions in Theorems 1.1 and 1.3 illustrate some methodological distinctions. Specifically, Theorem 1.1 provides an efficient pathway to proving computability when equilibrium states are unique and entropy functions are regular. In contrast, Theorem 1.3 becomes useful for systems with non-unique equilibrium states or those lacking upper semi-continuous entropy, though it demands more sophisticated analysis. Note that computable functions on some subsets are always continuous on the corresponding subsets. Thus, Theorem 1.3 can be applied to some Borel-measurable dynamical systems (not necessarily continuous on the whole space), for example, Pomeau–Manneville maps with geometric potentials. To discuss the topological pressures and equilibrium states in these Borel-measurable settings, we apply Mike Todd’s definition of them in Subsection 3.2.
In our future investigations, the strategies of Theorem 1.3 will be further extended to study the computability of equilibrium states with time complexity in the following settings:
  1. (i)
    Pomeau–Manneville maps with geometric potentials.
  2. (ii)
    Holomorphic endomorphisms of $\mathbb{P}^{k}$ with $\log^{q}$-continuous hyperbolic potential.
  3. (iii)
    Non-uniformly expanding maps on smooth manifolds with Hölder continuous hyperbolic potentials.
1.2. Strategy and organization
In thermodynamic formalism, identifying equilibrium states often hinges on verifying the identity $P(T,\phi)=h_{\mu}(T)+\int\!\phi\,\mathrm{d}\mu$. From the perspective of computable analysis, computing the topological pressure $P(T,\phi)$ and the integral $\int\!\phi\,\mathrm{d}\mu$ is straightforward. The core difficulty lies in evaluating the measure-theoretic entropy function $\mu\mapsto h_{\mu}(T)$, which lacks universal computational methods.
Our approaches reflect two distinct strategies for addressing this challenge. The first approach assumes that the measure-theoretic entropy $h_{\mu}(T)$ coincides with its upper semi-continuous regularization $\overline{h}_{\mu}(T)$ (see Section 4) and uses Lemma 4.6 to characterize $\overline{h}_{\mu}(T)$. The second approach avoids relying on semi-continuity of entropy by instead investigating Jacobians for equilibrium states, a technique motivated by Rokhlin’s formula (cf. Proposition 5.3).
The proofs of the two approaches (Theorems 1.1 and 1.3) follow the same general strategy: we prove the recursive compactness of a subset $\mathcal{K}\subseteq\mathcal{E}(T,\phi)$. Then the assumption that $\mathcal{K}=\{\mu\}$ implies that the equilibrium state $\mu$ is computable (cf. Proposition 3.20 (i)). The strategy for identifying a compact subset $\mathcal{K}\subseteq\mathcal{E}(T,\phi)$ can be summarized as follows: for some dynamical systems, the measure-theoretic entropy has the following upper and lower bounds:
$\overline{h}_{\mu}(T)\geqslant h_{\mu}(T)\geqslant\int\!\log(J_{\mu})\,\mathrm{d}\mu,$
where $J_{\mu}$ is a Jacobian of $T$. Hence, we obtain that
$\{\mu\in\mathcal{M}(X,T):\overline{h}_{\mu}(T)+\mathopen{\langle}\mu,\phi\mathclose{\rangle}=P(T,\phi)\}\supseteq\mathcal{E}(T,\phi)\supseteq\{\mu\in\mathcal{M}(X,T):\mathopen{\langle}\mu,\phi+\log(J_{\mu})\mathclose{\rangle}=P(T,\phi)\}.$
For simplicity, we denote the former (resp. latter) set by $\mathcal{E}_{1}(T,\phi)$ (resp. $\mathcal{E}_{2}(T,\phi)$). Indeed, by some available results in convex analysis (cf. Lemma 4.3 and Propositions 4.5), the set $\mathcal{E}_{1}(T,\phi)$ coincides the set $C(X)^{*}_{\phi,P_{T}}$. Moreover, by our investigations in ergodic theory (cf. Theorem 5.6), the set $\mathcal{E}_{2}(T,\phi)$ can be described by $\mathcal{M}(X,T;Y,J)$ (defined in 1.2), where $Y\subseteq X$ is a Borel set and $J\colon X\rightarrow\mathbb{R}$ is a Borel function satisfying some assumptions (indeed, $\mathcal{E}_{2}(T,\phi)\cap\mathcal{P}(X;Y)$ coincides with $\mathcal{M}(X,T;Y,J)\cap\mathcal{M}(X,T)\cap\mathcal{P}(X;Y)$). It is worth noting that the set $\mathcal{E}(T,\phi)$ of equilibrium states may not be weak compact. However, the sets $C(X)^{*}_{\phi,P_{T}}$ and $\mathcal{M}(X,T;Y,J)$ are always weak compact. Hence, we demonstrate the recursive compactness of these sets instead of investigating $\mathcal{E}(T,\phi)$ directly.
We now outline the proofs of the recursive compactness properties of $C(X)^{*}_{\phi,P_{T}}$ and $\mathcal{M}(X,T;Y,J)$. For the set $C(X)^{*}_{\phi,P_{T}}$, by some results in convex analysis, it is the set of measures $\mu\in\mathcal{M}(X,T)$ such that $\phi$ is the minimizing point of the operator defined by $h\mapsto P(T,h)-\mathopen{\langle}\mu,h\mathclose{\rangle}$. Hence, the recursive compactness of the set $C(X)^{*}_{\phi,P_{T}}$ can be derived from some computability properties of the topological pressure function. Moreover, due to the convexity of such operator, the conditions can be further reduced to the computability properties of the topological pressure function near the potential $\phi$. For the set $\mathcal{M}(X,T;Y,J)$, a local version of a method from [BBRY11, BHLZ25] is applied to check if a Jacobian for $T$ with respect to $\mu$ is greater than or equal to the given function $J$ in the “good” subset $Y$, which is a union of open and admissible sets. It is worth noting that we improve this method and relax the requirements for the domains of computability of dynamical systems.
Moreover, in the proof of Theorem 5.6, the existence of singular points prevents the direct application of Ruelle operators to characterize the prescribed Jacobians of equilibrium states. To address this, we instead use transfer operators to provide an equivalent description of the Jacobians of invariant measures in Theorem 5.5. Combined with the Variational Principle, this yields Theorem 5.6, which allows us to identify equilibrium states by verifying if a Jacobian for $T$ is greater than the prescribed function $J$ in the “good” subset $Y$.
Finally, we summarize the two approaches as follows. In the topological pressure approach, we assume the upper semi-continuity of the pressure function to ensure that $\mathcal{E}_{1}(T,\phi)=\mathcal{E}(T,\phi)$. In the Jacobian approach, the main challenge is constructing a recursively compact set $\mathcal{K}\subseteq\mathcal{P}(X)$ to exclude the measures with positive mass on a “bad” region. Since a uniform construction of such a set $\mathcal{K}$ is not feasible for all systems, we hypothesize its existence in Theorem 1.3 to obtain a general result, and provide explicit constructions for specific families in Theorems 5.95.10, and 1.4. More precisely, we consider the case where the “bad” regions are indeed sets of finitely many points. Theorem 5.9 uses a sequence of open sets containing these “bad” regions to eliminate the mass supported on them, Theorem 5.10 investigates the dynamical systems that are uniformly contracting near all the periodic singular points, and Theorem 1.4 studies expanding Thurston maps, a family of maps which are uniformly expanding near all the periodic singular points.
We now describe the structure of the article. After fixing some notation in Section 2, we review some notions and results in computable analysis and ergodic theory in Section 3.
Section 4 focuses on the first approach. We prove Theorem 1.1 by establishing the recursive compactness of the set of tangent functionals to the topological pressure function at the potential with upper semi-continuous entropy.
Section 5 is devoted to the second approach as stated in Theorem 1.3, which uses the prescribed Jacobian with respect to an equilibrium state. Through employing the transfer operators to establish Theorem 5.6 (in Subsection 5.1), which characterizes an equilibrium state in terms of its Jacobian, we complete the proof of Theorem 1.3 in Subsection 5.2. Theorems 5.9 and 5.10, presented as consequences of Theorem 1.3, are stated in Subsection 5.3.
Section 6 examines the computability of the equilibrium states for expanding Thurston maps. We first provide the definitions and properties of these maps in Subsection 6.1. Then in Subsection 6.2, we apply Theorem 1.1 to demonstrate the computability of the equilibrium state for a Misiurewicz–Thurston rational map and a Hölder continuous potential, thereby establishing Theorem 1.2. Finally, Subsection 6.3 addresses a broader class of expanding Thurston maps where entropy map may lack upper semi-continuity. Here, we use Theorem 1.3 to study the measures of maximal entropy and establish Theorem 1.4.
Acknowledgments. Q. He, Z. Li, and X. Shi were partially supported by Beijing Natural Science Foundation (JQ25001 and 1214021) and National Natural Science Foundation of China (12471083, 12101017, 12090010, and 12090015). I. Binder was partially supported by an NSERC Discovery grant. Q. He was also supported by Peking University Funding (7101303303 and 6201001846).
2. Notation
The chordal metric $\sigma$ on the Riemann sphere $\widehat{\mathbb{C}}$ is defined as follows: $\sigma(z,w)\coloneqq\frac{2\mathopen{\lvert}z-w\mathclose{\rvert}}{\sqrt{1+\mathopen{\lvert}z\mathclose{\rvert}^{2}}\sqrt{1+\mathopen{\lvert}w\mathclose{\rvert}^{2}}}$ for all $z,\,w\in\mathbb{C}$, and $\sigma(\infty,z)=\sigma(z,\infty)\coloneqq\frac{2}{\sqrt{1+\mathopen{\lvert}z\mathclose{\rvert}^{2}}}$ for all $z\in\mathbb{C}$. Let $S^{2}$ denote an oriented topological $2$-sphere. We use $\mathbb{N}$ to denote the set of integers greater than or equal to $1$ and $\mathbb{N}^{*}\coloneqq\bigcup_{k\in\mathbb{N}}\mathbb{N}^{k}$. We write $\mathbb{N}_{0}\coloneqq\{0\}\cup\mathbb{N}$ and $\mathbb{N}_{0}^{*}\coloneqq\{0\}\cup\mathbb{N}^{*}$. We denote by $\mathbb{Q}^{+}$ (resp. $\mathbb{R}^{+}$) the set of all positive rational (resp. real) numbers. The symbol $\log$ denotes the natural logarithm. For $x\in$, we define $\lfloor x\rfloor$ as the greatest integer $\leqslant x$, $\lceil x\rceil$ as the smallest integer $\geqslant x$, and $x^{+}=(x)^{+}\coloneqq\max\{x,\,0\}$. For a function $f\colon X\rightarrow\mathbb{R}$ on a set $X$, we define $f^{+}=(f)^{+}\colon X\rightarrow\mathbb{R}$ by $f^{+}(x)\coloneqq(f(x))^{+}$ for each $x\in X$. The cardinality of a set $A$ is denoted by $\operatorname{card}A$.
Consider a map $f\colon X\rightarrow X$ on a set $X$. We write $f^{n}$ for the $n$-th iterate of $f$, and $f^{-n}\coloneqq(f^{n})^{-1}$, for each $n\in\mathbb{N}$. We set $f^{0}\coloneqq\operatorname{id}_{X}$, the identity map on $X$. For a real-valued function $\phi\colon X\rightarrow$, we write $S_{n}\phi(x)=S^{f}_{n}\phi(x)\coloneqq\sum_{m=0}^{n-1}\phi(f^{m}(x))$ for $x\in X$ and $n\in\mathbb{N}_{0}$. We omit the superscript $f$ when the map $f$ is clear from the context. When $n=0$, by definition $S_{0}\phi=0$.
Let $(X,d)$ be a metric space. We denote by $\mathcal{B}(X)$ the $\sigma$-algebra of all Borel subsets of $X$. For each subset $Y\subseteq X$, we denote the diameter of $Y$ by $\operatorname{diam}_{d}Y\coloneqq\sup\{d(x,y):x,\,y\in Y\}$, the interior of $Y$ by $\operatorname{int}Y$, and the characteristic function of $Y$ by $\mathbbm{1}_{Y}$.
For each $r\in\mathbb{R}$ and each $x\in X$, we denote the open (resp. closed) ball of radius $r$ centered at $x$ by $B_{d}(x,r)\coloneqq\{y\in X:d(x,y)<r\}$. For each $r\in\mathbb{R}$ and each nonempty set $K\subseteq X$, we define $d(x,K)\coloneqq\inf_{y\in K}d(x,y)$, and $B_{d}(K,r)\coloneqq\{x\in X:d(x,K)<r\}$. We often omit the metric $d$ in the subscript when it is clear from the context.
For a compact metric space $(X,d)$, we denote by $C(X)$ the space of continuous functions from $X$ to , and by $\mathcal{M}(X)$ (resp. $\mathcal{P}(X)$) the set of finite signed Borel measures (resp. Borel probability measures) on $X$. Let $g\colon X\rightarrow X$ be a Borel-measurable transformation. We denote by $\mathcal{M}(X,g)$ the set of $g$-invariant Borel probability measures on $X$. Moreover, for each Borel subset $C\in\mathcal{B}(X)$, $\mathcal{P}(X;C)$ denotes the set $\{\mu\in\mathcal{P}(X):\mu(C)=1\}$. By the Riesz representation theorem we can identify the dual of $C(X)$ with the space $\mathcal{M}(X)$. For $\mu\in\mathcal{M}(X)$, we use $\mathopen{\|}\mu\mathclose{\|}$ to denote the total variation norm of $\mu$, $\operatorname{supp}{\mu}$ to denote the support of $\mu$, and
$\mathopen{\langle}\mu,u\mathclose{\rangle}\coloneqq\int\!u\,\mathrm{d}\mu$
for each $\mu$-integrable Borel function $u$ on $X$. If we do not specify otherwise, we equip $C(X)$ with the uniform norm $\mathopen{\|}\,\cdot\,\mathclose{\|}_{C\mathopen{(}X\mathclose{)}}\coloneqq\mathopen{\|}\,\cdot\,\mathclose{\|}_{\infty}$, and equip $\mathcal{M}(X)$, $\mathcal{P}(X)$, and $\mathcal{M}(X,g)$ with the weak topology.
The space of real-valued Hölder continuous functions with an exponent $\alpha\in(0,1]$ on a compact metric space $(X,d)$ is denoted as $C^{0,\alpha}\mathopen{(}X,d\mathclose{)}$. For each $\phi\in C^{0,\alpha}\mathopen{(}X,d\mathclose{)}$,
$\mathopen{\lvert}\phi\mathclose{\rvert}_{\alpha,d}\coloneqq\sup\{\mathopen{\lvert}\phi(x)-\phi(y)\mathclose{\rvert}/d(x,y)^{\alpha}:x,\,y\in X,\,x\neq y\}.$ (2.1)
For a complete separable metric space $(X,d)$, we recall the Wasserstein–Kantorovich metric $W_{d}$ on $\mathcal{P}(X)$ given by
$W_{d}(\mu,\nu)\coloneqq\sup\bigl\{\mathopen{\lvert}\mathopen{\langle}\mu,f\mathclose{\rangle}-\mathopen{\langle}\nu,f\mathclose{\rangle}\mathclose{\rvert}:f\in C^{0,1}(X,d),\,\mathopen{\lvert}f\mathclose{\rvert}_{1,d}\leqslant 1\bigr\}.$ (2.2)
Note that for Borel probability measures in $\mathcal{P}(X)$, the convergence in $W_{d}$ is equivalent to the convergence in the weak topology (see e.g. [Vi09, Corollary 6.13]).
3. Preliminaries
3.1. Computable analysis
We recall fundamental notions and results from recursion theory and computable analysis. [^4] Our notion of algorithm is consistent with Type-2 machines defined in [We00, Definition 2.1.1]. We present, in order, definitions and results concerning the computability of real numbers, computable structures on metric spaces, computability of open sets, functions, compact sets, and probability measures.
Computability over the reals. We begin by reviewing basic notations and concepts from classical recursion theory; for an introduction, see e.g. [Bri94, Section 3].
Definition 3.1 (Effective enumeration and recursively enumerable set).
Let $S\subseteq\mathbb{N}^{*}$ be a nonempty set. An effective enumeration of $S$ is a sequence $\{x_{i}\}_{i\in\mathbb{N}}$ with $S=\{x_{i}:i\in\mathbb{N}\}$ such that there exists an algorithm that, for each $i\in\mathbb{N}$, upon input $i$, outputs $x_{i}$.
Moreover, a set $I\subseteq\mathbb{N}^{*}$ is said to be a recursively enumerable set [^5] We emphasize that recursively enumerable sets in this article are subsets of $\mathbb{N}^{*}$. if $I=\emptyset$ or there exists an effective enumeration of $I$.
For brevity, the symbol $I$ denotes a nonempty recursively enumerable set throughout this subsection.
Note that $\mathbb{N}^{k}$, for $k\in\mathbb{N}$, and $\mathbb{N}^{*}$ are all recursively enumerable sets by Definition 3.1. We then define partial recursive functions and recursive functions.
Definition 3.2 (Partial recursive and recursive function).
Let $\{i_{n}\}_{n\in\mathbb{N}}$ be an effective enumeration of $I$. We say that $f\colon I\rightarrow\mathbb{N}_{0}^{*}$ is partial recursive if there exists an algorithm that, for each $n\in\mathbb{N}$, on input $n$, outputs $f(i_{n})$ if $f(i_{n})\in\mathbb{N}^{*}$, and runs forever otherwise, namely, if $f(i_{n})=0$. We say that $f\colon I\rightarrow\mathbb{N}_{0}^{*}$ is recursive if $f$ is a partial recursive function with $f(I)\subseteq\mathbb{N}^{*}$.
We now define the computability of real numbers.
Definition 3.3 (Computable real number).
A real number $x$ is called computable if there exist three recursive functions $f\colon\mathbb{N}\rightarrow\mathbb{N}$, $g\colon\mathbb{N}\rightarrow\mathbb{N}$, and $h\colon\mathbb{N}\rightarrow\mathbb{N}$ such that $\bigl\lvert(-1)^{h(n)}f(n)/g(n)-x\bigr\rvert<2^{-n}$ for all $i\in I$ and $n\in\mathbb{N}$.
Let $\{x_{i}\}_{i\in I}$ be a sequence of real numbers. We say that $\{x_{i}\}_{i\in I}$ is a sequence of uniformly computable real numbers if there exist three recursive functions $f\colon\mathbb{N}\times I\rightarrow\mathbb{N}$, $g\colon\mathbb{N}\times I\rightarrow\mathbb{N}$, and $h\colon\mathbb{N}\times I\rightarrow\mathbb{N}$ such that $\bigl\lvert(-1)^{h(n,i)}f(n,i)/g(n,i)-x_{i}\bigr\rvert<2^{-n}$ for all $i\in I$ and $n\in\mathbb{N}$.
Clearly, $x\in\mathbb{R}$ is computable if and only if $\{x_{i}\}_{i\in\mathbb{N}}$ defined by $x_{i}\coloneqq x$ for all $i\in\mathbb{N}$ is uniformly computable. For analogous concepts in the sequel, we will define the uniform sequence version and regard the individual case as the special case of constant sequences.
Computable metric spaces.
Definition 3.4 (Computable metric space).
A computable metric space is a triple $(X,\,\rho,\,\mathcal{S})$ satisfying that
  1. (i)
    $(X,\rho)$ is a separable metric space,
  2. (ii)
    $\mathcal{S}=\{s_{n}\}_{n\in\mathbb{N}}$ forms a countable dense subset $\{s_{n}:n\in\mathbb{N}\}$ of $X$, and
  3. (iii)
    $\{\rho(s_{m},s_{n})\}_{(m,n)\in\mathbb{N}^{2}}$ is a sequence of uniformly computable real numbers.
The points in $\mathcal{S}$ are called ideal. Since $\mathbb{N}^{3}$ is recursively enumerable, the collection $\mathcal{B}\coloneqq\{B(s_{i},m/n):i,\,m,\,n\in\mathbb{N}\}$ can be enumerated as $\{B_{l}\}_{l\in\mathbb{N}}$ satisfying the following: there exists an algorithm that, for each $l\in\mathbb{N}$, upon input $l$, outputs $i,\,m,\,n\in\mathbb{N}$ with $B_{l}=B(s_{i},m/n)$. We call the elements in $\mathcal{B}$ ideal balls and such an enumeration of $\mathcal{B}$ an effective enumeration of ideal balls in $(X,\,\rho,\,\mathcal{S})$.
We then define the computability of points in a computable metric space.
Definition 3.5 (Computable point).
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space with $\mathcal{S}=\{s_{i}\}_{i\in\mathbb{N}}$, and $\{x_{i}\}_{i\in I}$ be a sequence of points in $X$. Then $\{x_{i}\}_{i\in I}$ is called uniformly computable (in $(X,\,\rho,\,\mathcal{S})$) if there exists a recursive function $f\colon\mathbb{N}\times I\rightarrow\mathbb{N}$ such that $\rho\bigl(s_{f(n,i)},x_{i}\bigr)<2^{-n}$ for all $n\in\mathbb{N}$ and $i\in I$. Moreover, a point $x$ in $X$ is computable (in $(X,\,\rho,\,\mathcal{S})$) if $\{x_{i}\}_{i\in\mathbb{N}}$ defined by $x_{i}\coloneqq x$ for all $i\in\mathbb{N}$ is uniformly computable.
We now specify the computable structure on $\mathbb{R}$. Let $\mathcal{S}_{\mathbb{Q}}=\{q_{n}\}_{n\in\mathbb{N}}$ be the enumeration of $\mathbb{Q}$ induced by an effective enumeration of $\mathbb{N}^{3}$ via the mapping $(a,b,c)\mapsto(-1)^{c}a/b$. Note that $\{d_{\mathbb{R}}(q_{m},q_{n})\}_{(m,n)\in\mathbb{N}^{2}}$ is a sequence of uniformly computable real numbers, where $d_{\mathbb{R}}$ is the Euclidean metric. Then the triple $\bigl(\mathbb{R},\,d_{\mathbb{R}},\,\mathcal{S}_{\mathbb{Q}}\bigr)$ forms a computable metric space according to Definition 3.4. A similar construction provides a computable structure for $\mathbb{R}^{+}$. In this article, we fix these as the standard computability structures on $\mathbb{R}$ and $\mathbb{R}^{+}$. It is clear that under these structures, Definitions 3.3 and 3.5 are equivalent for the computability of real numbers. That is, a sequence of reals is uniformly computable in one sense if and only if it is in the other.
We also consider a weaker notion of computability over $\mathbb{R}$ that leverages its natural ordered structure.
Definition 3.6 (Semi-computable real number).
Let $\{x_{i}\}_{i\in I}$ be a sequence of real numbers. We say that $\{x_{i}\}_{i\in I}$ is uniformly lower (resp. upper) semi-computable if there exist three recursive functions $f\colon\mathbb{N}\times I\rightarrow\mathbb{N}$, $g\colon\mathbb{N}\times I\rightarrow\mathbb{N}$, and $h\colon\mathbb{N}\times I\rightarrow\mathbb{N}$ such that for each $i\in I$, $\bigl\{(-1)^{h(n,i)}f(n,i)/g(n,i)\bigr\}_{n\in\mathbb{N}}$ is non-decreasing (resp. non-increasing) and converges to $x_{i}$ as $n\to+\infty$. Moreover, a real number $x$ is called lower (resp. upper) semi-computable if the sequence $\{x_{i}\}_{i\in\mathbb{N}}$ defined by $x_{i}\coloneqq x$ for each $i\in\mathbb{N}$ is uniformly lower (resp. upper) semi-computable.
Lower semi-computable open sets. We define an effective version of open sets and collect some relevant results.
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space. Let $\mathcal{B}$ be the set of ideal balls, and $\{B_{l}\}_{l\in\mathbb{N}}$ be an effective enumeration of ideal balls in $(X,\,\rho,\,\mathcal{S})$. We define the set $\mathcal{B}_{0}\coloneqq\mathcal{B}\cup\{\emptyset\}$ of extended ideal balls and an enumeration $\{D_{l}\}_{l\in\mathbb{N}}$ of $\mathcal{B}_{0}$ such that $D_{1}=\emptyset$ and $D_{l}=B_{l-1}$ for each integer $l\geqslant 2$. We call such an enumeration an effective enumeration of extended ideal balls in $(X,\,\rho,\,\mathcal{S})$.
Definition 3.7 (Lower semi-computable open set).
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space, and $\{D_{l}\}_{l\in\mathbb{N}}$ be an effective enumeration of extended ideal balls. Then a sequence $\{U_{i}\}_{i\in I}$ of open sets in $X$ is said to be uniformly lower semi-computable open (in $(X,\,\rho,\,\mathcal{S})$) if there exists a recursive function $f\colon\mathbb{N}\times I\rightarrow\mathbb{N}$ such that $U_{i}=\bigcup_{n\in\mathbb{N}}D_{f(n,i)}$ for each $i\in I$. Moreover, an open set $U\subseteq X$ is called lower semi-computable open (in $(X,\,\rho,\,\mathcal{S})$) if the sequence $\{U_{i}\}_{i\in\mathbb{N}}$ defined by $U_{i}\coloneqq U$ for $i\in\mathbb{N}$ is uniformly lower semi-computable open.
The above definition of a lower semi-computable open set differs slightly from the ones in [BBRY11, Definition 3.4] and [BRY14, Definition 2.4]. In our definition, we use extended ideal balls, which include the empty set.
The term recursively open set in the literature (e.g. [GHR11, Subsection 2.2 and Definition 2.4] and [HR09, Subsection 3.3]) is equivalent to the notion of lower semi-computable open set defined above. A detailed discussion of this equivalence is provided in [He25].
Note that we can algorithmically decide whether $s\in B$ for each ideal point $s\in\mathcal{S}$ and each extended ideal ball $B\in\mathcal{B}_{0}$. The following result then follows immediately from Definition 3.7.
Proposition 3.8.
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space with $\mathcal{S}=\{s_{n}\}_{n\in\mathbb{N}}$. Assume that $\{U_{i}\}_{i\in I}$ is uniformly lower semi-computable open. Then there exists a recursively enumerable set $E\subseteq\mathbb{N}\times I$ such that $\{s_{n}:(n,i)\in E_{i}\}=\{s_{n}:n\in\mathbb{N}\}\cap U_{i}$, where $E_{i}\coloneqq\{(n,i)\in E:n\in\mathbb{N}\}$ for each $i\in I$.
The following two results are two classical results in computable analysis which both follow immediately from Definitions 3.1 and 3.7 (see e.g. [He25]).
Proposition 3.9.
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space. Assume that $H$ and $L$ are two nonempty recursively enumerable sets with $L\subseteq I\times H$, and that $\{U_{i,h}\}_{(i,h)\in L}$ is uniformly lower semi-computable open. Then $\{\bigcup\{U_{i,h}:(i,h)\in L_{h}\}\}_{h\in H}$ is uniformly lower semi-computable open, where $L_{h}\coloneqq\{(i,h)\in L:i\in I\}$ for each $h\in H$. In particular, if $\{U_{i}\}_{i\in I}$ is uniformly lower semi-computable open, then $\bigcup_{i\in I}U_{i}$ is lower semi-computable open.
Proposition 3.10.
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space. Assume that $\{r_{i}\}_{i\in I}$ is a sequence of uniformly lower semi-computable real numbers and $\{x_{i}\}_{i\in I}$ is uniformly computable in $(X,\,\rho,\,\mathcal{S})$. Then $\{B(x_{i},r_{i})\}_{i\in I}$ is uniformly lower semi-computable open.
Computability of functions. We begin with the definition of oracles for points.
Definition 3.11 (Oracle).
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space with $\mathcal{S}=\{s_{i}\}_{i\in\mathbb{N}}$, and $x\in X$. We say that a function $\tau\colon\mathbb{N}\rightarrow\mathbb{N}$ is an oracle for $x$ if $\rho\bigl(s_{\tau(n)},x\bigr)<2^{-n}$ for each $n\in\mathbb{N}.$
With the above definition, computable functions can be defined as follows.
Definition 3.12 (Computable function).
Let $(X,\,\rho,\,\mathcal{S})$ and $(X^{\prime},\,\rho^{\prime},\,\mathcal{S}^{\prime})$ be computable metric spaces with $\mathcal{S}=\{s_{n}\}_{n\in\mathbb{N}}$ and $\mathcal{S}^{\prime}=\{s^{\prime}_{n}\}_{n\in\mathbb{N}}$. Assume that $\{i_{n}\}_{n\in\mathbb{N}}$ is an effective enumeration of $I$, and $C_{i}\subseteq X$ for each $i\in I$. Then a sequence $\{f_{i}\}_{i\in I}$ of functions $f_{i}\colon X\rightarrow X^{\prime}$ is called a sequence of uniformly computable functions with respect to $\{C_{i}\}_{i\in I}$ if there exists an algorithm that, for all $m,\,n\in\mathbb{N}$, $x\in C_{i_{n}}$, and oracle $\tau$ for $x$, on input l,nl,\,n, and $\tau$, outputs $m\in\mathbb{N}$ with $\rho^{\prime}(s^{\prime}_{m},f_{i_{n}}(x))<2^{-l}$. We often omit the phrase “with respect to $\{C_{i}\}_{i\in I}$” when $C_{i}=X$ for all $i\in I$. Moreover, a function $f\colon X\rightarrow X^{\prime}$ is said to be a computable function on $C$ if $\{f_{i}\}_{i\in\mathbb{N}}$, defined by $f_{i}\coloneqq f$ for all $i\in\mathbb{N}$, is a sequence of uniformly computable functions with respect to $\{C_{i}\}_{i\in\mathbb{N}}$ defined by $C_{i}\coloneqq C$ for all $i\in\mathbb{N}$. We often omit the phrase “with respect to CC” when $C=X$.
Computable functions serve as an effective version of continuous functions. The following result provides examples of computable functions (see e.g. [We00, Examples 4.3.3 and 4.3.13.5]).
Example 3.13.
The exponential function $\exp\colon\mathbb{R}\rightarrow\mathbb{R}$ and the logarithmic function $\log\colon\mathbb{R}^{+}\rightarrow\mathbb{R}$ are computable functions.
We recall the following classical characterization of computable functions (cf. [RY21a, Proposition 5.2.14] and [BBRY11, Proposition 3.6]; see also [He25, Theorem 1.4]).
Proposition 3.14.
Let $(X,\,\rho,\,\mathcal{S})$ and $(X^{\prime},\,\rho^{\prime},\,\mathcal{S}^{\prime})$ be computable metric spaces. Suppose $\{B_{n}^{\prime}\}_{n\in\mathbb{N}}$ is an effective enumeration of ideal balls in $(X^{\prime},\,\rho^{\prime},\,\mathcal{S}^{\prime})$. Given $f_{i}\colon X\rightarrow X^{\prime}$ and $C_{i}\subseteq X$ for each $i\in I$, the following statements are equivalent:
  1. (i)
    The sequence $\{f_{i}\}_{i\in I}$ is a sequence of uniformly computable functions with respect to $\{C_{i}\}_{i\in I}$.
  2. (ii)
    There exists a sequence $\{U_{n,i}\}_{(n,i)\in\mathbb{N}\times I}$ of uniformly lower semi-computable open sets in $(X,\,\rho,\,\mathcal{S})$ such that $f_{i}^{-1}(B_{n}^{\prime})\cap C_{i}=U_{n,i}\cap C_{i}$ for all $i\in I$ and $n\in\mathbb{N}$.
  3. (iii)
    For each nonempty recursively enumerable set $M$ and each sequence $\{V_{m}^{\prime}\}_{m\in M}$ of uniformly lower semi-computable open sets, there exists a sequence $\{W_{m,i}\}_{(m,i)\in M\times I}$ of uniformly lower semi-computable open sets in $(X,\,\rho,\,\mathcal{S})$ such that $f_{i}^{-1}(V_{m}^{\prime})\cap C_{i}=W_{m,i}\cap C_{i}$ for all $m\in M$ and $i\in I$.
We now define a notion of weaker computability property for functions.
Definition 3.15 (Semi-computable function).
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space, and $C_{i}\subseteq X$ for each $i\in I$. A sequence $\{f_{i}\}_{i\in I}$ of functions $f_{i}\colon X\rightarrow\mathbb{R}$ is a sequence of uniformly upper (resp. lower) semi-computable functions with respect to $\{C_{i}\}_{i\in I}$ if there exists a sequence $\{g_{i,n}\}_{(i,n)\in I\times\mathbb{N}}$ of uniformly computable functions $g_{i,n}\colon X\rightarrow\mathbb{R}$ with respect to $\{C_{i}\}_{i\in I}$ such that for each $i\in I$ and each $x\in C_{i}$, $\{g_{i,n}(x)\}_{n\in\mathbb{N}}$ is non-increasing (resp. non-decreasing) and converges to $f_{i}(x)$ as $n\to+\infty$. We often omit the phrase “with respect to $\{C_{i}\}_{i\in I}$” when $C_{i}=X$ for each $i\in I$. Moreover, a function $f\colon X\rightarrow\mathbb{R}$ is said to be an upper (resp. a lower) semi-computable function on $C$ if $\{f_{i}\}_{i\in\mathbb{N}}$ defined by $f_{i}\coloneqq f$ for each $i\in\mathbb{N}$, is a sequence of uniformly upper (resp. lower) semi-computable functions with respect to $\{C_{i}\}_{i\in\mathbb{N}}$ defined by $C_{i}\coloneqq C$ for all $i\in\mathbb{N}$. We often omit the phrase “with respect to CC” when $C=X$.
The following proposition is an immediate consequence of Proposition 3.14 (see e.g. [He25]).
Proposition 3.16.
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space, and $\mathcal{S}_{\mathbb{Q}}=\{q_{n}\}_{n\in\mathbb{N}}$. Given $f_{i}\colon X\rightarrow\mathbb{R}$ and $C_{i}\subseteq X$ for all $i\in I$, the following statements are equivalent:
  1. (i)
    The sequence $\{f_{i}\}_{i\in I}$ is a sequence of uniformly upper (resp. lower) semi-computable functions with respect to $\{C_{i}\}_{i\in I}$.
  2. (ii)
    There exists a sequence $\{U_{n,i}\}_{(n,i)\in\mathbb{N}\times I}$ of uniformly lower semi-computable open sets in $(X,\,\rho,\,\mathcal{S})$ such that $f_{i}^{-1}(Q_{n})\cap C_{i}=U_{n,i}\cap C_{i}$ with $Q_{n}\coloneqq(-\infty,q_{n})$ (resp. $Q_{n}\coloneqq(q_{n},+\infty)$) for all $i\in I$ and $n\in\mathbb{N}$.
  3. (iii)
    For each nonempty recursively enumerable set $L$ and each sequence $\{r_{l}\}_{l\in L}$ of uniformly computable real numbers, there exists a sequence $\{W_{l,i}\}_{(l,i)\in L\times I}$ of uniformly lower semi-computable open sets in $(X,\,\rho,\,\mathcal{S})$ such that $f_{i}^{-1}(R_{l})\cap C_{i}=W_{l,i}\cap C_{i}$ with $R_{l}\coloneqq(-\infty,r_{l})$ (resp. $R_{l}\coloneqq(r_{l},+\infty)$) for all $l\in L$ and $i\in I$.
The following result indicates that the characteristic function of a lower semi-computable open set is a lower semi-computable function (see e.g. [He25]).
Proposition 3.17.
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space. Assume that $\{U_{i}\}_{i\in I}$ is uniformly lower semi-computable open. Then there exists a sequence $\{h_{n,i}\}_{(n,i)\in\mathbb{N}\times I}$ of uniformly computable functions $h_{n,i}\colon X\rightarrow\mathbb{R}$ such that for each $i\in I$, the following properties are satisfied:
  1. (i)
    For each $x\in X$, $\{h_{n,i}(x)\}_{n\in\mathbb{N}}$ is non-decreasing and converges to $\mathbbm{1}_{U_{i}}(x)$ as $n\to+\infty$.
  2. (ii)
    For each $n\in\mathbb{N}$, $h_{n,i}(x)\geqslant 0$ for each $x\in X$ and $h_{n,i}(x)=0$ for each $x\notin U_{i}$.
Recursively compact sets and recursively precompact metric spaces. Here we recall the definitions of recursive compactness and recursive precompactness. For a more detailed discussion, see [GHR11, Section 2].
Definition 3.18 (Recursively compact set).
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space with $\mathcal{S}=\{s_{i}\}_{i\in\mathbb{N}}$, and $\{i_{l}\}_{l\in\mathbb{N}}$ be an effective enumeration of $I$. A sequence $\{K_{i}\}_{i\in I}$ of compact sets in $X$ is called uniformly recursively compact (in $(X,\,\rho,\,\mathcal{S})$) if there exists an algorithm that, for each $n\in\mathbb{N}$, each sequence $\{m_{n}\}_{n=1}^{p}$ of integers, and each sequence $\{q_{n}\}_{n=1}^{p}$ of positive rational numbers, upon input, halts if and only if $K_{i_{l}}\subseteq\bigcup_{n=1}^{p}B(s_{m_{n}},q_{n})$. Moreover, a set $K\subseteq X$ is called recursively compact (in $(X,\,\rho,\,\mathcal{S})$) if the sequence $\{K_{i}\}_{i\in\mathbb{N}}$ defined by $K_{i}\coloneqq K$ for each $i\in\mathbb{N}$, is uniformly recursively compact.
Note that for each compact set $K$ and each function $f\colon\mathbb{N}\rightarrow\mathbb{N}$, $K\subseteq\bigcup_{n\in\mathbb{N}}D_{f(n)}$ if and only if $K\subseteq\bigcup_{n=1}^{k}D_{f(n)}$ for some $k\in\mathbb{N}$. This implies the following result.
Proposition 3.19.
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space. Suppose $\{h_{m}\}_{m\in\mathbb{N}}$ (resp. $\{l_{n}\}_{n\in\mathbb{N}}$) is an effective enumeration of a nonempty recursively enumerable set $H$ (resp. $L$). Assume that $\{K_{h}\}_{h\in H}$ is uniformly recursively compact and $\{U_{l}\}_{l\in L}$ is uniformly lower semi-computable open. Then there exists an algorithm that, for all $m,\,n\in\mathbb{N}$, upon input, halts if and only if $K_{h_{m}}\subseteq U_{l_{n}}$.
We collect some fundamental properties of recursively compact sets (cf. [GHR11, Propositions 1 & 3]; see also [He25]).
Proposition 3.20.
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space. Assume that $X$ is recursively compact, and $\{K_{i}\}_{i\in I}$ is uniformly recursively compact. Then the following statements are true:
  1. (i)
    Let $x_{i}\in X$ for each $i\in I$. Then $\{x_{i}\}_{i\in I}$ is uniformly computable if and only if the sequence $\{\{x_{i}\}\}_{i\in I}$ of singletons is uniformly recursively compact.
  2. (ii)
    $\{X\smallsetminus K_{i}\}_{i\in I}$ is uniformly lower semi-computable open.
  3. (iii)
    If $\{U_{i}\}_{i\in I}$ is uniformly lower semi-computable open, then $\{K_{i}\smallsetminus U_{i}\}_{i\in I}$ is uniformly recursively compact.
  4. (iv)
    If $\{f_{i}\}_{i\in I}$ is a sequence of uniformly lower (resp. upper) semi-computable functions $f_{i}\colon X\rightarrow\mathbb{R}$ with respect to $\{K_{i}\}_{i\in I}$, then $\{\inf_{x\in K_{i}}f_{i}(x)\}_{i\in I}$ (resp. $\{\sup_{x\in K_{i}}f_{i}(x)\}_{i\in I}$) is uniformly lower (resp. upper) semi-computable.
  5. (v)
    If $\{T_{i}\}_{i\in I}$ is a sequence of uniformly computable functions $T_{i}\colon X\rightarrow X$ with respect to $\{K_{i}\}_{i\in I}$, then $\{T_{i}(K_{i})\}_{i\in I}$ is uniformly recursively compact.
Next, we investigate whether the property of uniform computability for recursively compact sets is preserved under the union and intersection.
Proposition 3.21.
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space. Suppose $X$ is recursively compact, $H$ and $L$ are two nonempty recursively enumerable sets with $L\subseteq I\times H$, and $\{K_{i,h}\}_{(i,h)\in L}$ is uniformly recursively compact. Define $L_{h}\coloneqq\{(i,h)\in L:i\in I\}$ for each $h\in H$. Then the following statements are true:
  1. (i)
    $\{\bigcap\{K_{i,h}:(i,h)\in L_{h}\}\}_{h\in H}$ is uniformly recursively compact.
  2. (ii)
    If the function $F\colon H\rightarrow\mathbb{N}$ defined by $F(h)\coloneqq\operatorname{card}\mathopen{(}L_{h}\mathclose{)}$ for $h\in H$ is recursive, then $\{\bigcup\{K_{i,h}:(i,h)\in L_{h}\}\}_{h\in H}$ is uniformly recursively compact.
Proposition 3.21 (i) follows immediately from Proposition 3.9 and Proposition 3.20 (ii) and (iii). Moreover, Proposition 3.21 (ii) follows from Definition 3.18. As a corollary of Proposition 3.21 (ii), we obtain the following result, which is important in the proof of Theorem 1.4.
Corollary 3.22.
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space. Assume that $X$ is recursively compact, $T\colon X\rightarrow X$ is a computable function, and $\{U_{i}\}_{i\in I}$ is uniformly lower semi-computable open. Then $\{V_{n,i}\}_{(n,i)\in\mathbb{N}\times I}$ is uniformly lower semi-computable open, where $V_{n,i}$ is defined inductively by setting $V_{1,i}\coloneqq U_{i}$ and $V_{n+1,i}\coloneqq T^{-1}(V_{n,i})\cap U_{i}$ for each $n\in\mathbb{N}$ and each $i\in I$.
Proof.
Since $T$ is a computable function, by Definition 3.2, we obtain that $\{T^{n}\}_{n\in\mathbb{N}_{0}}$ is a sequence of uniformly computable functions. Then by Proposition 3.14, $\{T^{-n}(U_{i})\}_{(n,i)\in\mathbb{N}_{0}\times I}$ is a sequence of uniformly lower semi-computable open sets. Hence, since $X$ is recursively compact, by Proposition 3.20 (iii), $\{X\smallsetminus T^{-n}(U_{i})\}_{(n,i)\in\mathbb{N}_{0}\times I}$ is a sequence of uniformly recursively compact sets.
Define $L\subseteq\mathbb{N}_{0}\times\mathbb{N}\times I$ by $L\coloneqq\{(m,n,i)\in\mathbb{N}_{0}\times\mathbb{N}\times I:m<n\}$. Then $L$ is a recursively enumerable set and $\{X\smallsetminus T^{-m}(U_{i})\}_{(m,n,i)\in L}$ is uniformly recursively compact. Note that $\mathrm{card}\{m\in\mathbb{N}_{0}:(m,n,i)\in L\}=n$ for all $n\in\mathbb{N}$ and $i\in I$. Then the function $F\colon\mathbb{N}\times I\rightarrow\mathbb{N}_{0}$ given by $F(n,i)\coloneqq\mathrm{card}\{m\in\mathbb{N}_{0}:(m,n,i)\in L\}$ is a recursive function. Thus, by Proposition 3.21 (ii), we obtain that $\{\bigcup\{X\smallsetminus T^{-m}(U_{i}):m\in\mathbb{N}_{0},\,(m,n,i)\in L\}\}_{(n,i)\in\mathbb{N}\times I}$ is a sequence of uniformly recursively compact sets. Hence, since $X$ is a recursively compact set, by Proposition 3.20 (ii), $\{X\smallsetminus\bigcup\{X\smallsetminus T^{-m}(U_{i}):m\in\mathbb{N}_{0},\,(m,n,i)\in L\}\}_{(n,i)\in\mathbb{N}\times I}=\{\bigcap\{T^{-m}(U_{i}):m\in\mathbb{N}_{0},\,(m,n,i)\in L\}\}_{(n,i)\in\mathbb{N}\times I}$ is uniformly lower semi-computable open. Since $V_{1,i}=U_{i}$ and $V_{n+1,i}=T^{-1}(V_{n,i})\cap U_{i}$ for all $n\in\mathbb{N}$ and $i\in I$, it follows by induction that $V_{n+1,i}=\bigcap_{k=0}^{n}T^{-k}(U_{i})$ for each $n\in\mathbb{N}_{0}$. Therefore, $\{V_{n,i}\}_{(n,i)\in\mathbb{N}\times I}$ is uniformly lower semi-computable open. ∎
Moreover, given the recursive compactness of $X$, the computability of functions is preserved under a finite number of operations among additions and multiplications. We summarize this property in the following result (cf. [We00, Corollary 4.3.4]; see also [He25]).
Proposition 3.23.
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space in which $X$ is recursively compact, and $H$ be a nonempty recursively enumerable set. Assume that $\{f_{i}\}_{i\in I}$ (resp. $\{g_{h}\}_{h\in H}$) is a sequence of uniformly computable functions $f_{i}\colon X\rightarrow\mathbb{R}$ (resp. $g_{h}\colon X\rightarrow\mathbb{R}$). Then $\{f_{i}+g_{h}\}_{(i,h)\in I\times H}$, $\{f_{i}\cdot g_{h}\}_{(i,h)\in I\times H}$ are two sequences of uniformly computable functions.
Next, we recall the definition of recursively precompact metric space.
Definition 3.24 (Recursively precompact metric space).
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space with $\mathcal{S}=\{s_{i}\}_{i\in\mathbb{N}}$. Then $(X,\,\rho,\,\mathcal{S})$ is called recursively precompact if there exists an algorithm that, for each $n\in\mathbb{N}$, on input $n$, outputs a finite subset $\{r_{i}:1\leqslant i\leqslant m\}$ of $\mathbb{N}$ such that $X=\bigcup_{i=1}^{m}B(s_{r_{i}},2^{-n})$.
Finally, we record the following useful characterization of complete recursively precompact metric spaces (cf. [GHR11, Proposition 4]; see also [He25]).
Proposition 3.25.
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space. Then $X$ is recursively compact if and only if $(X,\rho)$ is complete and $(X,\,\rho,\,\mathcal{S})$ is recursively precompact.
Computability of probability measures. Building upon the theory of computable functions and recursively compact sets, we now discuss the computability of probability measures. We begin by reviewing the computable structure on the measure space $\mathcal{P}(X)$ introduced in [HR09, Section 4] (cf. [HR09, Proposition 4.1.3]; see also [He25]).
Proposition 3.26.
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space with $\mathcal{S}=\{s_{n}\}_{n\in\mathbb{N}}$. Assume that $X$ is recursively compact in $(X,\,\rho,\,\mathcal{S})$. Then the following statements are true:
  1. (i)
    There exists an enumeration $\mathcal{Q}_{\mathcal{S}}=\{\nu_{k}\}_{k\in\mathbb{N}}$ of the set of Borel probability measures that are supported on finitely many points in $\{s_{n}:n\in\mathbb{N}\}$ and assign rational values to them such that there exists an algorithm that, for each $k\in\mathbb{N}$, upon input $k$, outputs a sequence $\{n_{l}\}_{l=1}^{p}$ of integers and a sequence $\{q_{l}\}_{l=1}^{p}$ of positive rational numbers satisfying that $\sum_{l=1}^{p}q_{l}=1$ and $\nu_{k}=\sum_{l=1}^{p}q_{l}\delta_{s_{n_{l}}}$.
  2. (ii)
    $(\mathcal{P}(X),\,W_{\rho},\,\mathcal{Q}_{\mathcal{S}})$ is also a computable metric space in which $\mathcal{P}(X)$ is recursively compact, where $W_{\rho}$ is the Wasserstein–Kantorovich metric on $\mathcal{P}(X)$ (see (2.2)).
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space and assume that $X$ is recursively compact. We endow the measure space $\mathcal{P}(X)$ with the computable structure $(\mathcal{P}(X),\,W_{\rho},\,\mathcal{Q}_{\mathcal{S}})$ given by Proposition 3.26.
The computability of measures is then defined via Definition 3.5. Specifically, a sequence $\{\mu_{i}\}_{i\in I}$ of measures in $\mathcal{P}(X)$ is a sequence of uniformly computable measures if it is uniformly computable in $(\mathcal{P}(X),\,W_{\rho},\,\mathcal{Q}_{\mathcal{S}})$, and a single measure $\mu\in\mathcal{P}(X)$ is a computable measure if the corresponding constant sequence consisting of $\mu$ is uniformly computable.
We now recall a key result on the computability of the integration function (cf. [HR09, Corollary 4.3.2]; see also [He25]).
Proposition 3.27.
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space. Assume that $X$ is recursively compact in $(X,\,\rho,\,\mathcal{S})$, and that $\{f_{i}\}_{i\in I}$ is a sequence of uniformly computable functions $f_{i}\colon X\rightarrow\mathbb{R}$. Then the sequence $\{\mathcal{I}_{i}\}_{i\in I}$ of functions $\mathcal{I}_{i}\colon\mathcal{P}(X)\rightarrow\mathbb{R}$ defined by $\mathcal{I}_{i}(\mu)\coloneqq\mathopen{\langle}\mu,f_{i}\mathclose{\rangle}$ for $\mu\in\mathcal{P}(X)$ is a sequence of uniformly computable functions.
As immediate corollaries of Proposition 3.27, we have the following results.
Corollary 3.28.
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space. Assume that $X$ is recursively compact in $(X,\,\rho,\,\mathcal{S})$, and that $\{f_{i}\}_{i\in I}$ is a sequence of uniformly upper (resp. lower) semi-computable functions $f_{i}\colon X\rightarrow\mathbb{R}$. Then the sequence $\{\mathcal{I}_{i}\}_{i\in I}$ of functions $\mathcal{I}_{i}\colon\mathcal{P}(X)\rightarrow\mathbb{R}$ given by $\mathcal{I}_{i}(\mu)\coloneqq\mathopen{\langle}\mu,f_{i}\mathclose{\rangle}$ is a sequence of uniformly upper (resp. lower) semi-computable functions.
Corollary 3.29.
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space in which $X$ is recursively compact. Assume that $H$ and $L$ are two nonempty recursively enumerable sets with $L\subseteq I\times H$, $\{U_{i,h}\}_{(i,h)\in L}$ is uniformly lower semi-computable open in $(X,\,\rho,\,\mathcal{S})$, and $\{r_{i,h}\}_{(i,h)\in L}$ is a sequence of uniformly computable real numbers. Then $\{\mathcal{K}_{i}\}_{i\in I}$ is uniformly recursively compact in $(\mathcal{P}(X),\,W_{\rho},\,\mathcal{Q}_{\mathcal{S}})$. Here $\mathcal{K}_{i}\coloneqq\{\mu\in\mathcal{P}(X):\mu(U_{i,h})\leqslant r_{i,h}\text{ for each }(i,h)\in L_{i}\}$, and $L_{i}\coloneqq\{(i,h)\in L:h\in H\}$ for each $i\in I$.
Recall the definition of $\mathcal{M}(X,T;Y)$ from (1.1). The following result indicates the recursive compactness of the set $\mathcal{M}(X,T;Y)$ (cf. [BHLZ25, Lemma 4.12]; see also [He25]).
Proposition 3.30.
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space in which $X$ is recursively compact, and $\{U_{i}\}_{i\in I}$ is a sequence of uniformly lower semi-computable open sets. Assume that $\{T_{i}\}_{i\in I}$ is a sequence of uniformly computable functions $T_{i}\colon X\rightarrow X$ with respect to $\{U_{i}\}_{i\in I}$. Then $\{\mathcal{M}(X,T_{i};U_{i})\}_{i\in I}$ is uniformly recursively compact in $(\mathcal{P}(X),\,W_{\rho},\,\mathcal{Q}_{\mathcal{S}})$. In particular, if $\{T_{i}\}_{i\in I}$ is a sequence of uniformly computable functions, then $\{\mathcal{M}(X,T_{i})\}_{i\in I}$ is uniformly recursively compact.
The following result is useful in the proof of the main result of this article (see e.g. [He25]).
Proposition 3.31.
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space, and $X$ be a recursively compact set in $(X,\,\rho,\,\mathcal{S})$. Then there exists a sequence $\{\tau_{n}\}_{n\in\mathbb{N}}$ of uniformly computable functions $\tau_{n}\colon X\rightarrow\mathbb{R}$ such that $\{\tau_{n}:n\in\mathbb{N}\}$ is dense in $C(X)$. Moreover, for all $\mu,\,\nu\in\mathcal{M}(X)$, $\mu(A)\geqslant\nu(A)$ for each $A\in\mathcal{B}(X)$ if and only if $\bigl\langle\mu,\tau_{n}^{+}\bigr\rangle\geqslant\bigl\langle\nu,\tau_{n}^{+}\bigr\rangle$ for each $n\in\mathbb{N}$.
3.2. Thermodynamic formalism
We review basic concepts from ergodic theory. For more detailed discussions, we refer the reader to [Wa82, Section 4].
Let $(X,\mathcal{B},\mu)$ be a probability space. A partition $\xi=\{A_{h}:h\in H\}$ of $(X,\mathcal{B},\mu)$ is a disjoint collection of elements of $\mathcal{B}$ whose union is $X$, where $H$ is a countable index set. For each pair of partitions $\xi=\{A_{h}:h\in H\}$ and $\eta=\{B_{l}:l\in L\}$ of $X$, their join is the partition $\xi\vee\eta\coloneqq\{A_{h}\cap B_{l}:h\in H,\,l\in L\}$.
Assume that $T\colon X\rightarrow X$ is a measure-preserving transformation of $(X,\mathcal{B},\mu)$. Consider a partition $\xi=\{A_{h}:h\in H\}$ of $X$. For each $n\in\mathbb{N}$, $T^{-n}(\xi)$ denotes the partition $\bigl\{T^{-1}(A_{h}):h\in H\bigr\}$, and $\xi^{n}_{T}$ denotes the join $\xi\vee T^{-1}(\xi)\vee\cdots\vee T^{-(n-1)}(\xi)$. The entropy of $\xi$ is $H_{\mu}(\xi)\coloneqq-\sum_{h\in H}\mu(A_{h})\log(\mu(A_{h}))\in[0,+\infty]$, where $0\log 0$ is defined to be zero. One can show (see e.g. [Wa82, Chapter 4]) that if $H_{\mu}(\xi)<+\infty$, then $\lim\limits_{n\to+\infty}H_{\mu}(\xi^{n}_{T})/n$ exists. We denote this limit by $h_{\mu}(T,\xi)$ and call it the measure-theoretic entropy of $T$ relative to $\xi$. The measure-theoretic entropy of $T$ for $\mu$ is defined as
$h_{\mu}(T)\coloneqq\sup\{h_{\mu}(T,\xi):\xi\text{ is a partition of }X\text{ with }H_{\mu}(\xi)<+\infty\}.$ (3.1)
We now introduce thermodynamic formalism, a particular branch of ergodic theory. The main objects of study are the topological pressure and equilibrium states (see e.g. [PU10, Wa82]; for the general Borel-measurable setting used in Approach II, see e.g. [IT10, Definition 1.1], [DeT17, Section 2.3], and [DoT23, Chapter 1.4]).
Let $(X,\rho)$ be a compact metric space, $T\colon X\rightarrow X$ be a Borel-measurable transformation such that $\mathcal{M}(X,T)\neq\emptyset$, and $\phi\colon X\rightarrow[-\infty,+\infty]$ be a Borel function. Then the topological pressure of the potential $\phi$ with respect to the transformation $T$ is given by
$P(T,\phi)\coloneqq\sup\bigl\{h_{\mu}(T)+\mathopen{\langle}\mu,\phi\mathclose{\rangle}:\mu\in\mathcal{M}(X,T)\text{ and }\mathopen{\langle}\mu,\phi\mathclose{\rangle}>-\infty\bigr\}.$ (3.2)
A measure $\mu\in\mathcal{M}(X,T)$ that attains the supremum in (3.2) is called an equilibrium state for the transformation $T$ and the potential $\phi$. Denote the set of all such measures by $\mathcal{E}(T,\phi)$. In particular, when the potential $\phi$ is the constant function 0, we denote $h_{\mathrm{top}}(T)\coloneqq P(T,0)$ and say that a measure $\mu\in\mathcal{M}(X,T)$ is a measure of maximal entropy of $T$ if $\mu\in\mathcal{E}(T,0)$.
4. Approach I
In this section, we establish the computability of equilibrium states for certain dynamical systems with upper semi-continuous measure-theoretic entropy functions. We begin by recalling several definitions and results from functional analysis: a characterization of the set ${C(X)}^{*}_{\phi,P_{T}}$ in (4.1) and its relation to the set $\mathcal{E}(T,\phi)$ of equilibrium states in (4.2). We then apply these results to verify the recursive compactness of ${C(X)}^{*}_{\phi,P_{T}}$, thereby completing the proof of Theorem 1.1.
Let $(X,\rho)$ be a compact metric space, and $T\colon X\rightarrow X$ a continuous map with finite topological entropy. The measure-theoretic entropy function of $T$ is the function $\mu\mapsto h_{\mu}(T)$ (defined in (3.1)) on the space $\mathcal{M}(X,T)$ of $T$-invariant Borel probability measures, where $\mathcal{M}(X,T)$ is equipped with the weak topology. The topological pressure function of $T$, denoted by $P_{T}$, is the function $\phi\mapsto P(T,\phi)$ (defined in (3.2)) on $C(X)$. For each $\mu\in\mathcal{M}(X,T)$, we denote by $\overline{h}_{\mu}(T)$ the upper semi-continuous regularization of the measure-theoretic entropy function. More precisely, $\overline{h}_{\mu}(T)$ is the supremum of all limit suprema $\limsup_{n\to\infty}h_{\mu_{n}}(T)$, where $\{\mu_{n}\}_{n\in\mathbb{N}}$ ranges over all sequences in $\mathcal{M}(X,T)$ that converge to $\mu$ in the weak topology.
Definition 4.1.
Let $V$ be a real topological vector space, and $G\colon V\rightarrow\mathbb{R}$ be a convex continuous function. A continuous linear functional $F\colon V\rightarrow\mathbb{R}$ is tangent to $G$ at $x\in V$ if
$F(y)\leqslant G(x+y)-G(x)$
for each $y\in V$. We denote the set of all such functionals by $V_{x,G}^{*}$.
The following lemma is well-known (see e.g. [Wa82, Theorem 9.7 (iv) and (v)]).
Lemma 4.2.
Let $(X,\rho)$ be a compact metric space and $T\colon X\rightarrow X$ be a continuous map with finite topological entropy. Then the topological pressure function $P_{T}\colon C(X)\rightarrow\mathbb{R}$ is convex and continuous.
Lemma 4.2 ensures that the set $C(X)_{\phi,P_{T}}^{*}$ is well-defined for any continuous map $T$ on a compact metric space $X$ and any continuous function $\phi$. We now record a characterization of this space from [Wa92, Theorem 3 (i)].
Lemma 4.3.
Let $(X,\rho)$ be a compact metric space, $T\colon X\rightarrow X$ be a continuous map with finite topological entropy, and $\phi\colon X\rightarrow\mathbb{R}$ be a continuous function. A functional $F\in C(X)^{*}$ belongs to $C(X)_{\phi,P_{T}}^{*}$ if and only if there exists a measure $\mu_{F}\in\mathcal{M}(X,T)$ such that $F(f)=\mathopen{\langle}\mu_{F},f\mathclose{\rangle}$ for each $f\in C(X)$, and $\mu_{F}$ is the weak limit of a sequence of measures $\{\mu_{n}\}_{n\in\mathbb{N}}$ in $\mathcal{M}(X,T)$ satisfying $h_{\mu_{n}}(T)+\mathopen{\langle}\mu_{n},\phi\mathclose{\rangle}\to P(T,\phi)$ as $n\to+\infty$.
Remark 4.4.
Recall that $C(X)^{*}$ can be naturally identified with the space $\mathcal{M}(X)$ of finite signed Borel measures on $X$. Then Lemma 4.3 allows us to identify $C(X)_{\phi,P_{T}}^{*}$ as a subset of $\mathcal{M}(X,T)$. We will adopt this identification in the remaining part of this article.
The following proposition characterizes $C(X)_{\phi,P_{T}}^{*}$ and its relation to the set of equilibrium states $\mathcal{E}(T,\phi)$.
Proposition 4.5.
Let $(X,\rho)$ be a compact metric space, $T\colon X\rightarrow X$ be a continuous map with finite topological entropy, and $\phi\colon X\rightarrow\mathbb{R}$ be a continuous function. Then
$C(X)_{\phi,P_{T}}^{*}$ $=\bigl\{\mu\in\mathcal{M}(X,T):\overline{h}_{\mu}(T)+\mathopen{\langle}\mu,\phi\mathclose{\rangle}=P(T,\phi)\bigr\}\text{ and }$ (4.1)
$\mathcal{E}(T,\phi)$ $=C(X)_{\phi,P_{T}}^{*}\cap\bigl\{\mu\in\mathcal{M}(X,T):\overline{h}_{\mu}(T)\leqslant h_{\mu}(T)\bigr\}.$ (4.2)
The characterization (4.1) follows from Lemma 4.3, and (4.2) is a consequence of [Wa92, Theorem 5].
The final ingredient of our argument is the following variational characterization of the regularized entropy, which can be proved by the same method as [Wa82, Theorem 9.12].
Lemma 4.6.
Let $(X,\rho)$ be a compact metric space and $T\colon X\rightarrow X$ be a continuous map with finite topological entropy. Then $\overline{h}_{\mu}(T)=\inf\{P(T,\theta)-\mathopen{\langle}\mu,\theta\mathclose{\rangle}:\theta\in C(X)\}$ for each $\mu\in\mathcal{M}(X,T)$.
We now apply these results to establish the computability of $C(X)_{\phi,P_{T}}^{*}$.
Theorem 4.7.
Let $(X,\,\rho,\,\mathcal{S},\,\{X_{n}\}_{n\in\mathbb{N}},\,\{T_{n}\}_{n\in\mathbb{N}},\,\{\phi_{n}\}_{n\in\mathbb{N}})$ be a uniformly computable system with $X_{n}=X$ for all $n\in\mathbb{N}$. Suppose $T_{n}$ has finite topological pressure for each $n\in\mathbb{N}$. Assume that the sequence $\{T_{n}\}_{n\in\mathbb{N}}$ of transformations and the sequence $\{\phi_{n}\}_{n\in\mathbb{N}}$ of functions satisfy properties (i) and (ii) in Theorem 1.1. For each $n\in\mathbb{N}$, define
$\Gamma_{n}\coloneqq\bigl\{\mu\in\mathcal{M}(X,T_{n}):\overline{h}_{\mu}(T_{n})+\mathopen{\langle}\mu,\phi_{n}\mathclose{\rangle}=P(T_{n},\phi_{n})\bigr\}.$ (4.3)
Then $\{\Gamma_{n}\}_{n\in\mathbb{N}}$ is uniformly recursively compact in $(\mathcal{P}(X),\,W_{\rho},\,\mathcal{Q}_{\mathcal{S}})$.
Proof.
For each $n\in\mathbb{N}$, by (4.1) in Proposition 4.5, $\Gamma_{n}$ is indeed the space of functionals which are tangent to $P_{T_{n}}$ at $\phi_{n}$. Recall that the sequence $\{\psi_{n,i}\}_{(n,i)\in\mathbb{N}^{2}}$ of functions and the sequence $\{D_{n}\}_{n\in\mathbb{N}}$ defined by $D_{n}\coloneqq\{\psi_{n,i}:i\in\mathbb{N}\}$ are given in Theorem 1.1 property (i). In what follows, we establish the following characterization of the spaces $\Gamma_{n}$.
Claim 1. For each $n\in\mathbb{N}$, we have $\Gamma_{n}=\{\mu\in\mathcal{M}(X,T_{n}):\inf\{P(T_{n},\psi)-\mathopen{\langle}\mu,\psi\mathclose{\rangle}:\psi\in D_{n}\}\geqslant P(T_{n},\phi_{n})-\mathopen{\langle}\mu,\phi_{n}\mathclose{\rangle}\}$.
Proof of Claim 1. Fix an arbitrary $n\in\mathbb{N}$. By (4.3) and Lemma 4.6, we have that
$\Gamma_{n}=\{\mu\in\mathcal{M}(X,T_{n}):\inf\{P(T_{n},\theta)-\mathopen{\langle}\mu,\theta\mathclose{\rangle}:\theta\in C(X)\}=P(T_{n},\phi_{n})-\mathopen{\langle}\mu,\phi_{n}\mathclose{\rangle}\}.$
Hence, to show Claim 1, it suffices to show that for each $\mu\in\mathcal{M}(X,T_{n})$, the following two relations are equivalent:
$\inf\{P(T_{n},\theta)-\mathopen{\langle}\mu,\theta\mathclose{\rangle}:\theta\in C(X)\}$ $=P(T_{n},\phi_{n})-\mathopen{\langle}\mu,\phi_{n}\mathclose{\rangle},$ (4.4)
$\inf\{P(T_{n},\psi)-\mathopen{\langle}\mu,\psi\mathclose{\rangle}:\psi\in D_{n}\}$ $\geqslant P(T_{n},\phi_{n})-\mathopen{\langle}\mu,\phi_{n}\mathclose{\rangle}.$ (4.5)
First, assume $\mu\in\mathcal{M}(X,T_{n})$ satisfies (4.4). It follows from $D_{n}\subseteq C(X)$ that (4.4) implies (4.5). Conversely, assume $\mu\in\mathcal{M}(X,T_{n})$ satisfies (4.5). We proceed by contradiction. Suppose $\mu$ does not satisfy (4.4). Then there exists $\phi\in C(X)$ such that $P(T_{n},\phi)-\mathopen{\langle}\mu,\phi\mathclose{\rangle}<P(T_{n},\phi_{n})-\mathopen{\langle}\mu,\phi_{n}\mathclose{\rangle}$. Since $\overline{D}_{n}$ (the closure of $D_{n}\coloneqq\{\psi_{n,i}:i\in\mathbb{N}\}$) contains a neighborhood of $\phi_{n}$ by property (i) in Theorem 1.1, there exists $c\in(0,1]$ such that $c\phi+(1-c)\phi_{n}\in\overline{D}_{n}$. Thus, by the convexity and continuity of the pressure function $\phi\mapsto P(T_{n},\phi)$, we obtain that
$P(T_{n},\phi_{n})-\mathopen{\langle}\mu,\phi_{n}\mathclose{\rangle}$ $>P\bigl(T_{n},c\phi+(1-c)\phi_{n}\bigr)-\mathopen{\big\langle}\mu,c\phi+(1-c)\phi_{n}\mathclose{\big\rangle}$
$\geqslant\inf\{P(T_{n},\psi)-\mathopen{\langle}\mu,\psi\mathclose{\rangle}:\psi\in\overline{D}_{n}\}=\inf\{P(T_{n},\psi)-\mathopen{\langle}\mu,\psi\mathclose{\rangle}:\psi\in D_{n}\},$
which contradicts the assumption that $\mu$ satisfies (4.5). Hence, $\mu$ must satisfy (4.4), completing the proof of Claim 1.
For each $n\in\mathbb{N}$, define a function $f_{n}\colon\mathcal{P}(X)\rightarrow\mathbb{R}$ by
$f_{n}(\nu)\coloneqq\inf\{P(T_{n},\psi_{n,i})-\mathopen{\langle}\nu,\psi_{n,i}\mathclose{\rangle}:i\in\mathbb{N}\}+\mathopen{\langle}\nu,\phi_{n}\mathclose{\rangle}\quad\text{for }\nu\in\mathcal{P}(X)\text{ and }n\in\mathbb{N}.$ (4.6)
Claim 2. $\{f_{n}\}_{n\in\mathbb{N}}$ is a sequence of uniformly upper semi-computable functions.
Proof of Claim 2. We construct a sequence $\{F_{m,n}\}_{(m,n)\in\mathbb{N}^{2}}$ of uniformly computable functions such that for each $n\in\mathbb{N}$, the sequence $\{F_{m,n}\}_{m\in\mathbb{N}}$ is non-increasing and converges pointwise to $f_{n}$ as $m\to\infty$. By property (i) of Theorem 1.1, $\{P(T_{n},\psi_{n,i})\}_{(n,i)\in\mathbb{N}^{2}}$ is a sequence of uniformly upper semi-computable real numbers. Hence, by Definition 3.6, there exists an algorithm $\mathcal{A}_{p}$ such that for all $m,\,n,\,i\in\mathbb{N}$, on input m,n,im,\,n,\,i, the algorithm $\mathcal{A}_{p}$ outputs $p_{m,n,i}\in\mathcal{Q}$ such that $\{p_{m,n,i}\}_{(m,n,i)\in\mathbb{N}^{3}}$ is non-increasing in $m$ and converges to $P(T_{n},\psi_{n,i})$ as $m\rightarrow+\infty$ for all $n,i\in\mathbb{N}$. Define integral functionals $\mathcal{I}_{n,i}(\nu)\coloneqq\mathopen{\langle}\nu,\phi_{n}-\psi_{n,i}\mathclose{\rangle}$ for all $n,i\in\mathbb{N}$ and $\nu\in\mathcal{P}(X)$. Then by Proposition 3.27, it follows from the uniform computability of $\{\psi_{n,i}\}_{(n,i)\in\mathbb{N}^{2}}$ and $\{\phi_{n}\}_{n\in\mathbb{N}}$ that $\{\mathcal{I}_{n,i}\}_{(n,i)\in\mathbb{N}^{2}}$ is a sequence of uniformly computable functions. Define $F_{m,n}(\nu)\coloneqq\min\{p_{m,n,i}+\mathcal{I}_{n,i}(\nu):i\in\mathbb{N}\cap[1,m]\}$ for all $n,m\in\mathbb{N}$ and $\nu\in\mathcal{P}(X)$. By the uniform computability of $\{\mathcal{I}_{n,i}\}_{(n,i)\in\mathbb{N}^{2}}$, there exists an algorithm $\mathcal{A}_{\mathcal{I}}$ such that for all $n,\,i,\,k\in\mathbb{N}$ and $\nu\in\mathcal{P}(X)$, on input n,i,kn,\,i,\,k and an oracle for $\nu$, $\mathcal{A}_{\mathcal{I}}$ outputs $q_{i,k}$ with $\mathopen{\lvert}q_{i,k}-\mathcal{I}_{n,i}(\nu)\mathclose{\rvert}\leqslant 2^{-k}$. We can now design an algorithm $\mathcal{A}_{F}$ to compute $\{F_{m,n}\}_{(m,n)\in\mathbb{N}^{2}}$. For each $m,\,n,\,k\in\mathbb{N}$ and each $\nu\in\mathcal{P}(X)$, we apply $\mathcal{A}_{p}$ and $\mathcal{A}_{\mathcal{I}}$ to compute $\{p_{m,n,i}\}_{i=1}^{m}$ and $\{q_{i,k}\}_{i=1}^{m}$ such that $\mathopen{\lvert}q_{i,k}-\mathcal{I}_{n,i}(\nu)\mathclose{\rvert}\leqslant 2^{-k}$ for each integer $1\leqslant i\leqslant m$. Then we compute $\min\{p_{m,n,i}+q_{i,k}:i\in\mathbb{N}\cap[1,m]\}$ as the output of $\mathcal{A}_{F}$. Note that $\min\{p_{m,n,i}+q_{i,k}:i\in\mathbb{N}\cap[1,m]\}+2^{-k}\geqslant F_{m,n}(\nu)\geqslant\min\{p_{m,n,i}+q_{i,k}:i\in\mathbb{N}\cap[1,m]\}-2^{-k}$. It follows that $\mathcal{A}_{F}$ demonstrates the uniform computability of the sequence $\{F_{m,n}\}_{(m,n)\in\mathbb{N}^{2}}$ of functions.
Now fix an integer $n\in\mathbb{N}$ and a measure $\nu\in\mathcal{P}(X)$. Since $\{p_{m,n,i}\}_{(m,i)\in\mathbb{N}^{2}}$ is non-increasing in $m$ for each $i\in\mathbb{N}$, the sequence $\{F_{m,n}(\nu)\}_{m\in\mathbb{N}}$ is non-increasing; so its limit exists. Note that $p_{m,n,i}\geqslant P(T_{n},\psi_{n,i})$ for all $m,\,i\in\mathbb{N}$. Then by construction, we have $F_{m,n}(\nu)\geqslant f_{n}(\nu)$ for each $m\in\mathbb{N}$. Hence, $\lim_{m\to+\infty}F_{m,n}(\nu)\geqslant f_{n}(\nu)$. On the other hand, since $\lim_{m\to+\infty}p_{m,n,i}=P(T_{n},\psi_{n,i})$ for each $i\in\mathbb{N}$, it follows from (4.6) that for each $\epsilon>0$, there exist integers $i,j$ with $1\leqslant i\leqslant j$ such that $f_{n}(\nu)+\epsilon>P(T_{n},\psi_{n,i})+\mathcal{I}_{n,i}(\nu)+\epsilon/2>p_{j,n,i}+\mathcal{I}_{n,i}(\nu)\geqslant F_{j,n}(\nu)\geqslant\lim_{m\to+\infty}F_{m,n}(\nu)$. Consequently, $f_{n}(\nu)\geqslant\lim_{m\to+\infty}F_{m,n}(\nu)$. This shows that $\lim_{m\to+\infty}F_{m,n}(\nu)=f_{n}(\nu)$.
Since $\{F_{m,n}\}_{(m,n)\in\mathbb{N}^{2}}$ is a sequence of uniformly computable functions and $\{F_{m,n}(\nu)\}_{m\in\mathbb{N}}$ is non-increasing for all $\nu$, by Definition 3.15, $\{f_{n}\}_{n\in\mathbb{N}}$ is a sequence of uniformly upper semi-computable functions. This establishes Claim 2.
We now complete the proof by expressing $\Gamma_{n}$ as the complement of a uniformly lower semi-computable open set within the uniformly recursively compact set $\mathcal{M}(X,T_{n})$. By property (ii) of Theorem 1.1, $\{P(T_{n},\phi_{n})\}_{n\in\mathbb{N}}$ is a sequence of uniformly lower semi-computable real numbers. By Definition 3.6, there exists a sequence $\{q_{m,n}\}_{(m,n)\in\mathbb{N}^{2}}$ of uniformly computable real numbers such that, for each $n\in\mathbb{N}$, the sequence $\{q_{m,n}\}_{m\in\mathbb{N}}$ is non-decreasing in $m$ and converges to $P(T_{n},\phi_{n})$ as $m\to+\infty$. Hence, it follows from the definition of $\{D_{n}\}_{n\in\mathbb{N}}$, Claim 1, and (4.6) that
$\Gamma_{n}=\mathcal{M}(X,T_{n})\cap f_{n}^{-1}([P(T_{n},\phi_{n}),+\infty))=\mathcal{M}(X,T_{n})\smallsetminus\bigcup_{m\in\mathbb{N}}f_{n}^{-1}((-\infty,q_{m,n})).$ (4.7)
Since $\{q_{m,n}\}_{(m,n)\in\mathbb{N}^{2}}$ is a sequence of uniformly computable real numbers, by Claim 2 and the implication from (i) to (iii) in Proposition 3.16, $\bigl\{f_{n}^{-1}((-\infty,q_{m,n}))\bigr\}_{(m,n)\in\mathbb{N}^{2}}$ is uniformly lower semi-computable open in $(\mathcal{P}(X),\,W_{\rho},\,\mathcal{Q}_{\mathcal{S}})$. Hence, by Proposition 3.9, $\bigl\{\bigcup_{m\in\mathbb{N}}f_{n}^{-1}((-\infty,q_{m,n}))\bigr\}_{n\in\mathbb{N}}$ is uniformly lower semi-computable open. Moreover, since $\{T_{n}\}_{n\in\mathbb{N}}$ is a sequence of uniformly computable functions, by Proposition 3.30, $\{\mathcal{M}(X,T_{n})\}_{n\in\mathbb{N}}$ is uniformly recursively compact. Therefore, by Proposition 3.20 (iii) and (4.7), $\{\Gamma_{n}\}_{n\in\mathbb{N}}$ is uniformly recursively compact. ∎
Now we turn to prove Theorem 1.1.
Proof of Theorem 1.1.
Let $\Gamma_{n}$ be as defined in (4.3). Consider an arbitrary $n\in\mathbb{N}$. Then the measure-theoretic entropy map $\nu\mapsto h_{\nu}(T_{n})$ is upper semi-continuous by hypothesis. This implies that $\overline{h}_{\nu}(T_{n})=h_{\nu}(T_{n})$ for all $\nu\in\mathcal{M}(X,T_{n})$. It follows from (4.3) that the set $\Gamma_{n}$ coincides with the set of equilibrium states $\mathcal{E}(T_{n},\phi_{n})$. Hence, by property (iii), we have $\Gamma_{n}=\mathcal{E}(T_{n},\phi_{n})=\{\mu_{n}\}$.
By Theorem 4.7, the sequence $\{\{\mu_{n}\}\}_{n\in\mathbb{N}}$ of singletons is uniformly recursively compact in $(\mathcal{P}(X),\,W_{\rho},\,\mathcal{Q}_{\mathcal{S}})$. Therefore, by Proposition 3.20 (i), $\{\mu_{n}\}_{n\in\mathbb{N}}$ is uniformly computable. ∎
5. Approach II
This section is dedicated to the proof of Theorem 1.3 and its applications. Subsection 5.1 recalls essential notions and establishes Theorem 5.6, which states the relationship between the set $\mathcal{M}(X,T;Y,J)$ and $\mathcal{E}(T,\phi)$ under additional conditions. The proof of Theorem 1.3 follows in Subsection 5.2. Finally, in Subsection 5.3, we apply these results to demonstrate the computability of equilibrium states.
5.1. Jacobian and the transfer operator
We begin by recalling the definition of Jacobians and establishing their existence, uniqueness, and Rokhlin’s formula. We then define transfer operators and establish their key properties in our setting. Using these tools, we provide an equivalent characterization of Jacobians with respect to invariant measures in Theorem 5.5. Building on this, the subsection concludes with Theorem 5.6.
Definition 5.1 (Jacobian).
Let $(X,\rho)$ be a compact metric space, and $T\colon X\rightarrow X$ be a Borel-measurable transformation. We say that $A\subseteq X$ is admissible (for $T$) if $A,\,T(A)\in\mathcal{B}(X)$, and $T|_{A}$ is injective. Suppose $J\colon X\rightarrow[0,+\infty)$ is a Borel function, $\mu\in\mathcal{P}(X)$, and $E\in\mathcal{B}(X)$ with $\mu(E)=1$. Then $J$ is said to be a Jacobian on $E$ for $T$ with respect to $\mu$ if for all admissible sets $A\subseteq E$,
$\mu(T(A))=\int_{A}\!J\,\mathrm{d}\mu.$ (5.1)
Moreover, we say that $J$ is a Jacobian for $T$ with respect to $\mu$ if there exists $\widetilde{E}\in\mathcal{B}(X)$ with $\mu\bigl(\widetilde{E}\bigr)=1$ such that $J$ is a Jacobian on $\widetilde{E}$ for $T$ with respect to $\mu$.
Recall that $\mathcal{P}(X;Y)=\{\mu\in\mathcal{P}(X):\mu(Y)=1\}$ for $Y\in\mathcal{B}(X)$. We state below the hypotheses under which we will develop our theory in this section.
Definition 5.2.
We say that the sextuple $(X,\,\rho,\,T,\,Y,\,\{Y_{k}\}_{k\in\mathbb{N}},\,\mu)$ is admissible if it has the following properties:
  1. (i)
    $(X,\rho)$ is a compact metric space.
  2. (ii)
    $T\colon X\rightarrow X$ is a Borel-measurable transformation.
  3. (iii)
    $\{Y_{k}\}_{k\in\mathbb{N}}$ is a sequence of pairwise disjoint admissible sets for $T$.
  4. (iv)
    $Y=\bigcup_{k\in\mathbb{N}}Y_{k}$.
  5. (v)
    $\mu\in\mathcal{M}(X,T)\cap\mathcal{P}(X;Y)$.
The following proposition states the uniqueness of the Jacobian and provides a lower bound for the measure-theoretic entropy in terms of the Jacobian.
Proposition 5.3.
Let $(X,\,\rho,\,T,\,Y,\,\{Y_{k}\}_{k\in\mathbb{N}},\,\mu)$ be admissible. Assume that $J\colon X\rightarrow[0,+\infty)$ is a Jacobian for $T$ with respect to $\mu$. Then $J(x)\geqslant 1$ for $\mu$-a.e. $x\in X$, and $h_{\mu}(T)\geqslant\mathopen{\langle}\mu,\log(J)\mathclose{\rangle}$. Moreover, for each Borel function $\widetilde{J}\colon X\rightarrow[0,+\infty)$, $\widetilde{J}$ is a Jacobian for $T$ with respect to $\mu$ if and only if $J(x)=\widetilde{J}(x)$ for $\mu$-a.e. $x\in X$.
The lower bound given above is a classical result in ergodic theory known as the Rokhlin entropy formula. We refer the reader to [Sa99, Theorem 4.2] for its topological Markov shift version, [Co12, Corollary 12.1] for its finite admissible partitions version; see [He25] for the proof in our context. The uniqueness of the Jacobian immediately follows from [Ro49, Theorem 2.7], [PU10, Definition 2.9.2 & Proposition 2.9.5].
Next, we construct a specific Jacobian, define a useful operator, and establish its key properties in the following proposition. With these properties, this operator can be seen as a normalized transfer operator, which is essential for the proof of Theorem 5.5.
Proposition 5.4.
Let $(X,\,\rho,\,T,\,Y,\,\{Y_{k}\}_{k\in\mathbb{N}},\,\mu)$ be admissible. Then the following statements are true:
  1. (i)
    For each $k\in\mathbb{N}$, there exists a nonnegative $\mu$-integrable Borel function $\Phi_{k}$ on $X$ such that
    $\Phi_{k}(x)=0\quad\text{for each }x\notin T(Y_{k})\quad\text{and}\quad\mu\bigl(T^{-1}(B)\cap Y_{k}\bigr)=\int_{B}\!\Phi_{k}\,\mathrm{d}\mu\quad\text{for each }B\in\mathcal{B}(X).$ (5.2)
  2. (ii)
    Define a function $\Psi\colon X\rightarrow\mathbb{R}$ by
    $\Psi(x)\coloneqq 0\quad\text{for each }x\in Y^{c}\quad\text{and}\quad\Psi(x)\coloneqq\Phi_{k}(T(x))\quad\text{for each }k\in\mathbb{N}\text{ and each }x\in Y_{k}.$ (5.3)
    Write $\widetilde{Y}\coloneqq Y\smallsetminus\Psi^{-1}(0)$. Then there exists a Borel function $J_{\mu}\colon X\rightarrow[0,+\infty)$ that is a Jacobian on $\widetilde{Y}$ for $T$ with respect to $\mu$ and satisfies $J_{\mu}(x)\cdot\Psi(x)=1$ for $\mu$-a.e. $x\in X$.
  3. (iii)
    Denote by $L^{+}(X)$ the space of all Borel functions from $X$ to $0,+\infty]$. Then $\mathcal{L}_{\mu}\colon L^{+}(X)\rightarrow L^{+}(X)$ given by
    $\mathcal{L}_{\mu}(u)(x)\coloneqq\sum_{y\in T^{-1}(x)\cap Y}u(y)\Psi(y),\quad\text{for }u\in L^{+}(X)\text{ and }x\in X,$ (5.4)
    satisfies the following: for all $u,v\in L^{+}(X)$, $c\geqslant 0$, and $\mu$-a.e. $x\in X$,
    $\mathcal{L}_{\mu}(\mathbbm{1})(x)$ $=1,$ (5.5)
    $\mathcal{L}_{\mu}(u+v)(x)=\mathcal{L}_{\mu}(u)(x)+\mathcal{L}_{\mu}(v)$ $(x),\quad\mathcal{L}_{\mu}(cu)(x)=c\mathcal{L}_{\mu}(u)(x),\quad\text{and}$ (5.6)
    $\mathopen{\langle}\mu,\mathcal{L}_{\mu}(u)\mathclose{\rangle}$ $=\mathopen{\langle}\mu,u\mathclose{\rangle}.$ (5.7)
We prove (i) and (ii) by using the Radon–Nikodym theorem to construct $\Phi_{k}$} and $J_{\mu}$, then showing their reciprocal relationship with $\Psi$ via change of variables. To establish (iii), we shall apply the monotone convergence theorem repeatedly. We prove (5.5) by testing $\int_{A}\!\mathcal{L}_{\mu}(\mathbbm{1})\,\mathrm{d}\mu=\mu(A)$ for all Borel $A$. Finally, (5.6) follows directly from (5.4), and the integration formula (5.7) is checked on characteristic functions.
Proof.
(i) Fix an arbitrary $k\in\mathbb{N}$. Since $(X,\,\rho,\,T,\,Y,\,\{Y_{k}\}_{k\in\mathbb{N}},\,\mu)$ is admissible, by Definition 5.2 (ii), (iii), and (v), $T$ is Borel measurable, $Y_{k},\,T(Y_{k})\in\mathcal{B}(X)$, and $\mu\in\mathcal{M}(X,T)\cap\mathcal{P}(X;Y)$. Define $\mu_{k}(B)\coloneqq\mu(B)$ and $\widetilde{\mu}_{k}(B)\coloneqq\mu\bigl(\bigl(T|_{Y_{k}}\bigr)^{-1}(B)\bigr)=\mu\bigl(T^{-1}(B)\cap Y_{k}\bigr)$ for each Borel subset $B\subseteq T(Y_{k})$. Then $\mu_{k}$ and $\widetilde{\mu}_{k}$ are both $\sigma$-finite positive Borel measures on $T(Y_{k})$. Since $\mu\in\mathcal{M}(X,T)\cap\mathcal{P}(X;Y)$, we obtain that $\widetilde{\mu}_{k}(B)=\mu\bigl(T^{-1}(B)\cap Y_{k}\bigr)\leqslant\mu\bigl(T^{-1}(B)\bigr)=\mu(B)=\mu_{k}(B)$ for each Borel subset $B\subseteq T(Y_{k})$. This implies that $\widetilde{\mu}_{k}$ is absolutely continuous with respect to $\mu_{k}$. By the Radon–Nikodym theorem, there exists a nonnegative $\mu_{k}$-integrable derivative $\frac{\mathrm{d}\widetilde{\mu}_{k}}{\mathrm{d}\mu_{k}}$. Define $\Phi_{k}:X\rightarrow\mathbb{R}$ by
$\Phi_{k}(x)\coloneqq\begin{cases}\frac{\mathrm{d}\widetilde{\mu}_{k}}{\mathrm{d}\mu_{k}}(x)&\text{if }x\in T(Y_{k});\\ 0&\text{if }x\notin T(Y_{k}).\end{cases}$
By construction, $\Phi_{k}$ is a nonnegative $\mu$-integrable Borel function. Moreover, by the definition of the Radon–Nikodym derivative,
$\mu\bigl(T^{-1}(B)\cap Y_{k}\bigr)=\mu\bigl(T^{-1}(B\cap T(Y_{k}))\cap Y_{k}\bigr)=\int_{B\cap T(Y_{k})}\!\Phi_{k}\,\mathrm{d}\mu=\int_{B}\!\Phi_{k}\,\mathrm{d}\mu\quad\text{for each }B\in\mathcal{B}(X).$
Hence, $\Phi_{k}$ satisfies (5.2), establishing Proposition 5.4 (i).
(ii) By (5.3), $\Psi$ is a Borel function on $X$, and thus $\widetilde{Y}\in\mathcal{B}(X)$. We first prove that $\mu\bigl(\widetilde{Y}\bigr)=1$. By Proposition 5.4 (i) and (5.3), we have $\mu\bigl(\Psi^{-1}(0)\cap Y_{k}\bigr)=\mu\bigl(T^{-1}\bigl(\Phi_{k}^{-1}(0)\bigr)\cap Y_{k}\bigr)=\int_{\Phi_{k}^{-1}(0)}\!\Phi_{k}\,\mathrm{d}\mu=0$ for each $k\in\mathbb{N}$. Since $\{Y_{k}\}_{k\in\mathbb{N}}$ is a sequence of pairwise disjoint Borel subsets with $\mu(Y)=\mu\bigl(\bigcup_{k\in\mathbb{N}}Y_{k}\bigr)=1$ by Definition 5.2 (iv) and (v), we obtain $\mu\bigl(\widetilde{Y}\bigr)=\mu(Y)-\mu\bigl(\Psi^{-1}(0)\cap Y\bigr)=1-\sum_{k\in\mathbb{N}}\mu\bigl(\Psi^{-1}(0)\cap Y_{k}\bigr)=1$.
Now fix an arbitrary $k\in\mathbb{N}$ and write $\widetilde{Y}_{k}\coloneqq Y_{k}\smallsetminus\Psi^{-1}(0)$. Since $\Psi$ is a Borel function, we have $\widetilde{Y}_{k}\in\mathcal{B}(X)$. Define $\nu_{k}(B)\coloneqq\mu(B)$ and $\tilde{\nu}_{k}(B)\coloneqq\mu(T(B))$ for each Borel subset $B\subseteq\widetilde{Y}_{k}$. By [Ke95, Corollary 15.2], it follows from Definition 5.2 (iii) that $T$ is a Borel isomorphism of $Y_{k}$ with $T(Y_{k})$. Hence, by Definition 5.2 (iii) and (v), $\nu_{k}$ and $\tilde{\nu}_{k}$ are both $\sigma$-finite positive Borel measures on $\widetilde{Y}_{k}$. We next prove that $\tilde{\nu}_{k}$ is absolutely continuous with respect to $\nu_{k}$. To this end, consider an arbitrary Borel set $A\subseteq\widetilde{Y}_{k}$ with $\mu(A)=0$; we show that $\mu(T(A))=0$. Indeed, since $T$ is injective on $Y_{k}$, we have $T^{-1}(T(A))\cap Y_{k}=A$. Thus, by Proposition 5.4 (i),
$0=\mu(A)=\mu\bigl(T^{-1}(T(A))\cap Y_{k}\bigr)=\int_{T(A)}\!\Phi_{k}\,\mathrm{d}\mu.$ (5.8)
Since $A\subseteq\widetilde{Y}_{k}=Y_{k}\smallsetminus\Psi^{-1}(0)$ and $\Psi(x)=\Phi_{k}(T(x))$ for each $x\in Y_{k}$ by (5.3), we have $\Phi_{k}(x)>0$ for each $x\in T(A)$. Combined with (5.8), this implies that $\mu(T(A))=0$. Thus, $\tilde{\nu}_{k}$ is absolutely continuous with respect to $\nu_{k}$. By the Radon–Nikodym theorem, there exists a nonnegative $\mu$-integrable derivative $\frac{\mathrm{d}\tilde{\nu}_{k}}{\mathrm{d}\nu_{k}}$.
We define $J_{\mu}$ by
[…]
We now verify that $J_{\mu}$ is a Jacobian on $\widetilde{Y}$ for $T$ with respect to $\mu$. Since $\widetilde{Y}=\bigcup_{k\in\mathbb{N}}\widetilde{Y}_{k}$, by the definitions of $\{\nu_{k}\}_{k\in\mathbb{N}}$ and $\{\tilde{\nu}_{k}\}_{k\in\mathbb{N}}$, it follows from (5.9) that
$\mu(T(A))$ $=\sum_{k\in\mathbb{N}}\mu\bigl(T\bigl(A\cap\widetilde{Y}_{k}\bigr)\bigr)=\sum_{k\in\mathbb{N}}\tilde{\nu}_{k}\bigl(A\cap\widetilde{Y}_{k}\bigr)=\sum_{k\in\mathbb{N}}\int_{A\cap\widetilde{Y}_{k}}\!\frac{\mathrm{d}\tilde{\nu}_{k}}{\mathrm{d}\nu_{k}}\,\mathrm{d}\nu_{k}$
$=\sum_{k\in\mathbb{N}}\int_{A\cap\widetilde{Y}_{k}}\!\frac{\mathrm{d}\tilde{\nu}_{k}}{\mathrm{d}\nu_{k}}\,\mathrm{d}\mu=\sum_{k\in\mathbb{N}}\int_{A\cap\widetilde{Y}_{k}}\!J_{\mu}\,\mathrm{d}\mu=\int_{A\cap\widetilde{Y}}\!J_{\mu}\,\mathrm{d}\mu=\int_{A}\!J_{\mu}\,\mathrm{d}\mu,$
for each admissible set $A\subseteq\widetilde{Y}$ for $T$. Since $\mu\bigl(\widetilde{Y}\bigr)=1$, we conclude that $J_{\mu}$ is a Jacobian on $\widetilde{Y}$.
Finally, we verify that $J_{\mu}(x)\cdot\Psi(x)=1$ for $\mu$-a.e. $x\in X$. By [Ke95, Corollary 15.2], it follows from Definition 5.2 (iii) that $T$ is a Borel isomorphism of $Y_{k}$ with $T(Y_{k})$ for each $k\in\mathbb{N}$. Since $J_{\mu}$ is a Jacobian on $\widetilde{Y}$ for $T$ with respect to $\mu$, by the definition of $\bigl\{\tilde{\nu}_{k}\bigr\}_{k\in\mathbb{N}}$, we have
$\tilde{\nu}_{k}(A)=\mu(T(A))=\int_{A}\!J_{\mu}\,\mathrm{d}\mu\quad\text{for each }k\in\mathbb{N}\text{ and each Borel }A\subseteq\widetilde{Y}_{k}.$ (5.10)
Moreover, by Proposition 5.4 (i), we obtain that
$\mu(A)=\mu\bigl(T^{-1}(T(A))\cap Y_{k}\bigr)=\int_{T(A)}\!\Phi_{k}\,\mathrm{d}\mu=\int_{A}\!(\Phi_{k}\circ T)\,\mathrm{d}(\mu\circ T)=\int_{A}\!\Psi\,\mathrm{d}\tilde{\nu}_{k}$
for each $k\in\mathbb{N}$ and each Borel $A\subseteq\widetilde{Y}_{k}$. Thus, by (5.10) and the change-of-variable formula, we obtain $J_{\mu}(x)\cdot\Psi(x)=1$ for $\mu$-a.e. $x\in\bigcup_{k\in\mathbb{N}}\widetilde{Y}_{k}$. Since $\mu\bigl(\bigcup_{k\in\mathbb{N}}\widetilde{Y}_{k}\bigr)=\mu\bigl(\widetilde{Y}\bigr)=1$, we conclude that $J_{\mu}(x)\cdot\Psi(x)=1$ for $\mu$-a.e. $x\in X$.
(iii) By Definition 5.2 (iii) and (iv), we have that $Y=\bigcup_{k\in\mathbb{N}}Y_{k}$ and $T$ is injective on $Y_{k}$ for each $k\in\mathbb{N}$. Since $\Psi$ is nonnegative and Borel, the expression for $\mathcal{L}_{\mu}(u)(x)$ given in (5.4) is the sum of countably (possibly infinitely) many nonnegative terms for each $u\in L^{+}(X)$ and each $x\in X$. Hence, $\mathcal{L}_{\mu}(u)\in L^{+}(X)$ for each $u\in L^{+}(X)$, and $\mathcal{L}_{\mu}\colon L^{+}(X)\rightarrow L^{+}(X)$ is monotone.
We now establish (5.5). By (5.4), we have that
$\mathcal{L}_{\mu}(\mathbbm{1}_{A})(x)=\sum_{y\in T^{-1}(x)\cap Y}\mathbbm{1}_{A}(y)\Psi(y)=\sum_{y\in T^{-1}(x)\cap A\cap Y}\Psi(y)\quad\text{for each }A\in\mathcal{B}(X).$ (5.11)
Then by (5.11), (5.3), $Y=\bigcup_{k\in\mathbb{N}}Y_{k}$, (5.2), the monotone convergence theorem, and $\mu\in\mathcal{M}(X,T)\cap\mathcal{P}(X;Y)$, we obtain that
$\int_{A}\!\mathcal{L}_{\mu}(\mathbbm{1})\,\mathrm{d}\mu$ $=\int_{A}\!\biggl(\sum_{y\in T^{-1}(x)\cap Y}\Psi(y)\biggr)\,\mathrm{d}\mu(x)=\int_{A}\!\biggl(\sum_{k\in\mathbb{N}:T^{-1}(x)\cap Y_{k}\neq\emptyset}\Phi_{k}\biggr)\,\mathrm{d}\mu=\int_{A}\!\biggl(\sum_{k\in\mathbb{N}}\Phi_{k}\biggr)\,\mathrm{d}\mu$
$=\sum_{k\in\mathbb{N}}\int_{A}\!\Phi_{k}\,\mathrm{d}\mu=\sum_{k\in\mathbb{N}}\mu\bigl(T^{-1}(A)\cap Y_{k}\bigr)=\mu\bigl(T^{-1}(A)\cap Y\bigr)=\mu\bigl(T^{-1}(A)\bigr)=\mu(A)$
for each $A\in\mathcal{B}(X)$. This establishes (5.5).
To establish (5.6), consider arbitrary $u,\,v\in L^{+}(X)$ and $c\geqslant 0$. By (5.4), we have $\mathcal{L}_{\mu}(u+v)(x)=\mathcal{L}_{\mu}(u)(x)+\mathcal{L}_{\mu}(v)(x)$ for each $x\in X$. Moreover, $\mathcal{L}_{\mu}(cu)(x)=c\mathcal{L}_{\mu}(u)(x)$ for each $x\in X$ and each nonnegative $c\in\mathbb{Q}$. By the monotonicity of $\mathcal{L}_{\mu}$, it follows from the monotone convergence theorem that $\mathcal{L}_{\mu}(cu)(x)=c\mathcal{L}_{\mu}(u)(x)$ for each $c\geqslant 0$. This establishes (5.6).
Finally, we establish (5.7). Consider an arbitrary $D\in\mathcal{B}(X)$. Fix $k\in\mathbb{N}$ and write $D_{k}\coloneqq D\cap Y_{k}$. By (5.11), $\mathcal{L}_{\mu}\bigl(\mathbbm{1}_{D_{k}}\bigr)(x)=0$ for each $x\notin T(D_{k})$. Now consider an arbitrary $x\in T(D_{k})$. Note that $T$ is injective on $Y_{k}$ by Definition 5.2 (iii). Then the set $T^{-1}(x)\cap D_{k}$ contains a unique element, which we denote by $y_{k}(x)$. Using (5.11) and (5.3), we deduce that $\mathcal{L}_{\mu}\bigl(\mathbbm{1}_{D_{k}}\bigr)(x)=\Psi(y_{k}(x))=\Phi_{k}(x)$. Integrating this and applying Proposition 5.4 (i) yields
$\bigl\langle\mu,\mathcal{L}_{\mu}\bigl(\mathbbm{1}_{D_{k}}\bigr)\bigr\rangle=\int_{T(D_{k})}\!\Phi_{k}\,\mathrm{d}\mu=\mu\bigl(T^{-1}(T(D_{k}))\cap Y_{k}\bigr)=\mu(D_{k}).$
Thus it follows from (5.11) and Definition 5.2 (iv) that $\bigl\langle\mu,\mathcal{L}_{\mu}(\mathbbm{1}_{D})\bigr\rangle=\sum_{k\in\mathbb{N}}\bigl\langle\mu,\mathcal{L}_{\mu}\bigl(\mathbbm{1}_{D_{k}}\bigr)\bigr\rangle=\sum_{k\in\mathbb{N}}\mu(D_{k})=\mu(D\cap Y)=\mu(D)$, where the last equality holds as $\mu\in\mathcal{P}(X;Y)$. This establishes the identity for indicator functions. The general result (5.7) for all nonnegative measurable functions then follows from a standard argument using the monotone convergence theorem. ∎
The following theorem characterizes Jacobians in terms of an eigenfunction equation for the transfer operator. This characterization will be key in identifying equilibrium states.
Theorem 5.5.
Let $(X,\,\rho,\,T,\,Y,\,\{Y_{k}\}_{k\in\mathbb{N}},\,\mu)$ be admissible. Assume that $J_{\mu}$ is given as in Proposition 5.4, and $J$ is a positive Borel function on $X$. Then $J$ is a Jacobian on $Y$ for $T$ with respect to $\mu$ if and only if
$\sum_{y\in T^{-1}(x)\cap Y}\frac{1}{J(y)}$ $=1\quad\text{for }\mu\text{-a.e.\ }x\in X,$ (5.12)
$\log(J)\in L^{1}(\mu),\quad$ $\text{and}\quad\mathopen{\langle}\mu,\log(J)\mathclose{\rangle}=\bigl\langle\mu,\log\bigl(J_{\mu}\bigr)\bigr\rangle.$ (5.13)
Proof.
Let $\{\Phi_{k}\}_{k\in\mathbb{N}},\,\Psi$, and $\mathcal{L}_{\mu}$ be as in Proposition 5.4. Write $\widetilde{Y}\coloneqq Y\smallsetminus\Psi^{-1}(0)$.
We first prove the forward implication. Assume that $J$ is a Jacobian on $Y$. We show that $J$ satisfies (5.12) and (5.13). By Proposition 5.4 (ii), $J_{\mu}$ is a Jacobian on $\widetilde{Y}$. By Proposition 5.3, $J(x)=J_{\mu}(x)\geqslant 1$ for $\mu$-a.e. $x\in X$. Thus, $\log(J),\,\log\bigl(J_{\mu}\bigr)\in L^{1}(\mu)$ and $\mathopen{\langle}\mu,\log(J)\mathclose{\rangle}=\bigl\langle\mu,\log\bigl(J_{\mu}\bigr)\bigr\rangle$. This establishes (5.13).
We now establish (5.12). By Proposition 5.4 (ii), there exists a Borel set $H_{1}\subseteq\widetilde{Y}$ such that $\mu(H_{1})=1$ and $J(x)\cdot\Psi(x)=1$ for each $x\in H_{1}$. For each $x\notin T(Y\smallsetminus H_{1})$, we have $T^{-1}(x)\cap Y\subseteq H_{1}$, and hence $J(y)\cdot\Psi(y)=1$ for each $y\in T^{-1}(x)\cap Y$. Since $J$ is a Jacobian on $Y$, it follows from $\mu(Y\smallsetminus H_{1})=0$ that $\mu(T(Y\smallsetminus H_{1}))=0$. Thus, for $\mu$-a.e. $x\in X$, we have $J(y)\cdot\Psi(y)=1$ for each $y\in T^{-1}(x)\cap Y$. Combined with (5.5), this implies that $\sum_{y\in T^{-1}(x)\cap Y}\frac{1}{J(y)}=\sum_{y\in T^{-1}(x)\cap Y}\Psi(y)=\mathcal{L}_{\mu}(\mathbbm{1}_{X})(x)=1$ for $\mu$-a.e. $x\in X$.
We now establish the backward implication. Assume that $J$ satisfies (5.12) and (5.13). We show that $J$ is a Jacobian on $Y$. By Proposition 5.4 (ii), there exists a Borel set $H_{2}\subseteq\widetilde{Y}$ such that $\mu(H_{2})=1$ and $J_{\mu}(x)\cdot\Psi(x)=1$ for each $x\in H_{2}$. For each $x\notin T\bigl(\widetilde{Y}\smallsetminus H_{2}\bigr)$, we have $T^{-1}(x)\cap\widetilde{Y}\subseteq H_{2}$, and hence $J_{\mu}(y)\cdot\Psi(y)=1$ for each $y\in T^{-1}(x)\cap\widetilde{Y}$. Since $J_{\mu}$ is a Jacobian on $\widetilde{Y}$, it follows from $\mu\bigl(\widetilde{Y}\smallsetminus H_{2}\bigr)=0$ that $\mu\bigl(T\bigl(\widetilde{Y}\smallsetminus H_{2}\bigr)\bigr)=0$. Thus, we obtain that
$\text{for }\mu\text{-a.e.\ }x\in X,\quad J_{\mu}(y)\cdot\Psi(y)=1\quad\text{for each }y\in T^{-1}(x)\cap\widetilde{Y}$. (5.14)
Since $J$ is a positive Borel function and $J_{\mu}$ is a nonnegative Borel function, we have $J_{\mu}/J\in L^{+}(X)$. By (5.4), $\widetilde{Y}=Y\smallsetminus\Psi^{-1}(0)$, (5.14), and (5.12), we obtain that for $\mu$-a.e. $x\in X$,
$\mathcal{L}_{\mu}\bigl(J_{\mu}/J\bigr)(x)$ $=\sum_{y\in T^{-1}(x)\cap Y}\frac{J_{\mu}(y)\Psi(y)}{J(y)}=\sum_{y\in T^{-1}(x)\cap\widetilde{Y}}\frac{J_{\mu}(y)\Psi(y)}{J(y)}$ (5.15)
$=\sum_{y\in T^{-1}(x)\cap\widetilde{Y}}\frac{1}{J(y)}\leqslant\sum_{y\in T^{-1}(x)\cap Y}\frac{1}{J(y)}=1.$
By (5.7) and (5.13), we obtain that
$1=\mathopen{\langle}\mu,\mathbbm{1}_{X}\mathclose{\rangle}\geqslant\mathopen{\big\langle}\mu,\mathcal{L}_{\mu}\bigl(J_{\mu}\big/J\bigr)\mathclose{\big\rangle}=\mathopen{\big\langle}\mu,J_{\mu}\big/J\mathclose{\big\rangle}\geqslant 1-\mathopen{\langle}\mu,\log(J)\mathclose{\rangle}+\mathopen{\big\langle}\mu,\log\bigl(J_{\mu}\bigr)\mathclose{\big\rangle}=1.$ (5.16)
[…]
Combining Theorem 5.5 with the definition of equilibrium states yields the following result, which provides a verifiable criterion for identifying equilibrium states.
Recall the definition of $\mathcal{M}(X,T;Y,J)$ from (1.2). For each admissible sextuple $(X,\,\rho,\,T,\,Y,\,\{Y_{k}\}_{k\in\mathbb{N}},\,\mu)$ and each Borel function $\phi\colon X\rightarrow\mathbb{R}$, we define
$\mathcal{E}_{0}(T,\phi)\coloneqq\bigl\{\mu\in\mathcal{E}(T,\phi)\cap\mathcal{P}(X;Y):h_{\mu}(T)=\bigl\langle\mu,\log\bigl(J_{\mu}\bigr)\bigr\rangle\bigr\},$ (5.17)
where $J_{\mu}$ is given as in Proposition 5.4 for the measure $\mu$.
Theorem 5.6.
Let $(X,\,\rho,\,T,\,Y,\,\{Y_{k}\}_{k\in\mathbb{N}},\,\mu)$ be admissible, and $\phi\colon X\rightarrow\mathbb{R}$ be a Borel function with $\mathopen{\langle}\mu,\phi\mathclose{\rangle}\in\mathbb{R}$. Suppose that $J$ is a positive Borel function on $X$ which satisfies the following properties:
  1. (i)
    There exists a bounded Borel function $h\colon X\rightarrow\mathbb{R}$ such that for each $x\in Y$,
    $J(x)=\exp(P(T,\phi)-\phi(x)+h(T(x))-h(x)).$
  2. (ii)
    $\sum\limits_{y\in T^{-1}(x)\cap Y}\frac{1}{J(y)}=1$ for each $x\in\bigcap_{i\in\mathbb{N}_{0}}T^{i}(Y)$.
Then $\mu\in\mathcal{M}(X,T;Y,J)$ if and only if $\mu\in\mathcal{E}_{0}(T,\phi)$.
Proof.
Since $h$ is a bounded Borel function and $\mu\in\mathcal{M}(X,T)$, we have $\mathopen{\langle}\mu,h\mathclose{\rangle}=\mathopen{\langle}\mu,h\circ T\mathclose{\rangle}\in(-\infty,+\infty)$. Since $J$ satisfies property (i) of Theorem 5.6, $\mu\in\mathcal{P}(X;Y)$, and $\mathopen{\langle}\mu,\phi\mathclose{\rangle}<+\infty$, we obtain that
$\mathopen{\langle}\mu,\log(J)\mathclose{\rangle}=P(T,\phi)-\mathopen{\langle}\mu,\phi\mathclose{\rangle}+\mathopen{\langle}\mu,h\circ T\mathclose{\rangle}-\mathopen{\langle}\mu,h\mathclose{\rangle}=P(T,\phi)-\mathopen{\langle}\mu,\phi\mathclose{\rangle}.$ (5.18)
By Proposition 5.4 (ii), there exists a Borel function $J_{\mu}\colon X\rightarrow\mathbb{R}^{+}$ which is a Jacobian for $T$ with respect to $\mu$.
We first establish the forward implication. Let $\mu\in\mathcal{M}(X,T;Y,J)$. Since $\mu\in\mathcal{P}(X;Y)$, it follows from (1.2) that $J_{\mu}(x)\geqslant J(x)$ for $\mu$-a.e. $x\in X$. By (5.18), (3.2), and Proposition 5.3, we have
$\mathopen{\langle}\mu,\log(J)\mathclose{\rangle}=P(T,\phi)-\mathopen{\langle}\mu,\phi\mathclose{\rangle}\geqslant h_{\mu}(T)\geqslant\bigl\langle\mu,\log\bigl(J_{\mu}\bigr)\bigr\rangle.$ (5.19)
Since $J_{\mu}(x)\geqslant J(x)$ for $\mu$-a.e. $x\in X$, the inequalities in (5.19) must be equalities, which implies $\mu\in\mathcal{E}(T,\phi)$ and $h_{\mu}(T)=\bigl\langle\mu,\log\bigl(J_{\mu}\bigr)\bigr\rangle$. This implies $\mu\in\mathcal{E}_{0}(T,\phi)$ (recall (5.17)).
We now establish the backward implication. Suppose $\mu\in\mathcal{E}(T,\phi)$ and $h_{\mu}(T)=\bigl\langle\mu,\log\bigl(J_{\mu}\bigr)\bigr\rangle$. By (3.2) and (5.18), we obtain $\bigl\langle\mu,\log\bigl(J_{\mu}\bigr)\bigr\rangle=h_{\mu}(T)=P(T,\phi)-\mathopen{\langle}\mu,\phi\mathclose{\rangle}=\mathopen{\langle}\mu,\log(J)\mathclose{\rangle}$. Since $\mu\in\mathcal{M}(X,T)\cap\mathcal{P}(X;Y)$, we have $1\geqslant\mu\bigl(T^{i}(Y)\bigr)\geqslant\mu(Y)=1$ for each $i\in\mathbb{N}_{0}$. Thus $\mu\bigl(\bigcap_{i\in\mathbb{N}_{0}}T^{i}(Y)\bigr)=1$. Since $J$ satisfies property (ii) in Theorem 5.6, by Theorem 5.5, $J$ is a Jacobian on $Y$ for $T$ with respect to $\mu$. By (1.2), we conclude that $\mu\in\mathcal{M}(X,T;Y,J)$. ∎
5.2. Proof of Theorem 1.3
Theorem 1.3 follows immediately from the following theorem.
Theorem 5.7.
Let $(X,\,\rho,\,\mathcal{S},\,\{X_{n}\}_{n\in\mathbb{N}},\,\{T_{n}\}_{n\in\mathbb{N}})$ be a uniformly computable system, and $Y_{n}$ be an open subset of $X_{n}$ for each $n\in\mathbb{N}$. Assume that there exist two recursively enumerable sets $K$ and $L$ with $L\subseteq\mathbb{N}\times K$, and a sequence $\{Y_{n,k}\}_{(n,k)\in L}$ of uniformly lower semi-computable open sets in $(X,\,\rho,\,\mathcal{S})$ such that $Y_{n,k}$ is admissible for $T_{n}$, and $Y_{n}=\bigcup_{(n,k)\in L_{n}}Y_{n,k}$, where $L_{n}\coloneqq\{(n,k)\in L:k\in K\}$ for each $n\in\mathbb{N}$.
Assume that $\{J_{n}\}_{n\in\mathbb{N}}$ is a sequence of uniformly lower semi-computable functions $J_{n}\colon X\rightarrow[0,+\infty)$ with respect to $\{Y_{n}\}_{n\in\mathbb{N}}$ such that $J_{n}$ is nonnegative on $Y_{n}$ and Borel for each $n\in\mathbb{N}$. Then $\{\mathcal{M}(X,T_{n};Y_{n},J_{n})\}_{n\in\mathbb{N}}$ is uniformly recursively compact in $(\mathcal{P}(X),\,W_{\rho},\,\mathcal{Q}_{\mathcal{S}})$.
Proof of Theorem 1.3.
By Theorem 5.7, $\{\mathcal{M}(X,T_{n};Y_{n},J_{n})\}_{n\in\mathbb{N}}$ is uniformly recursively compact in $\bigl(\mathcal{P}(X),\,W_{\rho},\,\mathcal{Q}_{\mathcal{S}}\bigr)$. Since $\{\mathcal{K}_{n}\}_{n\in\mathbb{N}}$ is uniformly recursively compact, by Proposition 3.21 (i), $\{\mathcal{M}(X,T_{n};Y_{n},J_{n})\cap\mathcal{K}_{n}\}_{n\in\mathbb{N}}$ is uniformly recursively compact. Note that $\mathcal{M}(X,T_{n};Y_{n},J_{n})\cap\mathcal{K}_{n}=\{\mu_{n}\}$ for each $n\in\mathbb{N}$ by (1.3). Therefore, by Proposition 3.20 (i), $\{\mu_{n}\}_{n\in\mathbb{N}}$ is uniformly computable in $(\mathcal{P}(X),\,W_{\rho},\,\mathcal{Q}_{\mathcal{S}})$. ∎
We devote the rest of this subsection to a proof of Theorem 5.7. The proof proceeds by constructing a sequence of uniformly lower semi-computable open sets and expressing the set $\mathcal{M}(X,T_{n};Y_{n},J_{n})$ as the complement of their union. The uniform recursive compactness then follows from established properties of recursively compact spaces.
Proof of Theorem 5.7.
By Proposition 3.31, there exists a sequence $\{\tau_{s}\}_{s\in\mathbb{N}}$ of uniformly computable functions $\tau_{s}\colon X\rightarrow\mathbb{R}$ such that $\{\tau_{s}:s\in\mathbb{N}\}$ is dense in $C(X)$. By the computability of the absolute value function, the sequence $\bigl\{\tau_{s}^{+}\bigr\}_{s\in\mathbb{N}}$ is a sequence of uniformly computable functions. Since $\{Y_{n,k}\}_{(n,k)\in L}$ is uniformly lower semi-computable open, by Proposition 3.17, there exists a sequence $\{h_{m,n,k}\}_{(m,(n,k))\in\mathbb{N}\times L}$ of uniformly computable functions $h_{m,n,k}\colon X\rightarrow\mathbb{R}$ such that for each $(n,k)\in L$, the following properties are satisfied:
  1. (i)
    For each $x\in X$, $\{h_{m,n,k}(x)\}_{m\in\mathbb{N}}$ is non-decreasing and $h_{m,n,k}(x)\to\mathbbm{1}_{Y_{n,k}}(x)$ as $m\to+\infty$.
  2. (ii)
    For each $m\in\mathbb{N}$, $h_{m,n,k}(x)\geqslant 0$ for each $x\in X$ and $h_{m,n,k}(x)=0$ for each $x\notin Y_{n,k}$.
We define for all $m,s\in\mathbb{N},\,x\in X$, and $(n,k)\in L$,
$W_{m,s}^{n,k}(x)$ $\coloneqq\begin{cases}\sup\bigl\{\tau^{+}_{s}(y)\cdot h_{m,n,k}(y):y\in T_{n}^{-1}(x)\bigr\}&\textrm{if }x\in T_{n}(Y_{n,k});\\ 0&\textrm{otherwise},\end{cases}$ (5.20)
$V_{m,s}^{n,k}(x)$ $\coloneqq J_{n}(x)\cdot\tau^{+}_{s}(x)\cdot h_{m,n,k}(x),\qquad\qquad\qquad\,\,\,\text{and}$ (5.21)
$\Psi_{m,s}^{n,k}$ $\coloneqq\bigl\{\mu\in\mathcal{P}(X):\bigl\langle\mu,W_{m,s}^{n,k}-V_{m,s}^{n,k}\bigr\rangle<0\bigr\}.$ (5.22)
We first establish the following claim.
Claim 1. $\bigl\{\Psi_{m,s}^{n,k}\big\}_{(m,s,(n,k))\in\mathbb{N}^{2}\times L}$ is uniformly lower semi-computable open in $(\mathcal{P}(X),\,W_{\rho},\,\mathcal{Q}_{\mathcal{S}})$.
Proof of Claim 1. Let $\mathcal{S}_{\mathbb{Q}}=\{q_{v}\}_{v\in\mathbb{N}}$ (see Subsection 3.1). First, we show that $\bigl\{W_{m,s}^{n,k}\bigr\}_{(m,s,(n,k))\in\mathbb{N}^{2}\times L}$ is a sequence of uniformly upper semi-computable functions. Denote $Q_{v}\coloneqq(-\infty,q_{v})$ for each $v\in\mathbb{N}$. Indeed, it is not hard to derive from (5.20) and property (ii) of $\{h_{m,n,k}\}_{(m,(n,k))\in\mathbb{N}\times L}$ that
$\bigl(W_{m,s}^{n,k}\bigr)^{-1}(Q_{v}^{c})$ $=\begin{cases}X&\textrm{if }q_{v}\leqslant 0;\\ T_{n}\bigl(\bigl(\tau_{s}^{+}\cdot h_{m,n,k}\bigr)^{-1}(Q_{v}^{c})\bigr)&\textrm{otherwise}\end{cases}$ (5.23)
for all $m,\,s,\,v\in\mathbb{N}$, and $(n,k)\in L$.
By the uniform computability of $\bigl\{\tau_{s}^{+}\bigr\}_{s\in\mathbb{N}}$ and $\{h_{m,n,k}\}_{(m,(n,k))\in\mathbb{N}\times L}$, Proposition 3.23 implies that $\bigl\{\tau_{s}^{+}\cdot h_{m,n,k}\bigr\}_{(m,s,(n,k))\in\mathbb{N}^{2}\times L}$ is a sequence of uniformly computable functions. Since $\{Q_{v}\}_{v\in\mathbb{N}}$ is uniformly lower semi-computable open in $\bigl(\mathbb{R},\,d_{\mathbb{R}},\,\mathcal{S}_{\mathbb{Q}}\bigr)$, by Proposition 3.14, $\bigl\{\bigl(\tau_{s}^{+}\cdot h_{m,n,k}\bigr)^{-1}(Q_{v})\bigr\}_{(m,s,v,(n,k))\in\mathbb{N}^{3}\times L}$ is uniformly lower semi-computable open in $(X,\,\rho,\,\mathcal{S})$. Note that $\bigl(\tau_{s}^{+}\cdot h_{m,n,k}\bigr)^{-1}(Q_{v}^{c})=\bigl(\bigl(\tau_{s}^{+}\cdot h_{m,n,k}\bigr)^{-1}(Q_{v})\bigr)^{c}$ for all $m,\,s,\,v\in\mathbb{N}$, and $(n,k)\in L$. Then since $X$ is recursively compact in $(X,\,\rho,\,\mathcal{S})$, by Proposition 3.20 (iii), $\bigl\{\bigl(\tau_{s}^{+}\cdot h_{m,n,k}\bigr)^{-1}(Q_{v}^{c})\bigr\}_{(m,s,v,(n,k))\in\mathbb{N}^{3}\times L}$ is uniformly recursively compact. Note that by property (ii) of $\{h_{m,n,k}\}_{(m,(n,k))\in\mathbb{N}\times L}$ and the hypotheses of Theorem 5.7, we have that $\bigl(\tau_{s}^{+}\cdot h_{m,n,k}\bigr)^{-1}(Q_{v}^{c})\subseteq Y_{n,k}\subseteq X_{n}$ for all $m,\,s,\,v\in\mathbb{N}$, and $(n,k)\in L$. Hence, since $\{T_{n}\}_{n\in\mathbb{N}}$ is a sequence of uniformly computable functions with respect to $\{X_{n}\}_{n\in\mathbb{N}}$, by Proposition 3.20 (v), we obtain that $\bigl\{T_{n}\bigl(\bigl(\tau_{s}^{+}\cdot h_{m,n,k}\bigr)^{-1}(Q_{v}^{c})\bigr)\bigr\}_{(m,s,v,(n,k))\in\mathbb{N}^{3}\times L}$ is uniformly recursively compact. Since $X$ is recursively compact, and $\mathcal{S}_{\mathbb{Q}}=\{q_{v}\}_{v\in\mathbb{N}}$, by (5.23), $\bigl\{\bigl(W_{m,s}^{n,k}\bigr)^{-1}(Q_{v}^{c})\bigr\}_{(m,s,v,(n,k))\in\mathbb{N}^{3}\times L}$ is uniformly recursively compact. Thus, since $\bigl(W_{m,s}^{n,k}\bigr)^{-1}(Q_{v}^{c})=\bigl(\bigl(W_{m,s}^{n,k}\bigr)^{-1}(Q_{v})\bigr)^{c}$ for all $m,\,s,\,v\in\mathbb{N}$, and $(n,k)\in L$, by Proposition 3.20 (ii), the sequence $\bigl\{\bigl(W_{m,s}^{n,k}\bigr)^{-1}(Q_{v})\bigr\}_{(m,s,v,(n,k))\in\mathbb{N}^{3}\times L}$ is uniformly lower semi-computable open. Hence, it follows from Proposition 3.16 that $\bigl\{W_{m,s}^{n,k}\bigr\}_{(m,s,(n,k))\in\mathbb{N}^{2}\times L}$ is a sequence of uniformly upper semi-computable functions.
We next show that $\bigl\{V_{m,s}^{n,k}\bigr\}_{(m,s,(n,k))\in\mathbb{N}^{2}\times L}$ is a sequence of uniformly lower semi-computable functions. Denote $R_{v}\coloneqq(q_{v},+\infty)$ for each $v\in\mathbb{N}$. Since $\bigl\{\tau_{s}^{+}\bigr\}_{s\in\mathbb{N}}$ and $\{h_{m,n,k}\}_{(m,(n,k))\in\mathbb{N}\times L}$ are two sequences of uniformly computable nonnegative functions, and $\{J_{n}\}_{n\in\mathbb{N}}$ is a sequence of uniformly lower semi-computable functions with respect to $\{Y_{n}\}_{n\in\mathbb{N}}$, by (5.21), Definition 3.15, and Proposition 3.23, $\bigl\{V_{m,s}^{n,k}\bigr\}_{(m,s,(n,k))\in\mathbb{N}^{2}\times L}$ is a sequence of uniformly lower semi-computable functions with respect to $\{Y_{n}\}_{n\in\mathbb{N}}$. Hence, by Proposition 3.16, there exists a sequence $\bigl\{Y_{m,s,v}^{n,k}\bigr\}_{(m,s,v,(n,k))\in\mathbb{N}^{3}\times L}$ of uniformly lower semi-computable open sets such that $\bigl(V_{m,s}^{n,k}\bigr)^{-1}(R_{v})\cap Y_{n}=Y_{m,s,v}^{n,k}\cap Y_{n}$ for all $m,\,s,\,v\in\mathbb{N}$ and $(n,k)\in L$. Indeed, it is not hard to derive from (5.21) and property (ii) of $\{h_{m,n,k}\}_{(m,(n,k))\in\mathbb{N}\times L}$ that
$\bigl(V_{m,s}^{n,k}\bigr)^{-1}(R_{v})$ $=\begin{cases}X&\textrm{if }q_{v}<0;\\ Y_{m,s,v}^{n,k}\cap Y_{n}&\textrm{otherwise}\end{cases}\quad\text{ for all }m,\,s,\,v\in\mathbb{N},\text{ and }(n,k)\in L.$ (5.24)
Since $\{Y_{n}\}_{n\in\mathbb{N}}$ and $\bigl\{Y_{m,s,v}^{n,k}\bigr\}_{(m,s,v,(n,k))\in\mathbb{N}^{3}\times L}$ are both uniformly lower semi-computable open in $(X,\,\rho,\,\mathcal{S})$, by Proposition 3.20 (ii), (iii), and Proposition 3.21 (ii), it follows from the recursive compactness of $X$ that $\bigl\{Y_{m,s,v}^{n,k}\cap Y_{n}\bigr\}_{(m,s,v,(n,k))\in\mathbb{N}^{3}\times L}$ is uniformly lower semi-computable open. Since $\mathcal{S}_{\mathbb{Q}}=\{q_{v}\}_{v\in\mathbb{N}}$, by (5.24), $\bigl\{\bigl(V_{m,s}^{n,k}\bigr)^{-1}(R_{v})\bigr\}_{(m,s,v,(n,k))\in\mathbb{N}^{3}\times L}$ is uniformly lower semi-computable open. Thus, by Proposition 3.16, we conclude that $\bigl\{V_{m,s}^{n,k}\bigr\}_{(m,s,(n,k))\in\mathbb{N}^{2}\times L}$ is a sequence of uniformly lower semi-computable functions.
Since $\bigl\{W_{m,s}^{n,k}\bigr\}_{(m,s,(n,k))\in\mathbb{N}^{2}\times L}$ (resp. $\bigl\{V_{m,s}^{n,k}\bigr\}_{(m,s,(n,k))\in\mathbb{N}^{2}\times L}$) is a sequence of uniformly upper (resp. lower) semi-computable functions, by Definition 3.15, $\bigl\{W_{m,s}^{n,k}-V_{m,s}^{n,k}\bigr\}_{(m,s,(n,k))\in\mathbb{N}^{2}\times L}$ is a sequence of uniformly upper semi-computable functions. Thus, by Corollary 3.28, (5.22), and Proposition 3.16, we conclude that $\bigl\{\Psi_{m,s}^{n,k}\bigr\}_{(m,s,n,k)\in\mathbb{N}^{2}\times L}$ is uniformly lower semi-computable open in $(\mathcal{P}(X),\,W_{\rho},\,\mathcal{Q}_{\mathcal{S}})$. This establishes Claim 1.
We now characterize the set $\mathcal{M}(X,T_{n};Y_{n},J_{n})$ (recall (1.2)).
Claim 2. $\mathcal{M}(X,T_{n};Y_{n},J_{n})=\mathcal{P}(X)\smallsetminus\bigcup_{(m,s)\in\mathbb{N}^{2}}\bigcup_{(n,k)\in L_{n}}\Psi_{m,s}^{n,k}$ for each $n\in\mathbb{N}$.
Proof of Claim 2. We fix an arbitrary $n\in\mathbb{N}$ and establish our claim by showing that these two sets are mutually inclusive for $n$.
Suppose $\mu\in\mathcal{P}(X)\smallsetminus\bigcup_{(m,s)\in\mathbb{N}^{2}}\bigcup_{(n,k)\in L_{n}}\Psi_{m,s}^{n,k}$. By (5.22), we have $\bigl\langle\mu,W_{m,s}^{n,k}\bigr\rangle\geqslant\bigl\langle\mu,V_{m,s}^{n,k}\bigr\rangle$ for all $m,\,s\in\mathbb{N}$, and $(n,k)\in L_{n}$. Consider an arbitrary $(n,k)\in L_{n}$. Since $Y_{n,k}$ is admissible for $T_{n}$, by [Ke95, Corollary 15.2], $T_{n}$ is a Borel isomorphism of $Y_{n,k}$ with $T_{n}(Y_{n,k})$. It follows from (5.20) and (5.21) that
$\int_{T_{n}(Y_{n,k})}\!\bigl(\tau_{s}^{+}\cdot h_{m,n,k}\bigr)\circ\bigl(T_{n}|_{Y_{n,k}}\bigr)^{-1}\,\mathrm{d}\mu\geqslant\bigl\langle\mu,J_{n}\cdot\tau_{s}^{+}\cdot h_{m,n,k}\bigr\rangle\quad\text{for all }m,\,s\in\mathbb{N}.$ (5.25)
Letting $m\to+\infty$ in (5.25) and applying the monotone convergence theorem, we derive from properties (i) and (ii) of $\{h_{m,n,k}\}_{(m,(n,k))\in\mathbb{N}\times L}$ that
$\int_{T_{n}(Y_{n,k})}\!\tau_{s}^{+}\circ\bigl(T_{n}|_{Y_{n,k}}\bigr)^{-1}\,\mathrm{d}\mu\geqslant\int_{Y_{n,k}}\!\bigl(J_{n}\cdot\tau_{s}^{+}\bigr)\,\mathrm{d}\mu\quad\text{for each }s\in\mathbb{N}.$ (5.26)
For each $A\in\mathcal{B}(X)$, we define $\mu_{1,n,k}(A)\coloneqq\mu(T_{n}(A\cap Y_{n,k}))$ and $\mu_{2,n,k}(A)\coloneqq\int_{A\cap Y_{n,k}}\!J_{n}\,\mathrm{d}\mu$. Then $\mu_{1,n,k}$ and $\mu_{2,n,k}$ are finite Borel measures on $X$. By the change-of-variable formula and (5.26), we have $\bigl\langle\mu_{1,n,k},\tau_{s}^{+}\bigr\rangle\geqslant\bigl\langle\mu_{2,n,k},\tau_{s}^{+}\bigr\rangle$ for each $s\in\mathbb{N}$. By Proposition 3.31, for each $A\in\mathcal{B}(X)$, $\mu(T_{n}(A\cap Y_{n,k}))\geqslant\int_{A\cap Y_{n,k}}\!J_{n}\,\mathrm{d}\mu$. Thus, since $Y_{n}=\bigcup_{(n,k)\in L_{n}}Y_{n,k}$, we obtain that $\mu(T_{n}(A))\geqslant\int_{A}\!J_{n}\,\mathrm{d}\mu$ for each admissible set $A\subseteq Y_{n}$ for $T_{n}$. By (1.2), this implies that $\mu\in\mathcal{M}(X,T_{n};Y_{n},J_{n})$.
Conversely, consider an arbitrary $\mu\in M(X,T_{n};Y_{n},J_{n})$. Note that $T_{n}$ is a Borel isomorphism of $Y_{n,k}$ with $T_{n}(Y_{n,k})$ for each $(n,k)\in L$. Then by (5.20), (5.21), the change-of-variable formula, and property (ii) of $\{h_{m,n,k}\}_{(m,(n,k))\in\mathbb{N}\times L}$, we obtain that
$\int\!W_{m,s}^{n,k}\,\mathrm{d}\mu$ $=\int_{T_{n}(Y_{n,k})}\!\bigl(\tau_{s}^{+}\cdot h_{m,n,k}\bigr)\circ\bigl(T_{n}|_{Y_{n,k}}\bigr)^{-1}\,\mathrm{d}\mu=\int_{Y_{n,k}}\!\bigl(\tau_{s}^{+}\cdot h_{m,n,k}\bigr)\,\mathrm{d}\bigl(\mu\circ\bigl(T_{n}|_{Y_{n,k}}\bigr)\bigr)$
$\geqslant\int_{Y_{n,k}}\!\bigl(J_{n}\cdot\tau_{s}^{+}\cdot h_{m,n,k}\bigr)\,\mathrm{d}\mu=\int\!\bigl(J_{n}\cdot\tau_{s}^{+}\cdot h_{m,n,k}\bigr)\,\mathrm{d}\mu=\int\!V_{m,s}^{n,k}\,\mathrm{d}\mu.$
for all $m,\,s\in\mathbb{N}$ and $(n,k)\in L_{n}$. By (5.22), this implies that $\mu\notin\Psi_{m,s}^{n,k}$, establishing Claim 2.
By Proposition 3.9 and Claim 1, $\bigl\{\bigcup_{(m,s)\in\mathbb{N}^{2}}\bigcup_{(n,k)\in L_{n}}\Psi_{m,s}^{n,k}\bigr\}_{n\in\mathbb{N}}$ is uniformly lower semi-computable open in $(\mathcal{P}(X),\,W_{\rho},\,\mathcal{Q}_{\mathcal{S}})$. Since $X$ is recursively compact in $(X,\,\rho,\,\mathcal{S})$, by Proposition 3.26, we have that $\mathcal{P}(X)$ is recursively compact in $(\mathcal{P}(X),\,W_{\rho},\,\mathcal{Q}_{\mathcal{S}})$. By Proposition 3.20 (iii) and Claim 2, we conclude that $\{\mathcal{M}(X,T_{n};Y_{n},J_{n})\}_{n\in\mathbb{N}}$ is uniformly recursively compact in $(\mathcal{P}(X),\,W_{\rho},\,\mathcal{Q}_{\mathcal{S}})$. ∎
5.3. Applications of Theorem 1.3
In this subsection, we demonstrate how to apply Theorem 1.3 to different systems by constructing appropriate sequences $\{\mathcal{K}_{n}\}_{n\in\mathbb{N}}$, with the main applications being Theorems 5.9 and 5.10.
We state the following hypotheses under which we establish our results.
The Assumptions.
  1. (i)
    Let $(X,\,\rho,\,\mathcal{S},\,\{X_{n}\}_{n\in\mathbb{N}},\,\{T_{n}\}_{n\in\mathbb{N}})$ be a uniformly computable system. Assume that $\{Y_{n}\}_{n\in\mathbb{N}}$ is a sequence of open sets and $\{J_{n}\}$ is a sequence of Borel functions satisfying the hypotheses of Theorem 5.7.
  2. (ii)
    Let $\phi_{n}\colon X\rightarrow\mathbb{R}$ be a bounded Borel function for each $n\in\mathbb{N}$.
  3. (iii)
    For each $n\in\mathbb{N}$, properties (i) and (ii) in Theorem 5.6 are satisfied in the case where $T\coloneqq T_{n}$, $Y\coloneqq Y_{n}$, $J\coloneqq J_{n}$, and $\phi\coloneqq\phi_{n}$.
  4. (iv)
    $\{C_{n}\}_{n\in\mathbb{N}}$ is uniformly recursively compact in $(X,\,\rho,\,\mathcal{S})$ such that $\operatorname{card}\mathopen{(}C_{n}\smallsetminus Y_{n}\mathclose{)}<+\infty$ and $\mathcal{E}_{0}(T_{n},\phi_{n})\cap\mathcal{P}(X;C_{n}\cap Y_{n})=\{\mu_{n}\}$ for each $n\in\mathbb{N}$.
  5. (v)
    $H$ and $I$ are nonempty recursively enumerable sets with $H\subseteq\mathbb{N}\times I$. Moreover, $\{p_{n,i}\}_{(n,i)\in H}$ is a sequence of uniformly computable points such that $\{p_{n,i}:(n,i)\in H_{n}\}$ is the union of the set $C_{n}\smallsetminus X_{n}$ and the set $B_{n}$ of periodic points of $T_{n}$ in $C_{n}\cap(X_{n}\smallsetminus Y_{n})$ whose orbits for $T_{n}$ are contained in $C_{n}\cap X_{n}$, where $H_{n}\coloneqq\{(n,i)\in H:i\in I\}$ for each $n\in\mathbb{N}$.
We need the following lemma, whose proof is verbatim the same as that of [LS24b, Lemma 6.3 (iii)].
Lemma 5.8.
Let $(X,\rho)$ be a compact metric space, $T\colon X\rightarrow X$ be a Borel-measurable transformation, and $\mu\in\mathcal{M}(X,T)$. Assume that $x$ is a point in $X$ with $\mu(\{x\})>0$. Then $x$ is a periodic point of $T$. If we assume in addition that $\mu$ is ergodic, then $\mu=\frac{1}{n}\sum_{i=0}^{n-1}\delta_{T^{i}(x)}$, where $n$ is the period of $x$.
We now consider dynamical systems satisfying the hypotheses of Theorem 1.3 and the Assumptions. Employing Theorem 5.9, we can establish the computability of equilibrium states for some expanding Thurston maps and Hölder continuous potentials (see e.g.  Theorem 1.4).
Theorem 5.9.
Under the Assumptions, assume, in addition, that the following statement is true:
There exists a sequence $\{U_{n,i,k}\}_{((n,i),k)\in H\times\mathbb{N}}$ of uniformly lower semi-computable open sets in $(X,\,\rho,\,\mathcal{S})$ such that $p_{n,i}\in U_{n,i,k}$ for all $(n,i)\in H$ and $k\in\mathbb{N}$. Moreover, there exists a sequence $\{t_{n,i,k}\}_{((n,i),k)\in H\times\mathbb{N}}$ of uniformly computable real numbers such that for each $(n,i)\in H$,
$\inf_{k\in\mathbb{N}}t_{n,i,k}=0,\quad\text{ and}\quad\mu_{n}(U_{n,i,k})\leqslant t_{n,i,k}\quad\text{for each }k\in\mathbb{N}.$ (5.27)
Then $\{\mu_{n}\}_{n\in\mathbb{N}}$ is uniformly computable in $(\mathcal{P}(X),\,W_{\rho},\,\mathcal{Q}_{\mathcal{S}})$.
Proof.
Define for each $n\in\mathbb{N}$,
$\mathcal{K}_{n}\coloneqq\{\mu\in\mathcal{P}(X;C_{n})\cap\mathcal{M}(X,T_{n};X_{n}):\mu(U_{n,i,k})\leqslant t_{n,i,k}\text{ for all }k\in\mathbb{N}\text{ and }(n,i)\in H_{n}\}.$ (5.28)
First, we establish the uniformly recursive compactness of $\{\mathcal{K}_{n}\}_{n\in\mathbb{N}}$ as follows.
Claim 1. $\{\mathcal{K}_{n}\}_{n\in\mathbb{N}}$ is uniformly recursively compact in $(\mathcal{P}(X),\,W_{\rho},\,\mathcal{Q}_{\mathcal{S}})$.
Proof of Claim 1. Since $(X,\,\rho,\,\mathcal{S},\,\{X_{n}\}_{n\in\mathbb{N}},\,\{T_{n}\}_{n\in\mathbb{N}})$ satisfy the Assumptions, in $(X,\,\rho,\,\mathcal{S})$, $X$ is recursively compact, $\{X_{n}\}_{n\in\mathbb{N}}$ is uniformly lower semi-computable open, and $\{T_{n}\}_{n\in\mathbb{N}}$ is a sequence of uniformly computable functions with respect to $\{X_{n}\}_{n\in\mathbb{N}}$. Thus, by Proposition 3.30, $\{\mathcal{M}(X,T_{n};X_{n})\}_{n\in\mathbb{N}}$ is uniformly recursively compact in $(\mathcal{P}(X),\,W_{\rho},\,\mathcal{Q}_{\mathcal{S}})$.
Moreover, since in $(X,\,\rho,\,\mathcal{S})$, $X$ is recursively compact, and $\{C_{n}\}_{n\in\mathbb{N}}$ is uniformly recursively compact, by Proposition 3.20 (ii), $\{X\smallsetminus C_{n}\}_{n\in\mathbb{N}}$ is uniformly lower semi-computable open. Note that by definition, $\mathcal{P}(X;C_{n})=\{\mu\in\mathcal{P}(X):\mu(X\smallsetminus C_{n})\leqslant 0\}$ for each $n\in\mathbb{N}$. By Corollary 3.29, $\{\mathcal{P}(X;C_{n})\}_{n\in\mathbb{N}}$ is uniformly recursively compact in $(\mathcal{P}(X),\,W_{\rho},\,\mathcal{Q}_{\mathcal{S}})$.
Since in $(X,\,\rho,\,\mathcal{S})$, $X$ is recursively compact, $\{U_{n,i,k}\}_{((n,i),k)\in H\times\mathbb{N}}$ is uniformly lower semi-computable open, and $\{t_{n,i,k}\}_{((n,i),k)\in H\times\mathbb{N}}$ is a sequence of uniformly computable real numbers, by Corollary 3.29, $\{\{\mu\in\mathcal{P}(X):\mu(U_{n,i,k})\leqslant t_{n,i,k}\}\}_{((n,i),k)\in H\times\mathbb{N}}$ is uniformly recursively compact in $(\mathcal{P}(X),\,W_{\rho},\,\mathcal{Q}_{\mathcal{S}})$. Combined with the uniformly recursive compactness of $\{\mathcal{P}(X;C_{n})\}_{n\in\mathbb{N}}$ and $\{\mathcal{M}(X,T_{n};X_{n})\}_{n\in\mathbb{N}}$, by Proposition 3.21 and (5.28), this implies that $\{\mathcal{K}_{n}\}_{n\in\mathbb{N}}$ is uniformly recursively compact.
Next, we establish the following result.
Claim 2. $\mathcal{M}(X,T_{n};Y_{n},J_{n})\cap\mathcal{K}_{n}=\{\mu_{n}\}$ for each $n\in\mathbb{N}$.
Proof of Claim 2. We consider an arbitrary $n\in\mathbb{N}$ and prove our claim for $n$.
First, we show that
$\mathcal{M}(X,T_{n};Y_{n},J_{n})\cap\mathcal{M}(X,T_{n};X_{n})\cap\mathcal{P}(X;C_{n}\cap Y_{n})=\{\mu_{n}\}.$ (5.29)
Since $\{Y_{n}\}_{n\in\mathbb{N}}$ satisfies the hypotheses of Theorem 1.3, there exist two recursively enumerable sets K,LK,\,L with $L\subseteq\mathbb{N}\times K$, and a sequence $\{Y_{n,k}\}_{(n,k)\in L}$ of open sets such that $Y_{n,k}$ is admissible for $T_{n}$ for each $(n,k)\in L$. Since $L$ is a recursively enumerable set, there exists a function $f\colon\mathbb{N}\rightarrow L_{n}$ with $f(\mathbb{N})=L_{n}$. Define $Y_{m}^{\prime}\coloneqq Y_{f(m)}\smallsetminus\bigcup_{k=1}^{m-1}Y_{f(k)}$ for each $m\in\mathbb{N}$. Then by Definition 5.2, $(X,\,\rho,\,T_{n},\,Y_{n},\,\{Y^{\prime}_{m}\}_{m\in\mathbb{N}},\,\mu)$ is admissible. Note that $\phi_{n}$ is a bounded Borel function by item (ii) of the Assumptions. Then we have $\mathopen{\langle}\mu,\phi_{n}\mathclose{\rangle}\in\mathbb{R}$ for each $\mu\in\mathcal{P}(X)$ and each $n\in\mathbb{N}$. Hence, by item (iii) of the Assumptions and Theorem 5.6, we obtain that $\mathcal{M}(X,T_{n};Y_{n},J_{n})\cap\mathcal{M}(X,T_{n})\cap\mathcal{P}(X;C_{n}\cap Y_{n})=\mathcal{E}_{0}(T_{n},\phi_{n})\cap\mathcal{P}(X;C_{n}\cap Y_{n})$. Note that $Y_{n}\subseteq X_{n}$ by item (ii) of the Assumptions. Then by (1.1), we have $\mathcal{M}(X,T_{n};X_{n})\cap\mathcal{P}(X;C_{n}\cap Y_{n})=\mathcal{M}(X,T_{n})\cap\mathcal{P}(X;C_{n}\cap Y_{n})$. Thus, by item (iv) of the Assumptions, we obtain that $\mathcal{M}(X,T_{n};Y_{n},J_{n})\cap\mathcal{M}(X,T_{n};X_{n})\cap\mathcal{P}(X;C_{n}\cap Y_{n})=\mathcal{M}(X,T_{n};Y_{n},J_{n})\cap\mathcal{M}(X,T_{n})\cap\mathcal{P}(X;C_{n}\cap Y_{n})=\mathcal{E}_{0}(T_{n},\phi_{n})\cap\mathcal{P}(X;C_{n}\cap Y_{n})=\{\mu_{n}\}$.
Now we prove our claim by showing that these two sets are mutually inclusive. Indeed, by (5.27) and (5.28), $\mu_{n}\in\mathcal{K}_{n}$. Hence, by (5.29), we obtain that $\mu_{n}\in\mathcal{M}(X,T_{n};Y_{n},J_{n})\cap\mathcal{K}_{n}$.
Finally, we show that $\mathcal{M}(X,T_{n};Y_{n},J_{n})\cap\mathcal{K}_{n}\subseteq\{\mu_{n}\}$. We argue this for $n$ by contradiction and consider a measure $\mu\in\mathcal{M}(X,T_{n};Y_{n},J_{n})\cap\mathcal{K}_{n}$ with $\mu\neq\mu_{n}$. Then it is not hard to derive from (5.29) that $\mu\in\mathcal{P}(X;C_{n})\smallsetminus\mathcal{P}(X;C_{n}\cap Y_{n})$. Note that $\operatorname{card}\mathopen{(}C_{n}\smallsetminus Y_{n}\mathclose{)}<+\infty$ by item (iv) of the Assumptions. Thus, since $\mu\in\mathcal{P}(X;C_{n})\smallsetminus\mathcal{P}(X;C_{n}\cap Y_{n})$, we have $\mu(\{x_{0}\})>0$ for some $x_{0}\in C_{n}\smallsetminus Y_{n}$. Then we show that $\mu\in\mathcal{P}(X;X_{n})\cap\mathcal{M}(X,T_{n})$. By the assumptions in Theorem 5.9, $p_{n,i}\in U_{n,i,k}$ for all $(n,i)\in H_{n}$ and $k\in\mathbb{N}$. Hence, by $\mu\in\mathcal{K}_{n}$, it follows from (5.28) that $0\leqslant\mu(\{p_{n,i}\})\leqslant\mu(U_{n,i,k})\leqslant t_{n,i,k}$ for all $(n,i)\in H_{n}$ and $k\in\mathbb{N}$. Note that $\inf_{k\in\mathbb{N}}t_{n,i,k}=0$ for each $(n,i)\in H_{n}$ by (5.27). Then we have $\mu(\{p_{n,i}\})=0$ for each $(n,i)\in H_{n}$. Hence, by item (v) of the Assumptions, we obtain that $\mu(C_{n}\smallsetminus X_{n})=\mu(B_{n})=0$. Note that $\mu\in\mathcal{P}(X;C_{n})\cap\mathcal{M}(X,T_{n};X_{n})$ by $\mu\in\mathcal{K}_{n}$ and (5.28). Then by (1.1), we have $\mu\in\mathcal{P}(X;X_{n})\cap\mathcal{M}(X,T_{n};X_{n})=\mathcal{P}(X;X_{n})\cap\mathcal{M}(X,T_{n})$. Hence, by Lemma 5.8, it follows from $\mu(\{x_{0}\})>0$ that $x_{0}$ is a periodic point of $T_{n}$ in $C_{n}\cap(X_{n}\smallsetminus Y_{n})$. Moreover, by $\mu\in\mathcal{P}(X;C_{n}\cap X_{n})$, the orbit of $x_{0}$ for $T_{n}$ is contained in $C_{n}\cap X_{n}$. Thus, $x_{0}\in B_{n}$ by the definition of $B_{n}$. This implies that $\mu(B_{n})\geqslant\mu(\{x_{0}\})>0$, which contradicts that $\mu(B_{n})=0$. Hence, we obtain that $\mathcal{M}(X,T_{n};Y_{n},J_{n})\cap\mathcal{K}_{n}\subseteq\{\mu_{n}\}$. This completes the proof of Claim 2.
By Claims 1 and 2, $\{\mathcal{K}_{n}\}_{n\in\mathbb{N}}$ is uniformly recursively compact and satisfies (1.3). Hence, it follows from Theorem 1.3 that $\{\mu_{n}\}_{n\in\mathbb{N}}$ is uniformly computable in $(\mathcal{P}(X),\,W_{\rho},\,\mathcal{Q}_{\mathcal{S}})$. ∎
We now consider dynamical systems satisfying the hypotheses of Theorem 1.3 and the Assumptions whose maps $T_{n}$ are uniformly contracting near all periodic orbits in $B_{n}$ for $T_{n}$. Note that for a rational map, all periodic orbits containing critical points are super-attracting. Hence, Theorem 5.10 can be applied to establish the computability of equilibrium states for rational maps and potentials satisfying certain hypotheses.
Theorem 5.10.
Under the Assumptions, assume, in addition, that the following statements are true:
  1. (i)
    $C_{n}\subseteq X_{n}$ for each $n\in\mathbb{N}$.
  2. (ii)
    There exists a computable function $m\colon H\rightarrow\mathbb{N}$ and two sequences $\{r_{n,i}\}_{(n,i)\in H}$, $\{\lambda_{n,i}\}_{(n,i)\in H}$ of uniformly computable real numbers such that for each $(n,i)\in H$, $m(n,i)$ is the period of $p_{n,i}$ for $T_{n}$, and
    $\rho\bigl(p_{n,i},(T_{n})^{m(n,i)}(q)\bigr)\leqslant\lambda_{n,i}\cdot\rho(p_{n,i},q)\quad\text{for each }q\in B(p_{n,i},r_{n,i}).$
Then $\{\mu_{n}\}_{n\in\mathbb{N}}$ is a sequence of uniformly computable measures.
Proof.
We establish that $\mu_{n}(B(p_{n,i},r_{n,i}))=0$ for each $(n,i)\in H$. We consider an arbitrary pair $(n,i)\in H$ and define $B(l)\coloneqq(T_{n})^{lm(n,i)}(B(p_{n,i},r_{n,i}))$ for each $l\in\mathbb{N}_{0}$.
By item (iv) of the Assumptions, it follows from $C_{n}\subseteq X_{n}$ that $p_{n,i}\in B_{n}\subseteq C_{n}\cap(X_{n}\smallsetminus Y_{n})=C_{n}\smallsetminus Y_{n}$. Note that $\mu_{n}\in\mathcal{P}(X;C_{n}\cap Y_{n})$ by item (iv) of the Assumptions. Then we have $\mu_{n}(\{p_{n,i}\})=0$. Moreover, by the definition of $\{B(l)\}_{l\in\mathbb{N}_{0}}$, it follows from $\mu_{n}\in\mathcal{E}_{0}(T_{n},\phi_{n})\subseteq\mathcal{M}(X,T_{n})$ that $\mu_{n}(B(0))\leqslant\mu_{n}(B(l))$ for each $l\in\mathbb{N}_{0}$. Since $\lambda_{n,i}\in(0,1)$, we have $B(l)\subseteq B\bigl(p_{n,i},r_{n,i}\lambda_{n,i}^{l}\bigr)\subseteq B(0)$ for each $l\in\mathbb{N}_{0}$. Thus, we have $\mu_{n}(B(0))=\mu_{n}\bigl(B\bigl(p_{n,i},r_{n,i}\lambda_{n,i}^{l}\bigr)\bigr)=\mu_{n}(B(l))$, hence, $\mu_{n}\bigl(B(0)\smallsetminus B\bigl(p_{n,i},r_{n,i}\lambda_{n,i}^{l}\bigr)\bigr)=0$. Letting $l\rightarrow+\infty$, we obtain that $\mu_{n}(B(0)\smallsetminus\{p_{n,i}\})=0$. Combined with $\mu_{n}(\{p_{n,i}\})=0$, this proves that $\mu_{n}(B(p_{n,i},r_{n,i}))=0$.
By Proposition 3.10, it follows from the uniform computability of $\{r_{n,i}\}_{(n,i)\in H}$ and $\{p_{n,i}\}_{(n,i)\in H}$ that $\{B(p_{n,i},r_{n,i})\}_{(n,i)\in H}$ is a sequence of uniformly lower semi-computable open sets in $(X,\,\rho,\,\mathcal{S})$.
Therefore, by the above result, the hypotheses of Theorem 5.9 are satisfied with $U_{n,i,k}=B(p_{n,i},r_{n,i})$ and $t_{n,i,k}=0$ for each $(n,i)\in H$ and each $k\in\mathbb{N}$. Hence, it follows from Theorem 5.9 that $\{\mu_{n}\}_{n\in\mathbb{N}}$ is uniformly computable in $(\mathcal{P}(X),\,W_{\rho},\,\mathcal{Q}_{\mathcal{S}})$. ∎
6. Computability of equilibrium states for expanding Thurston maps
In this section, we consider a class of non-uniformly expanding maps on the topological $2$-sphere known as expanding Thurston maps. There has been active research on both thermodynamic formalism for expanding Thurston maps (see e.g. [BM10, BM17, HP09, Li18, Li15, Li17, LS24b]) and algorithmic aspects of these maps (see e.g. [SY15, RSY20]).
We first review the definition of expanding Thurston maps, along with some key concepts and results. Then in Subsection 6.2, we focus on Misiurewicz–Thurston rational maps and apply Approach I (see Theorem 1.1) to establish Theorem 1.2. Finally, in Subsection 6.3, we study computable expanding Thurston maps whose critical points are computable. By Theorem 6.4, their measure-theoretic entropy functions may not be upper semi-continuous, which prevents us from using Approach I (see Theorem 1.1) to demonstrate the computability of their corresponding equilibrium states. Instead, we apply Approach II (see Theorem 1.3) to prove Theorem 1.4.
6.1. Expanding Thurston maps
We review some key concepts and results concerning expanding Thurston maps. For a more thorough treatment of the subject, we refer the reader to the monographs [BM17, Li17].
Let $S^{2}$ denote an oriented topological $2$-sphere and $f\colon S^{2}\rightarrow S^{2}$ be a branched covering map. We denote by $\deg_{f}(x)$ the local degree of $f$ at $x\in S^{2}$. The degree of $f$ is $\deg{f}=\sum_{x\in f^{-1}(y)}\deg_{f}(x)$ for $y\in S^{2}$ and is independent of $y$.
A point $x\in S^{2}$ is a critical point of $f$ if $\deg_{f}(x)\geqslant 2$. The set of critical points of $f$ is denoted by $\operatorname{crit}{f}$. A point $y\in S^{2}$ is a postcritical point of $f$ if $y=f^{n}(x)$ for some $x\in\operatorname{crit}{f}$ and $n\in\mathbb{N}$. The set of postcritical points of $f$ is denoted by $\operatorname{post}{f}$. If $\operatorname{card}\mathopen{(}\operatorname{post}{f}\mathclose{)}<+\infty$, then $f$ is said to be postcritically-finite.
Definition 6.1 (Thurston maps).
A Thurston map is a branched covering map $f\colon S^{2}\rightarrow S^{2}$ with $\deg f\geqslant 2$ and $\operatorname{card}\mathopen{(}\operatorname{post}{f}\mathclose{)}<+\infty$.
We can now define expanding Thurston maps.
Definition 6.2 (Expanding Thurston maps).
A Thurston map $f\colon S^{2}\rightarrow S^{2}$ is called expanding if there exists a metric $d$ on $S^{2}$ that induces the standard topology on $S^{2}$ and a Jordan curve $\mathcal{C}\subseteq S^{2}$ containing $\operatorname{post}{f}$ such that
$\lim_{n\to+\infty}\sup\bigl\{\operatorname{diam}_{d}X:X\text{ is a connected component of }f^{-n}\bigl(S^{2}\smallsetminus\mathcal{C}\bigr)\bigr\}=0.$ (6.1)
For an expanding Thurston map $f$, we can fix a particular metric $d_{v}$ on $S^{2}$ called a visual metric for $f$. Such a metric induces the standard topology on $S^{2}$ ([BM17, Proposition 8.3]). For the existence of such a metric, see [BM17, Chapter 8]. For a visual metric $d_{v}$ for $f$, there exists a unique constant $\Lambda>1$ called the expansion factor of $d_{v}$ (see [BM17, Chapter 8] for more details).
We summarize the existence and the uniqueness of equilibrium states for expanding Thurston maps in the following theorem, which is part of [Li18, Theorem 1.1].
Theorem 6.3 (Li [Li18]).
Let $f\colon S^{2}\rightarrow S^{2}$ be an expanding Thurston map, and $d_{v}$ be a visual metric on $S^{2}$ for $f$. Assume that $\phi\in C^{0,\alpha}\bigl(S^{2},d_{v}\bigr)$ is a real-valued Hölder continuous function with an exponent $\alpha\in(0,1]$. Then there exists a unique equilibrium state $\mu_{\phi}$ for $f$ and $\phi$.
The main tool used in [Li18] to develop the thermodynamic formalism for expanding Thurston maps is the Ruelle operator. We recall the definition of the Ruelle operator below and refer the reader to [Li17, Chapter 3.3] for a detailed discussion.
Let $f\colon S^{2}\rightarrow S^{2}$ be an expanding Thurston map and $\phi\in C\bigl(S^{2}\bigr)$ be a real-valued continuous function. The Ruelle operator $\mathcal{L}_{\phi}$ (associated with $f$ and $\phi$) acting on $C\bigl(S^{2}\bigr)$ is given by
$\mathcal{L}_{\phi}(u)(x)\coloneqq\sum_{y\in f^{-1}(x)}\deg_{f}(y)u(y)\operatorname{exp}\mathopen{(}\phi(y)\mathclose{)}$ (6.2)
for each $u\in C\bigl(S^{2}\bigr)$. Note that $\mathcal{L}_{\phi}$ is a well-defined and continuous operator on $C\bigl(S^{2}\bigr)$.
Recall that the measure-theoretic entropy function of a continuous map $T\colon X\rightarrow X$ defined on a compact metrizable topological space $X$ is the function $\mu\mapsto h_{\mu}(T)$ defined on the space $\mathcal{M}(X,T)$ equipped with the weak topology.
The following result regarding the upper semi-continuity of the measure-theoretic entropy function is established in [LS24b, Theorem 1.1], extending [Li15, Corollary 1.3].
Theorem 6.4 (Li & Shi [LS24b]).
Let $f\colon S^{2}\rightarrow S^{2}$ be an expanding Thurston map. Then the measure-theoretic entropy function of $f$ is upper semi-continuous if and only if $f$ has no periodic critical points.
6.2. Misiurewicz–Thurston rational maps
We apply Theorem 1.1 to show the computability of equilibrium states for Misiurewicz–Thurston rational maps.
A Misiurewicz–Thurston rational map is a postcritically-finite rational map on the Riemann sphere $\widehat{\mathbb{C}}$ without periodic critical points. We remark that a postcritically-finite rational map is an expanding Thurston map (in the sense of Definition 6.2) if and only if it has no periodic critical point (see [BM17, Proposition 2.3]). To discuss the computability of rational maps, we give a computable structure of $\widehat{\mathbb{C}}$ as follows.
Proposition 6.5.
[…]
The following result states that there exists an algorithm that computes all zeros of a computable polynomial (cf. [BY09, Proposition 3.2]; see also [He25, Theorem 1.5]).
Lemma 6.6.
Let $\mathcal{S}\bigl(\widehat{\mathbb{C}}\bigr)=\{s_{n}\}_{n\in\mathbb{N}}$ be an effective enumeration of $\mathbb{Q}\bigl(\widehat{\mathbb{C}}\bigr)\coloneqq\{a+b\mathbf{i}:a,\,b\in\mathbb{Q}\}$. Then there exists an algorithm that satisfies the following property:
For all $k,\,l\in\mathbb{N}$, and complex polynomial $p$ of degree $k$, this algorithm outputs a sequence $\{q_{i}\}_{i=1}^{k}$ of integers such that there exists an enumeration $\{x_{i}\}_{i=1}^{k}$ of all the zeros of $p$ (counting with multiplicity) satisfying that $\sigma\bigl(s_{q_{i}},x_{i}\bigr)<2^{-l}$ for each integer $1\leqslant i\leqslant k$, after we input the following data into this algorithm:
  1. (i)
    the integers $k$ and $l$.
  2. (ii)
    an algorithm computing all the coefficients of the polynomial $p$.
In this subsection, we investigate computable Misiurewicz–Thurston rational maps, i.e., Misiurewicz–Thurston rational maps $f=h_{1}/h_{2}$ for which the coefficients of the polynomials $h_{1},\,h_{2}$ are all computable. Here $h_{1},\,h_{2}$ are two polynomials without common roots. We now design an algorithm to compute the Ruelle operator $\mathcal{L}_{\phi}$ (recall (6.2)).
Proposition 6.7.
There exists an algorithm that satisfies the following property:
For all $n,\,m\in\mathbb{N}$, $\phi,\,u\in C(\widehat{\mathbb{C}},\sigma)$, a Misiurewicz–Thurston rational map $f=h_{1}/h_{2}$, and $x\in\widehat{\mathbb{C}}\smallsetminus\bigl\{f^{i}(\infty):1\leqslant i\leqslant m\bigr\}$, the algorithm outputs a rational $2^{-n}$-approximation of $\mathcal{L}^{m}_{\phi}(u)(x)$, given the following input data:
  1. (i)
    an algorithm computing the function $\phi:\widehat{\mathbb{C}}\rightarrow\mathbb{R}$,
  2. (ii)
    an algorithm computing the function $u:\widehat{\mathbb{C}}\rightarrow\mathbb{R}$,
  3. (iii)
    an algorithm computing all the coefficients of the polynomials $h_{1}$ and $h_{2}$.
  4. (iv)
    an oracle $\tau:\mathbb{N}\rightarrow\mathbb{N}$ for the point $x$,
  5. (v)
    the integers $n$ and $m$.
Proof.
Since we can compute the function $\mathcal{L}_{\phi}^{m}(u)$ by iterating the operator $\mathcal{L}_{\phi}$ on the function $u$, by (6.2), it suffices to establish the algorithm in the case where $m\coloneqq 1$.
[…]
$\mathcal{L}_{\phi}(u)(x)=\sum_{y\in f^{-1}(x)}\deg_{f}(y)u(y)\exp(\phi(y))=\sum_{i=1}^{k}u(y_{i})\exp(\phi(y_{i})).$
Note that the exponential function $\exp\colon\mathbb{R}\rightarrow\mathbb{R}$ is computable by Example 3.13. Then we can use data (i) and (ii) to compute the value of $\mathcal{L}_{\phi}(u)(x)$. ∎
We now prove the computability of the topological pressure $P(f,\phi)$.
Proposition 6.8.
Let $f\colon\widehat{\mathbb{C}}\rightarrow\widehat{\mathbb{C}}$ be a computable Misiurewicz–Thurston rational map, $d_{v}$ be a visual metric on $\widehat{\mathbb{C}}$ for $f$, and $\alpha\in(0,1]$. Then there exists an algorithm that for all $n\in\mathbb{N}$ and $\phi\in C^{0,\alpha}\bigl(\widehat{\mathbb{C}},d_{v}\bigr)$, outputs a rational $2^{-n}$-approximation of $P(f,\phi)$, after inputting the following data:
  1. (i)
    an algorithm $\Phi$ computing the function $\phi$,
  2. (ii)
    a rational number $R$ with $\mathopen{\lvert}\phi\mathclose{\rvert}_{\alpha,d_{v}}\leqslant R$,
  3. (iii)
    the integer $n$.
Proof.
[…]
$|v-w|<2^{-n-1},\quad\text{where }w\coloneqq N^{-1}\log\bigl(\mathcal{L}_{\phi}^{N}\bigl(\mathbbm{1}_{\widehat{\mathbb{C}}}\bigr)(s_{m})\bigr).$
By [Li18, Lemma 5.15], there exists $C_{0}\in\mathbb{Q}$ such that
$\mathopen{\big\lvert}\log\mathopen{\big(}\mathcal{L}_{\overline{\phi}}^{n}\bigl(\mathbbm{1}_{\widehat{\mathbb{C}}}\bigr)(x)\mathclose{\big)}\mathclose{\big\rvert}\leqslant C_{0}\mathopen{\lvert}\phi\mathclose{\rvert}_{\alpha,d_{v}}\quad\text{for all }x\in\widehat{\mathbb{C}},\,n\in\mathbb{N}_{0},\text{ and }\phi\in C^{0,\alpha}\bigl(\widehat{\mathbb{C}},\sigma\bigr),$ (6.3)
where $\overline{\phi}(x)\coloneqq\phi(x)-P(f,\phi)$.
Finally, we demonstrate that the algorithm $\mathcal{M}_{C_{0}}(\Phi,R,n)$ outputs $v$ with $\mathopen{\lvert}v-P(f,\phi)\mathclose{\rvert}<2^{-n}$. Indeed, by the definition of $\mathcal{M}_{C_{0}}(\Phi,R,n)$ and (6.3), we have
$\mathopen{\lvert}w-P(f,\phi)\mathclose{\rvert}$ $=\mathopen{\big\lvert}N^{-1}\log\bigl(e^{-NP(f,\phi)}\mathcal{L}_{\phi}^{N}\bigl(\mathbbm{1}_{\widehat{\mathbb{C}}}\bigr)(x_{0})\bigr)\mathclose{\big\rvert}<N^{-1}\mathopen{\big\lvert}\log\mathopen{\big(}\mathcal{L}_{\overline{\phi}}^{N}\bigl(\mathbbm{1}_{\widehat{\mathbb{C}}}\bigr)(x_{0})\mathclose{\big)}\mathclose{\big\rvert}\leqslant N^{-1}C_{0}R<2^{-n-1}.$
Hence, by $\mathopen{\lvert}v-w\mathclose{\rvert}<2^{-n-1}$, this implies that $\mathopen{\lvert}v-P(f,\phi)\mathclose{\rvert}<2^{-n}$. ∎
Now we can apply Theorem 1.1 to prove Theorem 1.2.
Proof of Theorem 1.2.
Since there exists no periodic critical point of the Misiurewicz–Thurston rational map $f$, by Theorem 6.4, the measure-theoretic entropy function $\nu\mapsto h_{\nu}(f)$ is upper semi-continuous. By the hypothesis of Theorem 1.2, we have $\mathcal{E}(f,\phi_{n})=\{\mu_{n}\}$ for each $n\in\mathbb{N}$. Hence, to prove Theorem 1.2, by Theorem 1.1, it suffices to verify that the sequence $\{\phi_{n}\}_{n\in\mathbb{N}}$ of functions $\phi_{n}\colon\widehat{\mathbb{C}}\rightarrow\mathbb{R}$ satisfies properties (i) and (ii) in Theorem 1.1 in the case where $T_{n}\coloneqq f$ for each $n\in\mathbb{N}$.
We first apply Proposition 6.8 to show the uniform computability of $\{P(f,\phi_{n})\}_{n\in\mathbb{N}}$. Since $f$ is a Misiurewicz–Thurston rational map, by [BM17, Lemma 18.10], the identity map $\operatorname{id}_{\widehat{\mathbb{C}}}\colon\bigl(\widehat{\mathbb{C}},d_{v}\bigr)\rightarrow\bigl(\widehat{\mathbb{C}},\sigma\bigr)$ is a quasisymmetric homeomorphism. Note that $\bigl(\widehat{\mathbb{C}},\sigma\bigr)$ is a uniformly perfect space. By [Hei01, Corollary 11.5], the identity map $\operatorname{id}_{\widehat{\mathbb{C}}}$ is Hölder continuous on bounded sets. Hence, by the compactness of $\bigl(\widehat{\mathbb{C}},\sigma\bigr)$, there exist two constants $\beta\in(0,1]$ and $C\in\mathbb{Q}^{+}$ satisfying that
$\sigma(x,y)\leqslant Cd_{v}(x,y)^{\beta}\quad\text{ for all }x,y\in\widehat{\mathbb{C}}$. (6.4)
Thus, for each $n\in\mathbb{N}$, by $\phi_{n}\in C^{0,\alpha}\bigl(\widehat{\mathbb{C}},\sigma\bigr)$ and $\mathopen{\lvert}\phi_{n}\mathclose{\rvert}_{\alpha,\sigma}\leqslant Q_{n}$, we obtain that $\phi_{n}\in C^{0,\alpha\beta}\bigl(\widehat{\mathbb{C}},d_{v}\bigr)$ and $\mathopen{\lvert}\phi_{n}\mathclose{\rvert}_{\alpha\beta,d_{v}}\leqslant C^{\alpha}\mathopen{\lvert}\phi_{n}\mathclose{\rvert}_{\alpha,\sigma}\leqslant C^{\alpha}Q_{n}$. Hence, by Proposition 6.8, $\{P(f,\phi_{n})\}_{n\in\mathbb{N}}$ is a sequence of uniformly computable real numbers, i.e., $\{\phi_{n}\}_{n\in\mathbb{N}}$ satisfies property (ii) in Theorem 1.1.
[…]
$\psi_{k}=\sum_{t=1}^{l_{k}}q_{k,t}\prod_{j=1}^{m_{k,t}}f_{i_{j,k,t}},\quad\text{ where }l_{k},\,m_{k,t},\,i_{j,k,t}\in\mathbb{N}_{0}\text{ and }q_{k,t}\in\mathbb{Q}\text{ for all }j,\,k,\,t\in\mathbb{N}.$ (6.5)
Thus, since $\{f_{i}\}_{i\in\mathbb{N}_{0}}$ is a sequence of uniformly computable functions, so is the sequence $\{\psi_{k}\}_{k\in\mathbb{N}}$.
Claim. $\{P(f,\psi_{k})\}_{k\in\mathbb{N}}$ is a sequence of uniformly computable real numbers.
Proof of the claim. By the definition of the chordal metric $\sigma$ and the function $f_{i}$, we have $f_{i}(x)=\sigma(x,s_{i})\leqslant 2$ for all $i\in\mathbb{N}_{0}$ and $x\in\widehat{\mathbb{C}}$. Thus, since $f_{i}\in C^{0,1}\bigl(\widehat{\mathbb{C}},\sigma\bigr)$ and $\mathopen{\lvert}f_{i}\mathclose{\rvert}_{1,\sigma}=1$ for each $i\in\mathbb{N}_{0}$, we obtain that for all $k,\,t\in\mathbb{N}$ and $x,\,y\in\widehat{\mathbb{C}}$,
$\mathopen{\Bigg\lvert}\prod_{j=1}^{m_{k,t}}f_{i_{j,k,t}}(x)-\prod_{j=1}^{m_{k,t}}f_{i_{j,k,t}}(y)\mathclose{\Bigg\rvert}$ $=\mathopen{\Bigg\lvert}\sum_{j=1}^{m_{k,t}}\Biggl(\prod_{p=1}^{j-1}f_{i_{p,k,t}}(x)\cdot\prod_{q=j+1}^{m_{k,t}}f_{i_{q,k,t}}(y)\cdot\bigl(f_{i_{j,k,t}}(x)-f_{i_{j,k,t}}(y)\bigr)\Biggr)\mathclose{\Bigg\rvert}$
[…]
$\leqslant 2^{m_{k,t}-1}m_{k,t}\cdot\sigma(x,y).$
By (6.5), $\mathopen{\lvert}\psi_{k}\mathclose{\rvert}_{1,\sigma}\leqslant\sum_{t=1}^{l_{k}}2^{m_{k,t}-1}m_{k,t}\mathopen{\big\lvert}q_{k,t}\mathclose{\big\rvert}$ for each $k\in\mathbb{N}$. Hence, by Definition 3.12, the function $F\colon\mathbb{N}\rightarrow\mathbb{R}$ given by $F(k)\coloneqq\sum_{t=1}^{l_{k}}2^{m_{k,t}-1}m_{k,t}\mathopen{\big\lvert}q_{k,t}\mathclose{\big\rvert}$ for each $k\in\mathbb{N}$ is a computable function satisfying $F(k)\geqslant\mathopen{\lvert}\psi_{k}\mathclose{\rvert}_{1,\sigma}$ for each $k\in\mathbb{N}$. Thus, by (6.4), we have $\psi_{k}\in C^{0,\beta}\bigl(\widehat{\mathbb{C}},d_{v}\bigr)$ and $\mathopen{\lvert}\psi_{k}\mathclose{\rvert}_{\beta,d_{v}}\leqslant C\mathopen{\lvert}\psi_{k}\mathclose{\rvert}_{1,\sigma}\leqslant CF(k)$ for each $k\in\mathbb{N}$. By Proposition 6.8, $\{P(f,\psi_{k})\}_{k\in\mathbb{N}}$ is a sequence of uniformly computable real numbers. This completes the proof of the claim.
By the above claim, the sequence $\{\phi_{n}\}_{n\in\mathbb{N}}$ satisfies property (i) in Theorem 1.1. Therefore, by Theorem 1.1, $\{\mu_{n}\}_{n\in\mathbb{N}}$ is a sequence of uniformly computable measures. ∎
6.3. Computable expanding Thurston maps
[…]
Before the proof of Theorem 1.4, we establish the following results.
Lemma 6.9.
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space in which $X$ is recursively compact. Suppose that all balls in $(X,\rho)$ are connected. Assume that $I$ is a recursively enumerable set and that $\{U_{i}\}_{i\in I}$ is uniformly lower semi-computable open. Then there exists a recursively enumerable set $E\subseteq\mathbb{N}\times I$ and a sequence $\{V_{n,i}\}_{(n,i)\in E}$ of uniformly lower semi-computable open sets that are connected such that $U_{i}=\bigcup_{(n,i)\in E_{i}}V_{n,i}$, where $E_{i}\coloneqq\{(n,i)\in E:n\in\mathbb{N}\}$ for each $i\in I$.
Proof.
Without loss of generality, we assume that $\bigcup_{i\in I}U_{i}\neq\emptyset$. Write $\mathcal{S}=\{s_{n}\}_{n\in\mathbb{N}}$. Since $\{U_{i}\}_{i\in I}$ is uniformly lower semi-computable open, by Proposition 3.8, there exists a nonempty recursively enumerable set $E\subseteq\mathbb{N}\times I$ such that $\{s_{n}:(n,i)\in E_{i}\}=\{s_{n}:n\in\mathbb{N}\}\cap U_{i}$, where $E_{i}=\{(n,i)\in E:n\in\mathbb{N}\}$ for each $i\in I$. We define $V_{n,i}\coloneqq B\bigl(s_{n},\rho\bigl(s_{n},U_{i}^{c}\bigr)\bigr)$ for each $(n,i)\in E$. Since all open balls in $(X,\rho)$ are connected, $V_{n,i}$ is connected for each $(n,i)\in E$.
Now we fix $i\in I$ and check that $U_{i}=\bigcup_{(n,i)\in E_{i}}V_{n,i}$. Indeed, for each $(n,i)\in E_{i}$, since $B\bigl(s_{n},\rho\bigl(s_{n},U_{i}^{c}\bigr)\bigr)$ is connected and $s_{n}\in U_{i}$, we have $B\bigl(s_{n},\rho\bigl(s_{n},U_{i}^{c}\bigr)\bigr)\subseteq U_{i}$. By the definition of $\{V_{n,i}\}_{(n,i)\in E}$, this implies that $\bigcup_{(n,i)\in E}V_{n,i}\subseteq U_{i}$. Then we establish $U_{i}\subseteq\bigcup_{(n,i)\in E_{i}}V_{n,i}$. We consider an arbitrary $x_{0}\in U_{i}$, and show that $x_{0}\in V_{n_{0},i}$ for some $n_{0}\in\mathbb{N}$ with $(n_{0},i)\in E_{i}$. Since $\{s_{n}:n\in\mathbb{N}\}$ is dense in $X$ and $\{s_{n}:(n,i)\in E_{i}\}=\{s_{n}:n\in\mathbb{N}\}\cap U_{i}$, there exists $n_{0}\in\mathbb{N}$ with $(n_{0},i)\in E_{i}$ such that $\rho\bigl(s_{n_{0}},x_{0}\bigr)<\rho\bigl(x_{0},U_{i}^{c}\bigr)/2$. Thus, we have $x_{0}\in B\bigl(s_{n_{0}},\rho\bigl(x_{0},U_{i}^{c}\bigr)\big/2\bigr)\subseteq B\bigl(s_{n_{0}},\rho\bigl(x_{0},U_{i}^{c}\bigr)-\rho\bigl(x_{0},s_{n_{0}}\bigr)\bigr)\subseteq B\bigl(s_{n_{0}},\rho\bigl(s_{n_{0}},U_{i}^{c}\bigr)\bigr)=V_{n_{0},i}$. Therefore, we obtain that $U_{i}=\bigcup_{(n,i)\in E_{i}}V_{n,i}$ for each $i\in I$.
Finally, we show that $\{V_{n,i}\}_{(n,i)\in E}$ is uniformly lower semi-computable open. Indeed, since $X$ is recursively compact, by Proposition 3.20 (iii), $\bigl\{U_{i}^{c}\bigr\}_{i\in I}$ is uniformly recursively compact. Note that $E$ is a nonempty recursively enumerable set. Then by Proposition 3.20 (iv), the sequence $\bigl\{\rho\bigl(s_{n},U_{i}^{c}\bigr)\bigr\}_{(n,i)\in E}$ is uniformly lower semi-computable. Note that for each pair $(n,i)\in E$, $V_{n,i}=B\bigl(s_{n},\rho\bigl(s_{n},U_{i}^{c}\bigr)\bigr)$ by definition. Then by Proposition 3.10, $\{V_{n,i}\}_{(n,i)\in E}$ is uniformly lower semi-computable open. ∎
In order to prove Theorem 1.4, we need some notions. Let $f\colon\widehat{\mathbb{C}}\rightarrow\widehat{\mathbb{C}}$ be an expanding Thurston map and $\mathcal{C}\subseteq\widehat{\mathbb{C}}$ be a Jordan curve containing $\operatorname{post}{f}$. For $n\in\mathbb{N}_{0}$, we define the set of $n$-tiles as
$\mathbf{X}^{n}(f,\mathcal{C})\coloneq\bigl\{\ \overline{A}\,:\,A\text{ is a connected component of }\widehat{\mathbb{C}}\smallsetminus f^{-n}(\mathcal{C})\bigr\}$
and define the set of $n$-edges as
$\mathbf{E}^{n}(f,\mathcal{C})\coloneq\bigl\{\ \overline{A}\,:\,A\text{ is a connected component of }f^{-n}(\mathcal{C})\smallsetminus f^{-n}(\operatorname{post}{f})\bigr\}.$
For $v\in\operatorname{crit}{f^{n}}$, let $W^{n}(v)$ denote the $n$-flower of $v$, as defined in [BM17, Definition 5.27]. By [BM17, Lemmas 5.28 (i) and 5.9 (v)], this set is explicitly given by
$W^{n}(v)=\mathopen{\{}v\mathclose{\}}\cup\bigcup\mathopen{\{}\operatorname{int}X:X\in\mathbf{X}^{n}(f,\mathcal{C}),\,v\in X\mathclose{\}}\cup\bigcup\mathopen{\{}\operatorname{int}e:e\in\mathbf{E}^{n}(f,\mathcal{C}),\,v\in e\mathclose{\}}.$
Note that flowers are open (in the standard topology on $\widehat{\mathbb{C}}$). Let $\overline{W}^{n}(v)$ be the closure of $W^{n}(v)$.
Remark 6.10.
For each $n\in\mathbb{N}_{0}$ and each $v\in\operatorname{crit}{f^{n}}$, we have $\overline{W}^{n}(v)=X_{1}\cup X_{2}\cup\cdots\cup X_{m}$, where $m\coloneq 2\deg_{f^{n}}(v)$, and $X_{1},\,X_{2},\,\dots,\,X_{m}$ are all the $n$-tiles that contain $v$ as a vertex (see [BM17, Lemma 5.28]). In particular, $W^{n}(v)$ is a simply connected open set. Moreover, if $f(v)=v$ and $v\in\operatorname{crit}{f^{n-1}}$, then it follows from [BM17, Proposition 5.16] that
$f(W^{n}(v))=W^{n-1}(v).$ (6.6)
Proof of Theorem 1.4.
[…]
Now we show that there exists $k\in\mathbb{N}$ such that $f^{km}(x)=x$ for all $m\in\mathbb{N}$ and $x\in\mathrm{Per}\bigl(f^{km}\bigr)\cap\mathrm{post}\bigl(f^{km}\bigr)$. Indeed, by $\operatorname{card}\mathopen{(}\mathrm{Per}(f)\cap\operatorname{post}{f}\mathclose{)}<+\infty$, there exists $k\in\mathbb{N}$ satisfying that $f^{k}(x)=x$ for each $x\in\mathrm{Per}(f)\cap\operatorname{post}{f}$. Thus, since $\mathrm{Per}(f^{n})\cap\mathrm{post}(f^{n})\subseteq\mathrm{Per}(f)\cap\operatorname{post}{f}$ for each $n\in\mathbb{N}$, we have $f^{km}(x)=x$ for all $m\in\mathbb{N}$ and $x\in\mathrm{Per}\bigl(f^{km}\bigr)\cap\operatorname{post}{f^{km}}$.
Furthermore, by [BM17, Theorem 15.1] we can find $m\in\mathbb{N}$ such that there exists an $(f^{m})$-invariant Jordan curve $\mathcal{C}\subseteq S^{2}$ with $\mathrm{post}\bigl(f^{km}\bigr)=\operatorname{post}{f}\subseteq\mathcal{C}$. We set $F\coloneqq f^{km}$ and fix such a Jordan curve $\mathcal{C}$. Then by the previous discussion, we obtain that $F(\mathcal{C})\subseteq\mathcal{C}$ and all points in $\mathrm{Per}(F)\cap\operatorname{post}{F}$ are fixed points of $F$.
[…]
We begin by constructing a recursively enumerable set $K$ and a sequence $\{Y_{n,k}\}_{(n,k)\in L}$ of uniformly lower semi-computable open sets, where $L=\mathbb{N}\times K$. Now we write $\mathcal{S}\bigl(\widehat{\mathbb{C}}\bigr)=\{s_{i}\}_{i\in\mathbb{N}}$. Since $\operatorname{post}{F}$ is a finite set of computable points, by Proposition 3.20 (i), (ii), and (v), $\widehat{\mathbb{C}}\smallsetminus\operatorname{post}{F}$ is a nonempty lower semi-computable open set. By Proposition 3.8, there exists a recursively enumerable set $I\subseteq\mathbb{N}$ such that $\mathbb{Q}\bigl(\widehat{\mathbb{C}}\bigr)\smallsetminus\operatorname{post}{F}=\{s_{i}:i\in I\}$. We define $r_{i}\coloneqq\sigma(s_{i},\operatorname{post}{F})>0$ for each $i\in I$. Thus, by Definition 3.3, $\{r_{i}\}_{i\in I}$ is uniformly computable. Consequently, by Proposition 3.10, $\{D_{i}\}_{i\in I}$ is uniformly lower semi-computable open, where $D_{i}=B(s_{i},r_{i})$ for each $i\in I$. As $F$ is computable, $\bigl\{F^{-1}(D_{i})\bigr\}_{i\in I}$ is also uniformly lower semi-computable. By Lemma 6.9, there exists a recursively enumerable set $K\subseteq\mathbb{N}\times I$ and a sequence $\{V_{k}\}_{k\in K}$ of uniformly lower semi-computable open sets that are connected such that $\bigcup_{k\in K_{i}}V_{k}=F^{-1}(D_{i})$ for each $i\in I$.
Define $Y_{n,k}\coloneqq V_{k}$ for all $n\in\mathbb{N}$ and $k\in K$. Then we check that $Y_{n,k}$ is admissible for $T_{n}$, and $Y_{n,k}\subseteq X_{n}$ for each $(n,k)\in L$. By definitions, it suffices to show that $\bigcup_{k\in K}V_{k}=\widehat{\mathbb{C}}\smallsetminus F^{-1}(\operatorname{post}{F})$ and $V_{k}$ is admissible for $F$ for each $k\in\mathbb{N}$.
First, we verify that $\bigcup_{k\in K}V_{k}=\widehat{\mathbb{C}}\smallsetminus F^{-1}(\operatorname{post}{F})$. Indeed, by $\{s_{i}:i\in I\}=\mathbb{Q}\bigl(\widehat{\mathbb{C}}\bigr)\smallsetminus\operatorname{post}{F}$, $\operatorname{card}\mathopen{(}\operatorname{post}{F}\mathclose{)}<+\infty$, and the definition of $\{r_{i}\}_{i\in I}$, we obtain that $\bigcup_{i\in I}D_{i}=\widehat{\mathbb{C}}\smallsetminus\operatorname{post}{F}$. Hence, we have $\bigcup_{k\in K}V_{k}=\bigcup_{i\in I}\bigcup_{k\in K_{i}}V_{k}=\bigcup_{i\in I}F^{-1}(D_{i})=F^{-1}\bigl(\bigcup_{i\in I}D_{i}\bigr)=\widehat{\mathbb{C}}\smallsetminus F^{-1}(\operatorname{post}{F})$.
Next, we verify that $V_{k}$ is admissible for $F$ for each $k\in K$. Now we consider an arbitrary $k=(n,i)\in K$. Since $V_{k}$ is an open set and $F\colon\widehat{\mathbb{C}}\rightarrow\widehat{\mathbb{C}}$ is a Borel-measurable function, by [Ke95, Corollary 15.2], it suffices to show that $F$ is injective on $V_{k}$. Since $F$ is an expanding Thurston map, by [BM17, Lemma A.11], we obtain that the restriction $F|_{\widehat{\mathbb{C}}\smallsetminus F^{-1}(\operatorname{post}{F})}\colon\widehat{\mathbb{C}}\smallsetminus F^{-1}(\operatorname{post}{F})\rightarrow\widehat{\mathbb{C}}\smallsetminus\operatorname{post}{F}$ is a covering map.
Now we demonstrate that the restriction $F|_{F^{-1}(D_{i})}$ is a covering map onto $D_{i}$. We fix an arbitrary $x_{0}\in D_{i}$. Since $F$ is a finite map, $F^{-1}(x_{0})$ is a finite set, say $\{z_{n}:n\in[1,b]\cap\mathbb{N}\}$. Note that $D_{i}\subseteq\widehat{\mathbb{C}}\smallsetminus\operatorname{post}{F}$ and $D_{i}$ is path-connected and locally path-connected. Then by [BM17, Lemma A.6 (ii)], for each integer $1\leqslant n\leqslant b$, there exists a continuous map $h_{n}\colon D_{i}\rightarrow\widehat{\mathbb{C}}\smallsetminus F^{-1}(\operatorname{post}{F})$ such that $h_{n}(x_{0})=z_{n}$ and $F\circ h_{n}=\mathrm{id}_{D_{i}}$.
We argue by contradiction that for each pair of distinct integers $1\leqslant l,\,l^{\prime}\leqslant b$,
$h_{l}(D_{i})\cap h_{l^{\prime}}(D_{i})=\emptyset.$ (6.7)
We assume that (6.7) is not true for some pair of distinct integers $1\leqslant l,\,l^{\prime}\leqslant b$. Then there exist $z,\,z^{\prime}\in D_{i}$ with $h_{l}(z)=h_{l^{\prime}}(z^{\prime})$. Hence, $z=F(h_{l}(z))=F(h_{l^{\prime}}(z^{\prime}))=z^{\prime}$. Combined with [BM17, Lemma A.6 (i)], this implies that $h_{l}=h_{l^{\prime}}$, which contradicts $l\neq l^{\prime}$. Thus, (6.7) is true for each pair of distinct integers $1\leqslant l,\,l^{\prime}\leqslant b$.
Then we show that
$\bigcup_{n=1}^{b}h_{n}(D_{i})=F^{-1}(D_{i}).$ (6.8)
Indeed, since $F\circ h_{n}=\mathrm{id}_{D_{i}}$ for all integer $1\leqslant n\leqslant b$ and $x\in D_{i}$, we have $h_{n}(D_{i})\subseteq F^{-1}(D_{i})$ for each integer $1\leqslant n\leqslant b$. Hence, it suffices to show that $F^{-1}(D_{i})\subseteq\bigcup_{n=1}^{b}h_{n}(D_{i})$. We consider an arbitrary $y^{\prime}\in F^{-1}(D_{i})$. Then by [BM17, Lemma A.6 (ii)], there exists $h\colon D_{i}\rightarrow\widehat{\mathbb{C}}\smallsetminus F^{-1}(\operatorname{post}{F})$ such that $h(F(y^{\prime}))=y^{\prime}$ and $F\circ h=\mathrm{id}_{D_{i}}$. Note that $F^{-1}(x_{0})=\{z_{n}:n\in[1,b]\cap\mathbb{N}\}$ and $F(h(x_{0}))=x_{0}$. Then there exists an integer $n_{0}\in[1,b]$ satisfying that $h_{n_{0}}(x_{0})=z_{n_{0}}=h(x_{0})$. Combined with [BM17, Lemma A.6 (i)], this implies that $h_{n_{0}}=h$. Thus, we have $y^{\prime}=h(F(y^{\prime}))\subseteq h(D_{i})=h_{n_{0}}(D_{i})\subseteq\bigcup_{n=1}^{b}h_{n}(D_{i})$. Hence, we complete the proof of (6.8).
Since $h_{n}$ is continuous for each integer $1\leqslant n\leqslant b$, by (6.7) and (6.8), we have shown that the restriction $F|_{F^{-1}(D_{i})}$ is a covering map onto $D_{i}$.
Recall that $D_{i}$ is an open ball in $\widehat{\mathbb{C}}$. Hence, since $\widehat{\mathbb{C}}$ is locally path-connected, all connected components of $F^{-1}(D_{i})$ are always path-connected. Thus, since $D_{i}$ is a simply connected open set, for each connected component $V$ of $F^{-1}(D_{i})$, the restriction $F|_{V}$ is a homeomorphism onto $D_{i}$. Note that $V_{k}$ is connected and $V_{k}\subseteq F^{-1}(D_{i})$. Then there exists a connected component $V$ of $F^{-1}(D_{i})$ with $V\supseteq V_{k}$. Thus, $F$ is injective on $V_{k}$.
So far we have shown Claim 1.
We can enumerate $\mathrm{Per}(F)\cap\operatorname{post}{F}$ by $\{y_{i}\}_{i=1}^{N}$ as a sequence of uniformly computable points that are different from each other, since it is a finite set of computable points.
We now construct a sequence $\{\mathbb{B}_{i}\}_{i=1}^{N}$ of uniformly lower semi-computable open sets satisfying that $\{y_{i}\}\subseteq\mathbb{B}_{i}\subseteq W^{1}(y_{i})$ for each integer $1\leqslant i\leqslant N$. Fix an arbitrary integer $1\leqslant i\leqslant N$. By Remark 6.10, $W^{1}(y_{i})$ is an open set containing the point $y_{i}$. Then there exists $R_{i}\in\mathbb{Q}^{+}$ with $B(y_{i},R_{i})\subseteq W^{1}(y_{i})$. Since $y_{i}$ is a computable point, by Proposition 3.10, $\mathbb{B}_{i}\coloneqq B(y_{i},R_{i})$ is a lower semi-computable open set with $\{y_{i}\}\subseteq\mathbb{B}_{i}\subseteq W^{1}(y_{i})$. Hence, by Definition 3.7, $\{\mathbb{B}_{i}\}_{i=1}^{N}$ is uniformly lower semi-computable open.
[…]
Proof of Claim 2. Now we verify Assumptions (ii) to (v) in Subsection 5.3.
Assumption (ii): Since $\phi_{0}$ is the constant function, this assumption holds.
Assumption (iv): Note that $\widehat{\mathbb{C}}$ is recursively compact in $(X,\,\rho,\,\mathcal{S})$ by Proposition 6.5. Hence, by definitions, it suffices to show that $\operatorname{card}\mathopen{\big(}F^{-1}(\operatorname{post}{F})\mathclose{\big)}<+\infty$ and $\mathcal{E}_{0}(F,\phi_{0})\cap\mathcal{P}\bigl(X;\widehat{\mathbb{C}}\smallsetminus F^{-1}(\operatorname{post}{F})\bigr)=\{\mu_{0}\}$. Moreover, by Theorem 6.3 and [BM17, Corollary 17.3], we have $\mathcal{E}(F,\phi_{0})=\mathcal{E}(f,\phi_{0})=\{\mu_{0}\}$. Then we show that $\mu_{0}(\operatorname{post}{F})=0$. We argue by contradiction and assume that $\mu_{0}(\operatorname{post}{F})>0$. Combined with the fact that $\operatorname{card}\mathopen{(}\operatorname{post}{F}\mathclose{)}<+\infty$, this implies that $\mu_{0}(\{l_{0}\})>0$ for some $l_{0}\in\operatorname{post}{F}$. On the other hand, by [BM17, Theorem 17.1], the measure $\mu_{0}$ is mixing. Since mixing implies ergodicity, it follows that $\mu_{0}$ is ergodic. Hence, by [LS24b, Lemma 6.3 (iii)], it follows from $\mu_{0}(\{l_{0}\})>0$ that $l_{0}$ is a periodic point of $F$ and $\mu_{0}=\frac{1}{m}\sum_{k=0}^{m-1}\delta_{F^{k}(l_{0})}$ for some $m\in\mathbb{N}$. Thus, it follows from (3.1) that $h_{\mu_{0}}(F)=0$. However, since $\mu_{0}\in\mathcal{E}(F,\phi_{0})$, $h_{\mu_{0}}(F)=h_{\mathrm{top}}(F)=\log(\deg{F})>0$. This contradiction shows that $\mu_{0}(\operatorname{post}{F})=0$. Hence, by $\mu_{0}\in\mathcal{M}\bigl(\widehat{\mathbb{C}},F\bigr)$, we have $\mu_{0}\bigl(F^{-1}(\operatorname{post}{F})\bigr)=\mu_{0}(\operatorname{post}{F})=0$. Thus, we obtain that $\mathcal{E}_{0}(F,\phi_{0})\cap\mathcal{P}\bigl(X;\widehat{\mathbb{C}}\smallsetminus F^{-1}(\operatorname{post}{F})\bigr)=\{\mu_{0}\}$.
Assumption (iii): Since $F$ is an expanding Thurston map, by [BM17, Corollary 17.2], we have $h_{\mathrm{top}}(F)=\log(\deg{F})$. Suppose $h=\phi_{0}$. Then property (i) in Theorem 5.6 is satisfied in the case where $T\coloneqq F,\,Y\coloneqq\widehat{\mathbb{C}}\smallsetminus F^{-1}(\operatorname{post}{F}),\,J\coloneqq\deg{F}\cdot\mathbbm{1}_{\widehat{\mathbb{C}}}$, and $\phi\coloneqq\phi_{0}$. Moreover, since $F$ is an expanding Thurston map, we have $\operatorname{card}\mathopen{\big(}F^{-1}(x)\smallsetminus F^{-1}(\operatorname{post}{F})\bigr)\mathclose{\big)}=\deg{F}$ for each $x\in F\bigl(\widehat{\mathbb{C}}\smallsetminus F^{-1}(\operatorname{post}{F})\bigr)$. Hence, property (ii) in Theorem 5.6 is satisfied in the case where $T\coloneqq F,\,Y\coloneqq\widehat{\mathbb{C}}\smallsetminus F^{-1}(\operatorname{post}{F})$, and $J\coloneqq\deg{F}\cdot\mathbbm{1}_{\widehat{\mathbb{C}}}$. This verifies Assumption (iii).
Assumption (v): Now we demonstrate that $\mathrm{Per}(F)\cap\operatorname{post}{F}$ is the set of periodic points of $F$ in $F^{-1}(\operatorname{post}{F})$, namely, $\mathrm{Per}(F)\cap\operatorname{post}{F}=\mathrm{Per}(F)\cap F^{-1}(\operatorname{post}{F})$. Indeed, this immediately follows from the property that $F(x)=x$ for each $x\in\mathrm{Per}(F)\cap\operatorname{post}{F}$. Since $\{y_{i}\}_{i\in I}$ is uniformly computable, $\{p_{n,i}\}_{(n,i)\in H}$, with $p_{n,i}=y_{i}$ for all $(n,i)\in H$, is also uniformly computable. Moreover, by the definition of $\{y_{i}\}_{i=1}^{N}$, we have that $\{p_{n,i}:(n,i)\in H_{n}\}=\{y_{i}:i\in I\}=\mathrm{Per}(F)\cap\operatorname{post}{F}$. Therefore, Claim 2 follows.
Claim 3. The additional statement in Theorem 5.9 is true in the case where $(X,\,\rho,\,\mathcal{S})\coloneqq\bigl(\widehat{\mathbb{C}},\,\sigma,\,\mathcal{S}\bigl(\widehat{\mathbb{C}}\bigr)\bigr),\,I\coloneqq\mathbb{N}\cap[1,N],\,H\coloneqq\mathbb{N}\times I,\,\mu_{n}\coloneqq\mu_{0}$, and $p_{n,i}\coloneqq y_{i}$ for each $(n,i)\in H$.
Proof of Claim 3. Recall that $\{\mathbb{B}_{i}\}_{i=1}^{N}$ is uniformly lower semi-computable open and $F$ is a computable function. Define $U_{n,i,1}\coloneqq\mathbb{B}_{i}$ and $U_{n,i,k+1}\coloneqq F^{-1}(U_{n,i,k})\cap\mathbb{B}_{i}$ by induction for each $n,\,k\in\mathbb{N}$, and each $i\in I$. Then by Corollary 3.22, we have $\{U_{n,i,k}\}_{((n,i),k)\in H\times\mathbb{N}}$ is uniformly lower semi-computable open. Recall that $y_{i}$ is a fixed point of $F$ and $y_{i}\in\mathbb{B}_{i}$ for each $i\in I$. Then we obtain that $y_{i}\in U_{n,i,k}$ for all $n,\,k\in\mathbb{N}$, and $i\in I$.
Next, we set $t_{n,i,k}=\bigl(\deg_{F}(y_{i})\big/\deg{F}\bigr)^{k-1}$ for each $n,\,k\in\mathbb{N}$, and each $i\in I$. Then by [Li17, Lemma 4.27], $\deg_{F}(y_{i})<\deg{F}$ for each $i\in I$. Thus, for all $n\in\mathbb{N}$ and $i\in I$, we obtain that $\inf_{k\in\mathbb{N}}t_{n,i,k}=\inf_{k\in\mathbb{N}}\bigl(\deg_{F}(y_{i})\big/\deg{F}\bigr)^{k}=0$. Moreover, since $\deg_{F}(y_{i})\big/\deg{F}\in\mathbb{Q}$ for each $i\in I$, we have that $\{t_{n,i,k}\}_{((n,i),k)\in H\times\mathbb{N}}$ is a sequence of uniformly computable real numbers.
Finally, we consider an arbitrary pair of $n\in\mathbb{N}$ and $i\in I$ and prove that $\mu_{0}(U_{n,i,k})\leqslant t_{n,i,k}$ for each $k\in\mathbb{N}$. Recall that $y_{i}\in\operatorname{post}{F}$ and $F(y_{i})=y_{i}$. For each $m\in\mathbb{N}$, we define the set $\mathbf{\mathfrak{X}}^{m}(F,\mathcal{C},y_{i})$ of $m$-tiles of $F$ at $y_{i}$ as $\mathbf{\mathfrak{X}}^{m}(F,\mathcal{C},y_{i})\coloneq\{X\in\mathbf{X}^{m}(F,\mathcal{C}):y_{i}\in X\}$, then by Remark 6.10, we obtain that $\overline{W}^{m}(y_{i})=\bigcup\mathbf{\mathfrak{X}}^{m}(F,\mathcal{C},y_{i})$ and $\operatorname{card}\mathopen{(}\mathbf{\mathfrak{X}}^{m}(F,\mathcal{C},y_{i})\mathclose{)}=2\mathopen{\big(}\deg_{F}(y_{i})\mathclose{\big)}^{m}$. In particular, (6.6) implies that $F\bigl(W^{m+1}(y_{i})\bigr)=W^{m}(y_{i})$ for each $m\in\mathbb{N}$. Employing [BM17, Proposition 17.12] (recall that the Jordan curve $\mathcal{C}$ is $F$-invariant) and the fact that $\mu_{0}(W^{m}(y_{i}))=\mu_{0}\bigl(\overline{W}^{m}(y_{i})\bigr)$ for each $m\in\mathbb{N}$ (cf. [Li18, Proposition 7.1]), we deduce that $\mu_{0}\bigl(W^{m+1}(y_{i})\bigr)=\mu_{0}(W^{m}(y_{i}))\cdot\bigl(\deg_{F}(y_{i})\big/\deg{F}\bigr)$ for each $m\in\mathbb{N}$. Then it follows from the induction that $\mu_{0}\bigl(W^{m+1}(y_{i})\bigr)=\bigl(\deg_{F}(y_{i})\big/\deg{F}\bigr)^{m}\mu_{0}\bigl(W^{1}(y_{i})\bigr)\leqslant\bigl(\deg_{F}(y_{i})\big/\deg{F}\bigr)^{m}$ for each $m\in\mathbb{N}$. Since $\mathbb{B}_{i}\subseteq W^{1}(y_{i})$, by the definition of $\{U_{n,i,k}\}_{k\in\mathbb{N}}$, $U_{n,i,k}\subseteq W^{k}(y_{i})$ for each $k\in\mathbb{N}$. Hence, $\mu_{0}(U_{n,i,k})\leqslant\mu_{0}\bigl(W^{k}(y_{i})\bigr)\leqslant\bigl(\deg_{F}(y_{i})\big/\deg{F}\bigr)^{k-1}=t_{n,i,k}$ for each $k\in\mathbb{N}$.
Therefore, we have shown Claim 3.
Now we turn to show the original statement. Indeed, by Claims 1, 2, and 3, it follows from Theorem 5.9 that $\{\mu_{n}\}_{n\in\mathbb{N}}$, with $\mu_{n}=\mu_{0}$ for all $n\in\mathbb{N}$, is a sequence of uniformly computable measures. Thus, $\mu_{0}$ is a computable measure. ∎
As concrete applications of Theorem 1.4, we analyze two expanding Thurston maps $g_{1}$ and $g_{2}$ with periodic critical points. We verify that both maps satisfy the hypotheses of Theorem 1.4, thereby establishing the computability of their measures of maximal entropy $\mu_{1}$ and $\mu_{2}$ (Corollary 6.11). Crucially, since both maps have periodic critical points, their measure-theoretic entropy functions fail to be upper semi-continuous (Theorem 6.4), precluding a direct application of Theorem 1.1.
Our first example $g_{1}$ is derived from the barycentric subdivision rule as defined in [BM17, Example 12.21]. We construct a sphere $S_{\triangle}$ by gluing two equilateral (Euclidean) triangles along their boundaries (see Figure 6.1). These two triangles serve as 0-tiles, with their shared boundary forming a $g_{1}$-invariant Jordan curve containing the postcritical set $\{A,\,B,\,C\}$ of $g_{1}$. Subdividing each 0-tile via bisectors yields six smaller triangles, producing twelve $1$-tiles. The map $g_{1}$ sends each $1$-tile linearly onto a 0-tile, and has a fixed critical point. Recursively, $n$-tiles are generated by analogously subdividing (n1)(n-1)-tiles via bisectors, and are Euclidean triangles. It is not difficult to see that the diameters of $n$-tiles decay to zero as $n$ tends to $\infty$, confirming that $g_{1}$ is an expanding Thurston map.
[…]
Figure 6.1. An expanding Thurston map from the barycentric subdivision rule.
To further demonstrate the scope of Theorem 1.4 while illustrating the complexity of expanding Thurston maps, we construct a more intricate example $g_{2}$. The map $g_{2}$ acts on the same polyhedral sphere $S_{\triangle}$ with the postcritical set $\{A,\,B,\,C\}$. While $g_{1}$ subdivides each 0-tile into six congruent smaller triangles, $g_{2}$ implements a finer subdivision rule: each 0-tile splits into eight triangles following the structure illustrated in Figure 6.2. The map $g_{2}$ linearly maps each $1$-tile onto a 0-tile and has a fixed critical point. Recursive iteration of this subdivision generates $n$-tiles whose diameters decay to zero as $n$ tends to $\infty$, showing that $g_{2}$ is an expanding Thurston map.
Now we apply Theorem 1.4 to show that their measures of maximal entropy $\mu_{1}$ and $\mu_{2}$ are computable.
We first define the topology of these two equilateral triangles by embedding them into $\mathbb{R}^{2}$. Let $d_{\triangle}$ denote the geodesic metric on $S_{\triangle}$, and let $\mathcal{S}(S_{\triangle})$ be an effective enumeration of the set $\mathbb{Q}(S_{\triangle})$ of points in $S_{\triangle}$ with rational coordinates. Moreover, all marked points in Figure 6.1, specifically A,B,C,D,E,F,G,HA,\,B,\,C,\,D,\,E,\,F,\,G,\,H, are computable (recall Definition 3.5). Since $g_{1}$ acts as an affine map on each $1$-tile, we may compute $g_{1}(x)$ for each point $x$ within a $1$-tile by applying the corresponding matrix transformation. By Definition 3.12, this ensures that $g_{1}\colon S^{2}\rightarrow S^{2}$ is computable.
[…]
Similarly, we can verify all conditions in Theorem 1.4 in the case where $f\coloneqq\widetilde{g}_{2}$. Therefore, by a proof that is verbatim the same as the above, except for replacing $g_{1},\,\widetilde{g}_{1},\,\mu_{1},\,\widetilde{\mu}_{1}$ with $g_{2},\,\widetilde{g}_{2},\,\mu_{2},\,\widetilde{\mu}_{2}$, it follows that $\mu_{2}$ is computable.
[…]
Figure 6.2. Expanding Thurston map $g_{2}$.
Corollary 6.11.
For each $i\in\{1,\,2\}$, the measure of maximal entropy of the expanding Thurston map $g_{i}$ is computable.